### v6

Abstract
This paper studies the convergence of a spatial semidiscretization of a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. For non-smooth initial data, the regularity of the mild solution is investigated, and an error estimate is derived within the spatial $L^{2}$-norm setting. In the case of smooth initial data, two error estimates are established within the framework of general spatial $L^{q}$-norms.
Keywords: stochastic Allen-Cahn equation, spatial semidiscretization, discrete stochastic maximal $L^{p}$-regularity
1 Introduction
Let $\mathcal{O}$ be a bounded convex polyhedral domain in $\mathbb{R}^{3}$, and let $\Delta$ be the Laplace operator with homogeneous Dirichlet boundary condition. We consider the following stochastic Allen-Cahn equation:
$\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0<t\leqslant T,$ (1.1a)
$y(0)=v,$ (1.1b)
where the initial value $v$ is given and $W_{H}$ is an infinite-dimensional Wiener process.
The stochastic Allen-Cahn equations are widely used to model phase transitions in materials science; see e.g., [3, 9, 13, 23, 26]. They are also an important class of nonlinear stochastic partial differential equations from a mathematical perspective. Recently, researchers in the field of numerical analysis have shown considerable interest in this topic; see [1, 2, 4, 5, 8, 12, 19, 20, 22, 23, 24, 31] and the references therein. We briefly summarize some closely related works of numerical analysis of three-dimensional stochastic Allen-Cahn equations as follows. Feng et al. [8] analyzed finite element approximations of a stochastic Allen-Cahn equation with gradient-type multiplicative noise. Majee and Prohl [22] derived strong convergence rates for a stochastic Allen-Cahn equation driven by a multiplicative finite-dimensional Wiener process. Bréhier et al. [4] analyzed an explicit temporal splitting numerical scheme for a stochastic Allen-Cahn equation driven by an additive infinite-dimensional Wiener process. Jentzen and Pusnik [12] proposed a numerical method for approximating a class of semilinear stochastic evolution equations with non-globally Lipschitz continuous nonlinearities. Qi and Wang [24] derived strong approximation errors for a stochastic Allen-Cahn equation driven by an additive infinite-dimensional Wiener process. Liu and Qiao [20] established a general theory of the numerical analysis for the stochastic Allen-Cahn equations driven by a multiplicative infinite-dimensional Wiener process. Qi et al. [25] proposed an efficient approximation method to the stochastic Allen-Cahn equation with random diffusion coefficient field and multiplicative noise.
The main motivation of this work includes two components:
  • To the best of our knowledge, the existing works on the numerical analysis of the three-dimensional stochastic Allen-Cahn equation have not adequately addressed scenarios involving non-smooth initial data. This scenario is of practical significance, as initial phases may encompass randomly distributed particles. The inclusion of non-smooth initial data introduces heightened complexity to numerical analysis when compared to cases with smooth initial data. Firstly, the theoretical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise and non-smooth initial data is relatively scarce (see, e.g., [17, 18]). Secondly, the singularity introduced by non-smooth initial data significantly complicates the process of numerical analysis.
  • In the existing literature, the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise primarily focuses on the spatial $L^{2}$-norm. However, error estimates employing the more general spatial $L^{q}$-norms are relatively scarce.
Accordingly, the main contributions of this work are as follows:
  • For the model problem Eq. 1.1 with non-smooth initial data $v\in L^{\infty}$, we derive some regularity estimates within the framework of the general spatial $L^{q}$-norms, which are pivotal for the numerical analysis. Specifically, we affirm that
    $y\in L^{p}(\Omega;C([0,T];L^{q})),\quad\forall p,q\in(2,\infty).$
    Our analysis uses a suite of primary tools, including: (i) stochastic integration and Itô’s formula in UMD spaces [7]; (ii) the stochastic maximal $L^{p}$-regularity estimates developed in [29, 30]; (iii) the theory of stochastic partial differential equations with Lyapunov conditions from [17, 18]; (iv) the maximal $L^{p}$-regularity theory for deterministic parabolic equations [32].
  • To handle the singularity caused by the non-smooth initial data, by decomposing the error into a singular part and a relatively non-singular part, we derive the error estimate
    $\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\big{(}h^{\alpha/p}+h^{2/p}t^{-1/p}\big{)},\quad\forall t\in(0,T],$
    where $p\in(2,\infty)$ and $\alpha\in(1,2)$ can be arbitrarily chosen. This result is a stochastic counterpart of the well-known error estimate in [28, Theoerem 3.2].
  • Given smooth initial data $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we mainly use the standard theory of finite element methods (see, e.g., [6, 28]), the discrete maximal $L^{p}$-regularity [15], and the discrete stochastic maximal $L^{p}$-regularity [16] to derive two rigorous error estimates within the framework of general spatial $L^{q}$-norms. Specifically, for arbitrary parameters $p\in(2,\infty)$ and $q\in[2,\infty)$, and for any chosen $\epsilon>0$, we establish the following error estimates:
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},$
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
To achieve a second-order spatial convergence rate, the previous numerical analysis in the literature typically enforce stringent regularity assumptions on the diffusion term. However, in this study, we impose only some moderate regularity conditions on $F$ (see Section 2). Furthermore, this work introduces a novel approach to the numerical analysis of finite element method-based spatial semidiscretizations for a broad class of nonlinear stochastic parabolic equations.
The rest of this paper is organized as follows. Section 2 introduces some conventions and definitions used in this paper, and the spatial semidiscretization analyzed in this research. Section 3 presents the main results of this paper, including the regularity of Eq. 1.1 with non-smooth initial data and the convergence of the spatial semidiscretization under both non-smooth and smooth initial conditions. Section 4 provides a rigorous proof of the aforementioned results. Lastly, Section 5 summarizes the key findings of this work.
The rest of this paper is organized as follows. Section 2 establishes the foundational notation, definitions, and preliminaries that underpin the analysis within this study, as well as detailing the spatial semidiscretization that is the focus of our investigation. Section 3 is dedicated to the principal findings of this work, which encompass the treatment of the regularity of the solution to (1.1) with nonsmooth initial data, along with the convergence analysis of the spatial semidiscretization scheme under both nonsmooth and smooth initial conditions. Section 4 furnishes a detailed and exacting proof of the results previously outlined. Finally, Section 5 offers a concise synthesis of the pivotal insights and contributions that this research brings to the field.
2 Preliminaries
Notational Conventions. Throughout this paper, we will adhere to the following notational conventions. For Banach spaces $E_{1}$ and $E_{2}$, we denote the space of all bounded linear operators from $E_{1}$ to $E_{2}$ as $\mathcal{L}(E_{1},E_{2})$, and it is abbreviated to $\mathcal{L}(E_{1})$ when $E_{1}=E_{2}$. The identity map is denoted by $I$. For any $\theta\in(0,1)$ and $p\in(1,\infty)$, we use $(\cdot,\cdot)_{\theta,p}$ to represent the real interpolation space constructed by the $K$-method, as described in [21, Chapter 1]. The conjugate exponent of $p$ is denoted by $p^{\prime}$. For $q\in[1,\infty]$, we briefly write $L^{q}(\mathcal{O})$ as $L^{q}$. Let $W_{0}^{1,\infty}(\mathcal{O})$ and $W^{2,\infty}(\mathcal{O})$ denote two standard Sobolev spaces (see, e.g., [14]). Let $(\Omega,\mathcal{F},\mathbb{P})$ represent a complete probability space equipped with a filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ is defined. The expectation operator associated with this probability space is denoted by $\mathbb{E}$. For $p\in(1,\infty)$ and a Banach space $X$, we use $L^{p}_{\mathbb{F}}(\Omega\times(0,T);X)$ to denote the space of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega\times(0,T);X)$. For any Banach space $X$, the notation $\gamma(H,X)$ denotes the space of all $\gamma$-radonifying operators from $H$ to $X$ (see [11, Chapter 9]). Finally, for each $t\in\mathbb{R}_{+}$, define $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ by
$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H,$
where $H$ is a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$.
Definition of Operator $F$. Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of continuous functions mapping from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$, satisfying the following conditions for every $n\in\mathbb{N}$:
$f_{n}(x,0)=0,$ $\forall x\in\partial\mathcal{O},$ (2.1)
$\lvert{f_{n}(x,y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert),$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.2)
$\lvert{f_{n}(x,y_{1})-f_{n}(x,y_{2})}\rvert\leqslant C_{F}\lvert{y_{1}-y_{2}}\rvert,$ $\forall x\in\mathcal{O},\,\forall y_{1},y_{2}\in\mathbb{R},$ (2.3)
$\lvert{f_{n}(x_{1},y)-f_{n}(x_{2},y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert)\lvert{x_{1}-x_{2}}\rvert,$ $\forall x_{1},x_{2}\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.4)
where $C_{F}$ is a positive constant independent of $n$, $x$, $y$, $x_{1}$, $x_{2}$, $y_{1}$, and $y_{2}$. Hereafter, with a slight abuse of notation, we will also employ $f_{n}$ to denote the Nemytskii operator defined by
$f_{n}(v):=f_{n}(\cdot,v(\cdot)),\quad\forall v\in L^{1}.$
Let $\{\lambda_{n}\}_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying
$\sum_{n\in\mathbb{N}}\lambda_{n}<\infty.$ (2.5)
Given $v\in L^{2}$, we define $F(v)\in\gamma(H,L^{2})$ via
$F(v):=\sum_{n=0}^{\infty}\sqrt{\lambda}_{n}h_{n}\otimes f_{n}(v).$ (2.6)
Consequently, the model equation Eq. 1.1 can be explicitly stated as follows:
$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+\sum_{n=0}^{\infty}\sqrt{\lambda_{n}}f_{n}(y(t))\,\mathrm{d}\beta_{n}(t),\quad 0<t\leqslant T,\\ y(0)=v.\end{cases}$
Spatial Semidiscretization. Let $\mathcal{K}_{h}$ represent a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, consisting of three-dimensional simplexes. We define the finite element space $X_{h}$ as
$X_{h}:=\left\{v_{h}\in C(\overline{\mathcal{O}})\,\middle|\,v_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\text{ and }v_{h}\text{ vanishes on }\partial\mathcal{O}\right\},$
where $C(\overline{\mathcal{O}})$ is the space of continuous functions over the closure of $\mathcal{O}$, $r\geqslant 1$ is an integer denoting the polynomial degree, and $P_{r}(K)$ is the space of polynomials defined on simplex $K$ up to degree $r$. Let $h$ denote the maximum diameter of the elements in $\mathcal{K}_{h}$. Let $P_{h}$ be the $L^{2}$-orthogonal projection operator onto $X_{h}$, and we define the discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ by
$\int_{\mathcal{O}}(\Delta_{h}u_{h})\cdot v_{h}\,\mathrm{d}\mu=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}\mu,\quad\text{for all }u_{h},v_{h}\in X_{h},$
where $\mu$ represents the three-dimensional Lebesgue measure on $\mathcal{O}$. The spatial semidiscretization considered in this work is given by
$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}(y_{h}^{3})\right)(t)\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad t>0,$ (2.7a)
$y_{h}(0)=P_{h}v,$ (2.7b)
where $v$ is the given initial value.
Mild Solutions. Let $p\in(1,\infty)$ and $q\in[2,\infty)$. We define the convolution operator $S_{0}$ acting on any function $g\in L^{p}(0,T;L^{q})$ as follows:
$(S_{0}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad 0\leqslant t\leqslant T.$ (2.8)
Similarly, for any $g\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,L^{q})))$, we introduce the stochastic convolution operator $S_{1}$ defined by
$(S_{1}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad 0\leqslant t\leqslant T.$ (2.9)
An $\mathbb{F}$-adapted process $y$ is called a mild solution to Eq. 1.1 if $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ and satisfies
$y=G+S_{0}(y-y^{3})+S_{1}F(y),$ (2.10)
where $G$ is defined as
$G(t):=e^{t\Delta}v,\quad t\in[0,T].$ (2.11)
In the context of the spatial semidiscretization, for any $g_{h}\in L^{p}(0,T;X_{h})$, we define
$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(t)\,\mathrm{d}t,\quad t\in[0,T].$
For any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega;L^{2}(0,T;\gamma(H,X_{h})))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(t)\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
An $\mathbb{F}$-adapted $X_{h}$-valued stochastic process $y_{h}$ is called a mild solution to the spatial semidiscretization Eq. 2.7 if $y_{h}\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);X_{h})$ and it satisfies that
$y_{h}=G_{h}+S_{0,h}(y_{h}-P_{h}y_{h}^{3})+S_{1,h}P_{h}F(y_{h}),$
where $G_{h}$ is given by
$G_{h}(t):=e^{t\Delta_{h}}P_{h}v,\quad t\in[0,T].$ (2.12)
3 Main results
In the remainder of this paper, the symbol $c$ represents a generic positive constant that depends solely on $T$, $\mathcal{O}$, $\sum_{n\in\mathbb{N}}\lambda_{h}$, the regularity parameters of the triangulation $\mathcal{K}_{h}$, the indices of the pertinent Lebesgue, Sobolev, and interpolation spaces, the initial value $v$ in Eqs. 1.1 and 2.7, and the exponents related to the spatial mesh size $h$. The value of $c$ may differ from line to line. In addition, for any $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the standard norm defined as follows:
$\lVert{v}\rVert_{\dot{H}^{\alpha,q}}:=\lVert{(-\Delta)^{\alpha/2}v}\rVert_{L^{q}},\quad\forall v\in\dot{H}^{\alpha,q}.$
The main results of this paper are summarized in the following three theorems.
Theorem 3.1.
Assume that $v\in L^{\infty}$. Then equation Eq. 1.1 admits a unique mild solution $y$, which satisfies that
$y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\qquad\forall p,q\in(2,\infty).$ (3.1)
Furthermore, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the solution $y$ possesses the following regularity properties:
$y\in L^{p}\big{(}\Omega;C([0,T];L^{q})\big{)},$ (3.2)
$y\in L^{p}\big{(}\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}\big{)},$ (3.3)
$y-G\in L^{p}\big{(}\Omega;C([0,T];\dot{H}^{1-\epsilon,q})\big{)},$ (3.4)
$y-G\in L^{p}\big{(}\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q}\big{)},$ (3.5)
where $G$ is defined by Eq. 2.11.
Theorem 3.2.
Assume that $v\in L^{\infty}$. Let $y$ and $y_{h}$ be the mild solutions of Eqs. 1.1 and 2.7, respectively. Then, for any $\alpha\in(1,2)$ and $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\left(h^{\alpha/p}+h^{2/p}t^{-1/p}\right).$ (3.6)
Theorem 3.3.
Assume that $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Let $y$ and $y_{h}$ be the mild solutions of Eqs. 1.1 and 2.7, respectively. Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the following error estimates hold:
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},$ (3.7)
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$ (3.8)
Remark 3.1.
For Theorem 3.1, we have the following remarks.
  1. 1.
    Liu [17] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to our model yields the result that
    $y\in L^{p}(\Omega;C([0,T];\dot{H}^{1,2})),\quad\forall p\in(2,\infty),$
    provided that $v\in\dot{H}^{1,2}$. However, the theory in [17] does not cover the case of non-smooth initial data. In the context of the stochastic Allen-Cahn equation with additive noise, Bréhier et al. [4, Lemma 3.1] provide some a priori estimates with the general spatial $L^{q}$ norms.
  2. 2.
    Assume that the condition Eq. 2.1 is removed. By an approximation argument we conclude that Eq. 3.1 remains valid. Moreover, for any $p\in(2,\infty)$ and $q\in[2,\infty)$, the following assertions hold:
    $y-G\in L^{p}(\Omega\times(0,T);\dot{H}^{1,q}),$
    $y-G\in L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p})).$
Remark 3.2.
For Theorems 3.2 and 3.3, we have the following remarks.
  1. 1.
    In the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative infinite-dimensional Wiener process, the existing literature typically imposes stringent regularity conditions on the diffusion coefficient. Our chosen formulation of the function $F$ adheres to the criteria specified in Assumption 2.2 of Liu and Qiao’s work [20]; however, it fails to satisfy the more stringent requirements imposed by their Assumption 2.3. Examining the numerical approximations of Eq. 1.1 under less restrictive conditions on $F$, particularly the relaxation of condition (2.1), holds significant theoretical and practical implications.
  2. 2.
    In Section 4.3, we demonstrate that the norm of $y_{h}$ in $L^{p}(\Omega\times(0,T);L^{q})$ is uniformly bounded with respect to the spatial mesh size $h$ for all $p,q\in[2,\infty)$, provided that $v$ is an element of $W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ (see Eq. 4.51). It can be rigorously shown that this result remains valid even if $v$ is an element of $W_{0}^{\epsilon,\infty}(\mathcal{O})$, where $\epsilon>0$ can be any arbitrarily small positive number. This conclusion is crucial for the numerical analysis of the fully discrete approximations of Eq. 1.1. However, the question of whether this conclusion extends to the case where $v$ is an element of $L^{\infty}$ remains unresolved.
4 Proofs
The objective of this section is to affirm the validity of Theorems 3.1, 3.2, and 3.3. The structure of this section is as follows:
  • Subsection 4.1 introduces key lemmas that serve as foundational results.
  • Subsection 4.2 focuses on the proof of Theorem 3.1, which establishes the regularity of the mild solution of Eq. 1.1.
  • Subsection 4.3 is dedicated to the proofs of Theorems 3.2 and 3.3, which concern the error estimates of the spatial semidiscretization.
4.1 Preliminary lemmas
Lemma 4.1.
Let $0\leqslant\beta<\gamma\leqslant 1$ and $q\in(1,\infty)$. Then the following continuous embeddings hold:
  • (i)
    $\dot{H}^{\gamma,q}$ is continuously embedded in $(L^{q},\dot{H}^{1,q})_{\beta,q}$;
  • (ii)
    $(L^{q},\dot{H}^{1,q})_{\gamma,q}$ is continuously embedded in $\dot{H}^{\beta,q}$.
Lemma 4.2.
Let $0\leqslant\alpha\leqslant\beta<\infty$ and $q\in(1,\infty)$. Then for all $t\in(0,T]$ we have
$\lVert{e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})}\leqslant ct^{(\alpha-\beta)/2}.$
Lemma 4.3.
Let $p,q\in(1,\infty)$. Then for any $g\in L^{p}(0,T;L^{q})$, the following maximal $L^{p}$-regularity estimate holds:
$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0}g}\Big{\rVert}_{L^{p}(0,T;L^{q})}+\lVert{S_{0}g}\rVert_{L^{p}(0,T;\dot{H}^{2,q})}+\lVert{S_{0}g}\rVert_{C([0,T];(L^{q},\dot{H}^{2,q})_{1-1/p,p})}$
[…]
Lemma 4.4.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, the following stochastic maximal $L^{p}$-regularity estimate holds:
$\lVert{S_{1}g}\rVert_{L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1}g}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1,q})}$
[…]
For any $\alpha\in\mathbb{R}$ and $q\in(1,\infty)$, let $\dot{H}_{h}^{\alpha,q}$ denote the space $X_{h}$ equipped with the norm
$\lVert{v_{h}}\rVert_{\dot{H}_{h}^{\alpha,q}}:=\lVert{(-\Delta_{h})^{\alpha/2}v_{h}}\rVert_{L^{q}},\quad\forall v_{h}\in X_{h}.$
Lemma 4.5.
Let $p,q\in(1,\infty)$. Then for any $g_{h}\in L^{p}(0,T;X_{h})$, the following discrete maximal $L^{p}$-regularity estimate holds:
$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0,h}g_{h}}\Big{\rVert}_{L^{p}(0,T;\dot{H}_{h}^{0,q})}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\\ \leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
Lemma 4.6.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,X_{h}))$, the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
$\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,q})}$ (4.1)
[…]
Lemma 4.7.
The following assertions are valid:
  1. (i)
    For all $q\in[1,\infty]$, it holds that
    $\lVert{P_{h}}\rVert_{\mathcal{L}(L^{q})}\text{ is uniformly bounded with respect to $h$}.$ (4.2)
  2. (ii)
    For all $\alpha\in[0,2]$ and $q\in(1,\infty)$, $P_{h}$ has the following approximation property:
    $\lVert{I-P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha}.$ (4.3)
  3. (iii)
    For all $0\leqslant\alpha<\beta\leqslant 2$ and $q\in(1,\infty)$, the following inequality holds:
    $\lVert{\Delta_{h}P_{h}-P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{\beta,q},\dot{H}_{h}^{\alpha-2,q})}\leqslant ch^{\beta-\alpha}.$ (4.4)
  4. (iv)
    For all $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$, the following inequalities are satisfied:
    $\lVert{e^{t\Delta_{h}}}\rVert_{\mathcal{L}(\dot{H}_{h}^{0,q},\dot{H}_{h}^{\alpha,q})}\leqslant ct^{-\alpha/2},$ (4.5)
    $\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}+\lVert{(I-P_{h})e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}$
    $\leqslant c\min\Big{\{}h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\Big{\}}.$ (4.6)
Remark 4.1.
Lemma 4.1 is established by Proposition 1.4 from [21] and Theorem C.4.1 in [10]. The inequality asserted in Lemma 4.2 is a direct consequence of Theorem 2.12 and Equation (2.128) as presented in [33]. The proof of Lemma 4.3 is provided in [32] and Corollary 1.14 from [21]. The proof of Lemma 4.4 is detailed in [30]. The proof of Lemma 4.5 is accessible in [15], and the proof of Lemma 4.6 can be found in [16]. The assertions in Lemma 4.7 are widely recognized and can be found in standard references such as [6, 28].
Lemma 4.8.
For any $q\in(1,\infty)$, the following inequalities hold:
$\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),$ $\forall v\in L^{q};$ (4.7)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}),$ $\forall v\in\dot{H}^{1,q};$ (4.8)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{s_{1},q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{s_{2},q}}),$ $\forall v\in\dot{H}^{s_{2},q},\,\forall 0\leqslant s_{1}<s_{2}<1;$ (4.9)
$\lVert{F(u)-F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},$ $\forall u,v\in L^{q}.$ (4.10)
Proof.
We only prove Eq. 4.9, the proof of Eqs. 4.7, 4.8 and 4.10 being easier. Let $\Omega_{\gamma}$ denote a probability space on which a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of independent standard Gaussian random variables is defined, and let $\mathbb{E}_{\gamma}$ denote the expectation operator with respect to $\Omega_{\gamma}$. The rest of the proof is divided into two steps.
Step 1. We proceed to demonstrate that for any $\theta\in(0,1)$ and $N\in\mathbb{N}_{>0}$, the following inequality holds:
$\lVert{A_{N}v}\rVert_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}\big{)},\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q},$ (4.11)
where
$A_{N}v:=\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}f_{n}(v),\quad\forall v\in L^{q}.$
For any $v\in\dot{H}^{1,q}$, we infer that
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\Big{(}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}\mu\Big{)}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\Big{(}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q}$
$=c\Big{(}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\Big{(}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}(1+\lvert{v}\rvert+\lvert{\nabla v}\rvert)^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\Big{)}^{1/q},$
where step $(\mathrm{i})$ invokes Theorem 6.2.6 from [11], and step $(\mathrm{ii})$ relies on Lipschitz conditions Eqs. 2.3 and 2.4. Using the summability condition Eq. 2.5 and the standard inequality
$\left(\int_{\mathcal{O}}|v|^{q}+|\nabla v|^{q}\,\mathrm{d}\mu\right)^{1/q}\leqslant c\lVert{v}\rVert_{\dot{H}^{1,q}},$
we further deduce that
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}).$
Analogously, employing Eq. 2.2 yields
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),\quad\forall v\in L^{q},$
and utilizing Eq. 2.3 results in
$\lVert{A_{N}u-A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},\quad\forall u,v\in L^{q}.$
Thus, following the methodology outlined in Lemma 28.1 of [27], one obtains
$\lVert{A_{N}v}\rVert_{(L^{q}(\Omega_{\gamma};L^{q}),L^{q}(\Omega_{\gamma};\dot{H}^{1,q}))_{\theta,q}}\leqslant c(1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}),\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q}.$
This result, when combined with the continuous embedding property
$\left(L^{q}(\Omega_{\gamma};L^{q}),L^{q}(\Omega_{\gamma};\dot{H}^{1,q})\right)_{\theta,q}\hookrightarrow L^{q}\left(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q}\right),$
yields
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}\big{)},\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q}.$
Applying Theorem 6.2.6 from [11] once again, we arrive at the assertion in Eq. 4.11.
Step 2. For any $\theta\in(0,1)$ and $v\in(L^{q},\dot{H}^{1,q})_{\theta,q}$, we consider the equality
$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}=\lVert{\lim_{N\to\infty}A_{N}v}\rVert_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})},$
which stems directly from the definition of $F$ as stated in (2.6) and is justified by Theorem 9.1.19 in [11]. By invoking the inequality (4.11), we apply Fatou’s Lemma to deduce that
$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c(1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}).$
Consequently, the desired inequality (4.9) is established through the application of Lemma 4.1. This concludes the proof. ∎
Remark 4.2.
As established by Theorem 4.36 in [21], the interpolation space $(L^{2},\dot{H}^{1,q})_{s,2}$ coincides with $\dot{H}^{s,2}$, sharing equivalent norms across the interval $s\in(0,1)$. Consequently, we refine the estimate Eq. 4.9 to the following form:
$\|F(v)\|_{\gamma(H,\dot{H}^{s,2})}\leqslant c(1+\|v\|_{\dot{H}^{s,2}}),\quad\forall v\in\dot{H}^{s,2},\,\forall s\in(0,1).$
4.2 Proof of Theorem 3.1
It is a well-established fact that the space $L^{2}$ admits an orthonormal basis $\{\phi_{n}\}_{n\in\mathbb{N}}$, comprised of eigenvectors of the negative Laplace operator $-\Delta$. Furthermore, this basis $\{\phi_{n}\}_{n\in\mathbb{N}}$ forms a complete orthogonal system within the space $\dot{H}^{s,2}$ for all $s\in(0,\infty)$.
Lemma 4.9.
For each $n\in\mathbb{N}$ and $v_{n}\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}$, the stochastic Allen-Cahn equation
$\mathrm{d}y_{n}(t)=\big{(}\Delta y_{n}+y_{n}-y_{n}^{3}\big{)}(t)\,\mathrm{d}t+F(y_{n}(t))\,\mathrm{d}W_{H}(t),\quad 0<t\leqslant T,$ (4.12a)
$y_{n}(0)=v_{n}$ (4.12b)
has a unique mild solution. Moreover, for all $p,q\in(2,\infty)$, the following inequality holds:
$\lVert{y_{n}}\rVert_{L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})}\leqslant c(1+\lVert{v_{n}}\rVert_{L^{q}}).$ (4.13)
Proof.
We split the proof into the following two steps.
Step 1. Let $G_{n}(t):=e^{t\Delta}v_{n}$, $t\in[0,T]$. Since $v_{n}\in\dot{H}^{2,q}$ for each $q\in(2,\infty)$, by Lemma 4.2 with $\alpha=\beta=2$, we have
$G_{n}\in L^{p}(0,T;\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.14)
In view of Eq. 4.8 and the fact that $v_{n}\in\dot{H}^{1,2}$, we use [17, Theorem 1.1] to deduce that Eq. 4.12 has a unique mild solution $y_{n}$ and
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,2}),\quad\forall p\in(2,\infty).$ (4.15)
The continuous embedding of $\dot{H}^{1,2}$ in $L^{6}$ then leads to
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{6}),\quad\forall p\in(2,\infty),$
which implies $y_{n}-y_{n}^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for each $p\in(2,\infty)$. Hence, using Lemma 4.3 gives
$S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.16)
Since Eqs. 4.8 and 4.15 imply that $F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for each $p\in(2,\infty)$, by Lemma 4.4 we obtain
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.17)
Considering the mild solution representation
$y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n}),$ (4.18)
and combining Eqs. 4.14, 4.16 and 4.17, we conclude that
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$
From the continuous embeddings of $\dot{H}^{2,2}$ into $\dot{H}^{1,6}$ and $\dot{H}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$, it follows that
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\quad\forall p,q\in(2,\infty),$
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,6}),\quad\forall p\in(2,\infty).$
Hence, Eqs. 4.16 and 4.17 can be improved to
$S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$ (4.19)
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$ (4.20)
Combining Eqs. 4.14, 4.18, 4.19 and 4.20 yields
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty),$
thereby implying, through the continuous embedding of $\dot{H}^{2,6}$ into $\dot{H}^{1,q}$ for all $q\in(2,\infty)$, that
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q}),\quad\forall p,q\in(2,\infty).$
This regularity result allows us to enhance Eq. 4.17 to
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$
which, combined with Eqs. 4.14, 4.18 and 4.19, implies
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.21)
Step 2. Let $p,q\in(2,\infty)$ be arbitrary but fixed. Considering the regularity assertion Eq. 4.21 and the fact
$y_{n}(t)=v_{n}+\int_{0}^{t}(\Delta y_{n}+y_{n}-y_{n}^{3})(s)\,\mathrm{d}s+\int_{0}^{t}F(y_{n}(s))\,\mathrm{d}W_{H}(s),\quad t\in[0,T],$
we use Itô’s formula in UMD Banach spaces (see [7, Theorem 2.4]) to conclude that, for any $t\in(0,T]$,
$\mathbb{E}\Big{(}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}-\lVert{v_{n}}\rVert_{L^{q}}^{p}\Big{)}$
[…]
$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lambda_{k}\lvert{f_{k}(y_{n}(s))}\rvert^{2}\,\mathrm{d}\mu\,\mathrm{d}s$
$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)f_{k}(y_{n}(s))\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s$
[…]
Considering the inequality
$\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)(\Delta y_{n}-y_{n}^{3})(s)\,\mathrm{d}\mu$
[…]
we infer the bound for $I_{1}$:
$I_{1}\leqslant p\,\mathbb{E}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s.$
Regarding $I_{2}+I_{3}$, by using Hölder’s inequality in conjunction with Young’s inequality, and employing the summability condition Eq. 2.5 alongside the growth condition Eq. 2.2, a straightforward calculation leads to the following bound for $I_{2}+I_{3}$:
$I_{2}+I_{3}\leqslant c\,\mathbb{E}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s.$
Combining the above bounds for $I_{1}$ and $I_{2}+I_{3}$, we arrive at
$\mathbb{E}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{v_{n}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s,\quad\forall t\in[0,T].$
Applying Gronwall’s lemma to the above inequality yields
$\lVert{y_{n}(t)}\rVert_{L^{p}(\Omega;L^{q})}\leqslant c\big{(}1+\lVert{v_{n}}\rVert_{L^{q}}\big{)},\quad\forall t\in[0,T],$
which implies the desired inequality Eq. 4.13. This completes the proof. ∎
Proof of Theorem 3.1. We split the proof into the following two steps.
Step 1. Let us prove that Eq. 1.1 admits a unique mild solution $y$ and Eq. 3.1 holds. Let $p,q\in(2,\infty)$ be arbitrary but fixed. For each $n\in\mathbb{N}$, let $y_{n}$ be the mild solution of Eq. 4.12 with $v_{n}$ defined by
$v_{n}:=\mathop{\mathrm{argmin}}_{w\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}}\lVert{v-w}\rVert_{L^{3q}}.$
For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y_{m}-y_{n}$. By definition, we infer that, for any $t\in[0,T]$,
$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+y_{n}^{3}-y_{m}^{3})(t)\,\mathrm{d}t+\big{(}F(y_{m}(t))-F(y_{n}(t))\big{)}\,\mathrm{d}W_{H}(t).$
Observing that the integral over $\mathcal{O}$ of $\lvert{e_{m,n}(t)}\rvert^{q-2}e_{m,n}(t)(y_{n}^{3}-y_{m}^{3})(t)$ is non-positive for all $t\in[0,T]$, and following the proof of Eq. 4.13, we can deduce that
$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{v_{m}-v_{n}}\rVert_{L^{3q}}.$
On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $\{\phi_{n}\}_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $\{v_{n}\}_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$, converging to $v$. Consequently, there exists a unique $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
$\lim_{n\to\infty}y_{n}=y\quad\text{ in }L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q}).$ (4.22)
For any $n\in\mathbb{N}$, application of Hölder’s inequality yields
$\lVert{y_{n}^{3}-y^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
Therefore, using the convergence established in Eq. 4.22, we deduce that
$\lim_{n\to\infty}y_{n}^{3}=y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),$
consequently leading to
$\lim_{n\to\infty}y_{n}-y_{n}^{3}=y-y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Using Lemma 4.3, this directly implies the following convergence for $S_{0}(y_{n}-y_{n}^{3})$:
$\lim_{n\to\infty}S_{0}(y_{n}-y_{n}^{3})=S_{0}(y-y^{3})\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Furthermore, invoking Eqs. 4.10 and 4.22 yields
$\lim_{n\to\infty}F(y_{n})=F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q})),$
and subsequently by Lemma 4.4 we infer the following convergence for $S_{1}F(y_{n})$:
$\lim_{n\to\infty}S_{1}F(y_{n})=S_{1}F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Additionally, applying Lemma 4.2 with $\alpha=\beta=0$ and noting the convergence $\lim_{n\in\infty}\lVert{v_{n}-v}\rVert_{L^{q}}=0$, we establish the following convergence for $G_{n}$:
$\lim_{n\to\infty}G_{n}=G\quad\text{ in }L^{p}(0,T;L^{q}),$
where $G$ is defined by Eq. 2.11, and $G_{n}(t):=e^{t\Delta}v_{n}$ for each $t\in[0,T]$. Finally, combining the derived convergences of $y_{n}$, $G_{n}$, $S_{0}(y_{n}-y_{n}^{3})$, and $S_{1}F(y_{n})$, and recalling that $y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n})$, by passing to the limit $n\to\infty$ we obtain
$y=G+S_{0}(y-y^{3})+S_{1}F(y)\quad\text{in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
This confirms that $y$ is a mild solution of Eq. 1.1, and it is evident that Eq. 3.1 holds. The uniqueness of the mild solution can be proved similarly as in [17, Theorem 1.1].
Step 2. Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. By virtue of the regularity assertion in Eq. 3.1 and the growth estimate Eq. 4.7 for $F$, we deduce that $y-y^{3}$ belongs to $L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})$ and $F(y)$ is an element of $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$. Hence, the application of Lemma 4.3, in conjunction with Lemma 4.1, gives that
$S_{0}(y-y^{3})\in L^{p}\big{(}\Omega;L^{p}(0,T;\dot{H}^{2,q})\cap C([0,T];\dot{H}^{2-2/p-\epsilon,q})\big{)},\quad\forall\epsilon>0,$ (4.23)
while Lemma 4.4 implies $S_{1}F(y)\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q})$. Additionally, Lemma 4.2 ensures that $G\in L^{p}(0,T;\dot{H}^{2/p-\epsilon,q})$ for all $\epsilon>0$. Consequently, by the mild solution representation Eq. 2.10, we arrive at
$y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}),\quad\forall\epsilon>0,$
which confirms the regularity assertion Eq. 3.3. In conjunction with Eq. 4.9, we further infer that
$F(y)\in L_{\mathbb{F}}^{p}\big{(}\Omega\times(0,T);\gamma(H,\dot{H}^{2/p-\epsilon,q})\big{)},\quad\forall\epsilon>0,$
leading to the enhanced conclusion via Lemmas 4.1 and 4.4 that
$S_{1}F(y)\in L^{p}\big{(}\Omega;L^{p}(0,T;\dot{H}^{1+2/p-\epsilon,q})\cap C([0,T];\dot{H}^{1-\epsilon,q})\big{)},\quad\forall\epsilon>0.$ (4.24)
Finally, combining Eqs. 2.10, 4.23 and 4.24 establishes the regularity assertions in Eqs. 3.4 and 3.5. Moreover, the regularity assertion in Eq. 3.2 follows directly from Eq. 3.4 and the observation that $G\in C([0,T];L^{q})$. This completes the proof of Theorem 3.1.
$\blacksquare$
4.3 Proof of Theorems 3.2 and 3.3
We first summarize some standard results as follows.
Lemma 4.10.
Let $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;X_{h})$, the following inequality holds for all $\epsilon>0$:
$\max_{t\in[0,T]}\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-\frac{2}{p}-\epsilon,q}}\leqslant c\|g_{h}\|_{L^{p}(0,T;L^{q})}.$
Proof.
For any $t\in[0,T]$, recognizing that $S_{0,h}g_{h}(t)=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s$, we proceed as follows:
$\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-\frac{2}{p}-\epsilon,q}}$ $=\left\|\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}$
$\leqslant\int_{0}^{t}\left\|e^{(t-s)\Delta_{h}}g_{h}(s)\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\text{(i)}}}{{\leqslant}}c\int_{0}^{t}(t-s)^{-1+\frac{1}{p}+\frac{\epsilon}{2}}\|g_{h}(s)\|_{L^{q}}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\text{(ii)}}}{{\leqslant}}c\Big{(}\int_{0}^{t}(t-s)^{(-1+\frac{1}{p}+\frac{\epsilon}{2})\frac{p}{p-1}}\,\mathrm{d}s\Big{)}^{1-\frac{1}{p}}\|g_{h}\|_{L^{p}(0,T;L^{q})},$
where step (i) invokes Eq. 4.5 with $\alpha=2-2/p-\epsilon$, and step (ii) uses Hölder’s inequality. The uniform boundedness of the integral
$\int_{0}^{t}(t-s)^{\left(-1+\frac{1}{p}+\frac{\epsilon}{2}\right)\frac{p}{p-1}}\,\mathrm{d}s$
with respect to $t\in[0,T]$ then establishes the lemma. ∎
Lemma 4.11.
Assume that $v\in L^{\infty}$. Let $y$ and $y_{h}$ be the mild solutions of Eqs. 1.1 and 2.7, respectively, and let $G$ and $G_{h}$ be defined by Eqs. 2.11 and 2.12, respectively. For simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$. Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following inequalities hold:
$\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}+\lVert{P_{h}G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}\leqslant ch^{\frac{2}{p}},$ (4.25)
$\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\varepsilon},$ (4.26)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\varepsilon},$ (4.27)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{1-\varepsilon},$ (4.28)
$\mathbb{E}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\leqslant ch^{2-\varepsilon}.$ (4.29)
Furthermore, for any $\beta\in\Big{(}1,\frac{3p}{3p-4}\Big{)}$, the following inequalities hold:
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p},$ (4.30)
$\mathbb{E}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\varepsilon}.$ (4.31)
Proof.
The inequality Eq. 4.25 is evident by Eq. 4.6, and the inequality Eq. 4.26 follows by Eq. 4.3 and the regularity assertion Eq. 3.5. The inequality Eq. 4.27 is derived as follows:
[…]
The inequality Eq. 4.28 is derived in a similarly manner:
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
[…]
The inequality Eq. 4.29 is straightforward to verify by applying Hölder’s inequality, in conjunction with the approximation property of $P_{h}$ from Eq. 4.3 and the regularity assertions in Theorem 3.1.
Now let us prove Eq. 4.30. By invoking Eq. 4.2 and replicating the reasoning behind [18, Equation (2.2)], we establish that
$\mathbb{E}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{\dot{H}^{1,2}}^{2}\,\mathrm{d}t$
is uniformly bounded with respect to $h$. Using the continuous embedding of $\dot{H}^{1,2}$ into $L^{6}$, this uniform bound extends to
$\mathbb{E}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2}\,\mathrm{d}t.$
Employing Hölder’s inequality, we deduce for all $t\in[0,T]$ that
$\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{\frac{4\beta}{3\beta-3}}\leq\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2},$
which in turn confirms the uniform bound on
$\lVert{y_{h}}\rVert_{L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})}.$
Subsequently, applying Hölder’s inequality once more yields
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t$
$\leqslant\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}\lVert{y_{h}}\rVert_{L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})}^{p}$
$\leqslant c\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}.$
Invoking the inequality Eq. 4.25 with $p$ replaced by $4p\beta(4\beta-(3\beta-3)p)$, we find
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}(1-\frac{1}{\beta})p}.$
Given that Theorem 3.1 ensures $y$ resides within $L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})$, we can mirror the preceding argument to deduce that
$\mathbb{E}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\leqslant ch^{2-\frac{3}{2}(1-1/\beta)p}.$
The inequality Eq. 4.30 is thus substantiated by combining the above derived inequalities.
Finally, by invoking Eq. 4.27 and adhering to the same logical progression as demonstrated in the proof of Eq. 4.30, we deduce Eq. 4.31. This concludes the proof. ∎
Proof of Theorem 3.2. Fix any $\alpha\in(1,2)$ and $p\in(2,\infty)$. Let
$z(t)$ $:=y(t)-G(t),\quad 0\leqslant t\leqslant T,$ (4.32)
$z_{h}(t)$ $:=y_{h}(t)-G_{h}(t),\quad 0\leqslant t\leqslant T,$ (4.33)
where $G$ and $G_{h}$ are defined by Eqs. 2.11 and 2.12, respectively. From Eq. 1.1 and the fact $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
$\mathrm{d}z(t)=\big{(}\Delta z+z+G-y^{3}\big{)}(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Consequently,
$\mathrm{d}P_{h}z(t)=P_{h}(\Delta z+z+G-y^{3})(t)\,\mathrm{d}t+P_{h}F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Similarly, from Eq. 2.7 and the fact $G_{h}^{\prime}(t)=\Delta_{h}G_{h}(t)$ for $t\in[0,T]$, it follows that
$\mathrm{d}z_{h}(t)=(\Delta_{h}z_{h}\!+\!z_{h}\!+\!G_{h}\!-\!P_{h}y_{h}^{3})(t)\,\mathrm{d}t+P_{h}F\big{(}y_{h}(t)\big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Defining $e_{h}:=z_{h}-P_{h}z$, we then obtain
$\mathrm{d}e_{h}(t)$ $=\Big{(}\Delta_{h}e_{h}+\Delta_{h}P_{h}z-P_{h}\Delta z+e_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.34)
$\quad P_{h}\Big{(}F\big{(}y_{h}(t))-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Defining $\eta_{h}:=e_{h}-\xi_{h}$, where $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}z-P_{h}\Delta z)$, and considering the differential equation
$\Big{(}\frac{\mathrm{d}}{\mathrm{d}t}-\Delta_{h}\Big{)}\xi_{h}(t)=(\Delta_{h}P_{h}z-P_{h}\Delta z)(t),\quad t\geqslant 0,$
we use Eq. 4.34 to conclude that
$\mathrm{d}\eta_{h}(t)$ $=\Big{(}\Delta_{h}\eta_{h}+\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.35)
$\quad P_{h}\Big{(}F\big{(}y_{h}(t))-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
By applying Itô’s formula and considering the initial condition $\eta_{h}(0)=0$ and the definition of $F$ given by Eq. 2.6, we deduce that for any $0\leqslant t\leqslant T$,
$\mathbb{E}\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ $=p\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Big{[}\Delta_{h}\eta_{h}(s)+\xi_{h}(s)+\eta_{h}(s)$
$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{]}\,\mathrm{d}\mu\,\mathrm{d}s$
$\qquad{}+\frac{p}{2}\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s$
$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s$
$=I_{1}+I_{2}+I_{3}+I_{4}+I_{5}+I_{6},$
where
$I_{1}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{2}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{3}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{4}$ $:=p\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s,$
$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s,$
$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{[}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{]}^{2}\,\mathrm{d}s.$
Next, let us estimate $I_{1}$, $I_{2}$, $I_{3}$, and $I_{4}$. For the term $I_{1}$, by the definition of $\Delta_{h}$, we establish that
$I_{1}=-p\mathbb{E}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\|{\eta_{h}(s)}\|_{\dot{H}^{1,2}}^{2}\,\mathrm{d}s.$ (4.36)
Regarding $I_{2}$, we can deduce that
$I_{2}$ $\leqslant\,c\mathbb{E}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\big{(}\|{\xi_{h}(s)}\|_{L^{2}}^{2}+\|{\eta_{h}(s)}\|_{L^{2}}^{2}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{2}\big{)}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\,\mathbb{E}\int_{0}^{t}\|{\xi_{h}(s)}\|_{L^{2}}^{p}+\|{\eta_{h}(s)}\|_{L^{2}}^{p}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{p}\,\mathrm{d}s$
=$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{=}}c\int_{0}^{t}\|{\xi_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}+\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}+\|{(G_{h}-P_{h}G)(s)}\|_{L^{2}}^{p}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,$ (4.37)
where (i) follows from Young’s inequality, (ii) uses the determinism of $G$ and $G_{h}$, and (iii) is justified by the inequalities (4.25) and (4.27). For $I_{3}$, by employing Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ as stated in Eq. 4.2, a straightforward calculation yields
$I_{3}$ $\leqslant c\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\Big{(}\lVert{y(s)}\rVert_{L^{8}}^{2}+\lVert{P_{h}y(s)}\rVert_{L^{8}}^{2}\Big{)}\,\mathrm{d}s$
$\leqslant c\,\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\lVert{y(s)}\rVert_{L^{8}}^{2}\,\mathrm{d}s$
$\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\,\mathrm{d}s.$
By utilizing Eq. 4.29, we arrive at the following bound:
$I_{3}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,\quad\forall\epsilon>0.$ (4.38)
To estimate $I_{4}$, we proceed as follows. Using the definitions in (4.32) and (4.33), along with the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we deduce that $P_{h}y-y_{h}=-\xi_{h}-\eta_{h}+P_{h}G-G_{h}$. For any $s\in[0,t]$, we can then write:
$\int_{\mathcal{O}}\eta_{h}(s)\left[(P_{h}y)^{3}-y_{h}^{3}\right](s)\,\mathrm{d}\mu$
[…]
Hence, applying Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ from (4.2), we obtain the following bound for $I_{4}$:
$I_{4}$ $\leqslant c\mathbb{E}\int_{0}^{t}\bigg{[}\|\eta_{h}(s)\|_{L^{2}}^{p}+\left(\|\xi_{h}(s)\|^{p}_{L^{2\beta^{\prime}}}+\|(P_{h}G-G_{h})(s)\|^{p}_{L^{2\beta^{\prime}}}\right)$
×$\qquad\qquad\qquad{}\times\Big{(}\|y(s)\|^{p}_{L^{2\beta}}+\|y_{h}(s)\|^{p}_{L^{2\beta}}\Big{)}\bigg{]}\,\mathrm{d}s,\quad\forall\beta\in\big{(}1,\frac{3p}{3p-4}\big{)}.$
Using (4.30), we further refine this bound to:
$I_{4}\leqslant c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s+ch^{2-\frac{3}{2}p\left(1-\frac{1}{\beta}\right)},\quad\forall\beta\in\left(1,\frac{3p}{3p-4}\right).$
By selecting $\beta$ sufficiently small within $\big{(}1,\frac{3p}{3p-4}\big{)}$ for any $\epsilon>0$, we arrive at the final bound for $I_{4}$:
$I_{4}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s.$ (4.39)
Now, let us proceed to estimate $I_{5}+I_{6}$. Using Hölder’s inequality and the stability property of $P_{h}$ from Eq. 4.2, we obtain the following bound:
$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s.$
Invoking the Lipschitz continuity condition Eq. 2.3 and the summability condition Eq. 2.5, we further deduce that
$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{2}\,\mathrm{d}s.$
Subsequently, by Young’s inequality, we obtain
$I_{5}+I_{6}\leqslant c\mathbb{E}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s.$
Recalling the definitions in Eqs. 4.32 and 4.33 and considering the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
$(y_{h}-y)(s)=\xi_{h}(s)+\eta_{h}(s)-(z-P_{h}z)(s)-(G-G_{h})(s),\quad\forall s\in[0,T],$ (4.40)
which, in conjunction with Eq. 4.25, Eq. 4.26, and Eq. 4.27, yields
$\mathbb{E}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\leqslant ch^{2}+\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
Consequently, we obtain the following bound for $I_{5}+I_{6}$:
$I_{5}+I_{6}\leqslant ch^{2}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$ (4.41)
Finally, by integrating the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as given by Eqs. 4.36, 4.37, 4.38, 4.39 and 4.41, we obtain the following inequality:
$\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}^{p}+\mathbb{E}\int_{0}^{t}\|\eta_{h}(s)\|_{L^{2}}^{p-2}\|\eta_{h}(s)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}s$
[…]
Applying Gronwall’s lemma to the above inequality, we infer that
$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\left[\mathbb{E}\int_{0}^{T}\|\eta_{h}(t)\|_{L^{2}}^{p-2}\|\eta_{h}(t)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}t\right]^{1/p}\leqslant ch^{\alpha/p}.$
Moreover, for any $\epsilon>0$ and $t\in(0,T]$, we establish the following bounds:
$\lVert{(z-P_{h}z)(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{1-\epsilon}$ (by Eqs. 4.3 and 3.4),
[…]
Since the equality Eq. 4.40 implies
$\|(y-y_{h})(t)\|_{L^{p}(\Omega;L^{2})}$ $\leqslant\|(z-P_{h}z)(t)\|_{L^{p}(\Omega;L^{2})}+\|(G-G_{h})(t)\|_{L^{p}(\Omega;L^{2})}$
$\qquad{}+\|\xi_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})},\quad\forall t\in[0,T],$
the combination of these bounds yields the desired error estimate Eq. 3.6. This completes the proof of Theorem 3.2.
$\blacksquare$
Proof of Theorem 3.3. Let $G$ be defined by Eq. 2.11, and let $z$ be defined by Eq. 4.32. Extending the reasoning from the proof of Theorem 3.1, we establish the following enhanced regularity properties for $y$ and $z$:
$y\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0;$ (4.42)
$z\in L^{p}(\Omega\times(0,T);\dot{H}^{2,q}),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty);$ (4.43)
$z\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0.$ (4.44)
Let $G_{h}$ be defined by Eq. 2.12 and let $z_{h}$ be defined by Eq. 4.33. We adopt the same notation introduced in the proof of Theorem 3.2:
$e_{h}$ $:=z_{h}-P_{h}z,$
$\xi_{h}$ $:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z,$
$\eta_{h}$ $:=e_{h}-\xi_{h}.$
It is straightforward to confirm the following decomposition for $y-y_{h}$:
$y-y_{h}=(z-P_{h}z)+(G-G_{h})-(\xi_{h}+\eta_{h}).$ (4.45)
The remainder of the proof is structured into the following five steps.
Step 1. Let us present some key estimates that form the foundation of our subsequent analysis. We begin by establishing the following estimates:
$\left\|z-P_{h}z\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty),$ (4.46)
$\left\|G-G_{h}\right\|_{C([0,T];L^{q})}+\left\|P_{h}G-G_{h}\right\|_{C([0,T];L^{q})}\leqslant ch^{2},\quad\forall q\in[2,\infty).$ (4.47)
Estimate (4.46) is a direct consequence of (4.3) combined with the regularity property (4.43), and estimate (4.47) follows from (4.6) and the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Analogous to the estimate Eq. 4.27, we employ Lemma 4.5, along with (4.4) and the regularity assertion (4.43), to bound $\xi_{h}$ as follows:
$\left\|\xi_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.48)
Adopting the reasoning from Theorem 3.2, we establish the following bound for $\eta_{h}$:
$\max_{t\in[0,T]}\left\|\eta_{h}(t)\right\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.49)
In light of the decomposition Eq. 4.45, combining Eqs. 4.48, 4.49, 4.47 and 4.46 leads to the error estimate
$\left\|y-y_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.50)
Furthermore, for any $p\in(2,\infty)$ and $q\in[2,\infty)$, we deduce the following chain of inequalities:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}{}$ $\lVert{P_{h}y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{3(1/q-1/2)}\lVert{P_{h}(y-y_{h})}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{3(1/q-1/2)}\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{2+3(1/q-1/2)},$
where in (i) we apply the inverse estimate $\lVert{v_{h}}\rVert_{L^{q}}\leqslant ch^{3(1/q-1/2)}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$; in (ii) we utilize the stability property of $P_{h}$ as stated in Eq. 4.2; and in (iii) we employ the error estimate Eq. 4.50. By invoking the regularity assertion Eq. 4.42, we subsequently conclude a uniform boundedness with respect to the spatial mesh size $h$:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\text{ is uniformly bounded w.r.t~{}$h$ for all $p,q\in[2,\infty)$}.$ (4.51)
Step 2. Let us prove
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.52)
Using the equality Eq. 4.35, along with the definitions of the operators $S_{0,h}$ and $S_{1,h}$, we deduce the following decomposition:
$\eta_{h}=I_{1}+I_{2}+I_{3},$ (4.53)
where
$I_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)},$
$I_{2}$ $:=S_{0,h}P_{h}(y^{3}-y_{h}^{3}),$
$I_{3}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
Fix an arbitrary $p\in(2,\infty)$. For the term $I_{1}$, we establish the following bound for any $q\in[2,\infty)$:
$\lVert{I_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,2})}$
$\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\leqslant ch^{2},$
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{q}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,2}}$ for all $v_{h}\in X_{h}$; the subsequent inequality relies on Lemma 4.5; and the final inequality is justified by Eqs. 4.49, 4.47 and 4.48. Turning to $I_{2}$, we establish the bound: for any $q\in[2,\infty)$,
$\lVert{I_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
$\leqslant c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{5/3})}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2})}\Big{(}\lVert{y}\rVert_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}+\lVert{y_{h}}\rVert_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}\Big{)}$
$\leqslant ch^{2},$
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{q}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,5/3}}$ for all $v_{h}\in X_{h}$; the second inequality uses Lemma 4.5 and Eq. 4.2; the third inequality uses Hölder’s ineqality; and the last inequality uses Eqs. 4.42, 4.50 and 4.51. For $I_{3}$ we demonstrate the following bound:
$\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
[…]
where the first inequality employs the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{6}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{1,2}}$ for all $v_{h}\in X_{h}$, and the second inequality invokes Lemma 4.6, along with Eqs. 4.2 and 4.10. Given that the decomposition Eq. 4.53 implies
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant\lVert{I_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}+\lVert{I_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
$\qquad{}+\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})},$
combining the above bounds for $I_{1}$, $I_{2}$ and $I_{3}$ yields Eq. 4.52.
Step 3. Let us proceed to prove the error estimate Eq. 3.7. Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. We start by establishing the following error estimate:
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}+\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{6})}+{}$
[…]
Employing this error estimate and following the proof methodology for bounding $I_{3}$ in Step 2, we further obtain an enhanced bound for $I_{3}$:
$\lVert{I_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2}.$
Integrating this refined bound with those previously derived bounds for $I_{1}$ and $I_{2}$ in Step 2, and considering the decomposition in Eq. 4.53, we conclude that
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2}.$ (4.54)
Furthermore, we infer that
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}+{}$
[…]
Consequently, combining these estimates directly yields the desired error estimate Eq. 3.7.
Step 4. Let us demonstrate the following inequality:
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.55)
To this end, let $p\in(2,\infty)$ and $q\in[2,\infty)$ be arbitrarily chosen. Invoking the definitions of $I_{1}$ and $I_{2}$ from (4.53), in conjunction with Lemma 4.10 and the continuous embedding of $\dot{H}_{h}^{\varepsilon,q}$ into $L^{q}$ for any $\varepsilon>0$, we deduce
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
Relying on (4.47), (4.48), and (4.54), it follows that
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
which, combined with the stability property of $P_{h}$ as stated in (4.2), leads to
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
Moreover, using Hölder’s inequality yields
$\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
which, in light of the regularity result (4.42) and the stability assertion (4.51), simplifies to
$\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
Consequently, we obtain the following bound for $I_{1}+I_{2}$:
$\lVert{I_{1}+I_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
We proceed to bound the term $I_{3}$ from Eq. 4.53 as follows:
$\lVert{I_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
$\leqslant c\lVert{F(y_{h})-F(y)}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
where the first inequality is justified by Lemma 4.6 and the continuous embedding of $(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}$ in $L^{q}$; the second inequality is a consequence of the stability of the projection $P_{h}$ from Eq. 4.2; and the final inequality relies on the Lipschitz continuity of $F$ as stated in Eq. 4.10. Combining the above bounds for $I_{1}+I_{2}$ and $I_{3}$, we arrive at
$\lVert{I_{1}+I_{2}+I_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
[…]
Given the decomposition Eq. 4.53 and the derived error estimate Eq. 3.7 for all $p\in(2,\infty)$ and $q\in[2,\infty)$, the desired inequality Eq. 4.55 is thereby established.
Step 5. We proceed to prove the error estimate Eq. 3.8. Fix arbitrarily $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$. Using the regularity assertion Eq. 4.43 and following the proof technique used for Eq. 4.28, we deduce the following bound for $\xi_{h}$:
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
Employing Eq. 4.3 and the regularity assertion Eq. 4.44, we establish the following bound for $z-P_{h}z$:
$\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
Furthermore, combining Eqs. 4.47 and 4.55 leads to the following bound:
$\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}.$
Given that the decomposition Eq. 4.45 implies
$\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant\|\xi_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}+\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}$
$\qquad{}+\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))},$
the desired error estimate Eq. 3.8 is derived by combining the above bounds. This concludes the proof of Theorem 3.3.
$\blacksquare$
5 Conclusions
In this work, we have rigorously investigated the regularity properties of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise in the presence of non-smooth initial data. For the spatial semidiscretization using the finite element method, we have established a convergence rate of $O(h^{\alpha/p}+h^{2/p}t^{-1/p})$ with non-smooth initial data. Meanwhile, under the presence of smooth initial data, employing the discrete stochastic maximal $L^{p}$-regularity theory, we have achieved a convergence rate of $O(h^{2})$ in the norm of $L^{p}(\Omega\times(0,T);L^{q})$, alongside a pathwise uniform convergence rate of $O(h^{2-\epsilon})$. Significantly, these findings are inclusive of diverse spatial $L^{q}$-norms, enhancing their adaptability and universality. The derived theoretical results set a foundation for the numerical analysis of fully discrete approximations based on the finite element method for the three-dimensional stochastic Allen-Cahn equation with multiplicative noise.