### v7

Abstract
This paper studies the convergence of a spatial semidiscretization of a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. For non-smooth initial data, the regularity of the mild solution is investigated, and an error estimate is derived within the spatial $L^{2}$-norm setting. In the case of smooth initial data, two error estimates are established within the framework of general spatial $L^{q}$-norms.
Keywords: stochastic Allen-Cahn equation, spatial semidiscretization, discrete stochastic maximal $L^{p}$-regularity
1 Introduction
Let $\mathcal{O}$ be a bounded convex domain in $\mathbb{R}^{3}$ with a smooth boundary $\partial\mathcal{O}$, and let $\Delta$ denote the Laplace operator with homogeneous Dirichlet boundary conditions. We consider the following stochastic Allen-Cahn equation:
$\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (1.1a)
$y(0)=v,$ (1.1b)
where the initial value $v$ is given, and $W_{H}$ is an infinite-dimensional $H$-cylindrical Brownian motion.
The stochastic Allen-Cahn equations are widely used to model phase transitions in materials science; see e.g., [4, 10, 15, 27, 33]. They are also an important class of nonlinear stochastic partial differential equations from a mathematical perspective. Recently, researchers in the field of numerical analysis have shown considerable interest in this topic; see [2, 3, 5, 6, 9, 14, 20, 21, 23, 27, 31, 36] and the references therein. We briefly summarize some closely related works of numerical analysis of three-dimensional stochastic Allen-Cahn equations as follows. Feng et al. [9] analyzed finite element approximations of a stochastic Allen-Cahn equation with gradient-type multiplicative noise. Majee and Prohl [23] derived strong convergence rates for a stochastic Allen-Cahn equation driven by a multiplicative finite-dimensional Wiener process. Bréhier et al. [5] analyzed an explicit temporal splitting numerical scheme for a stochastic Allen-Cahn equation driven by an additive infinite-dimensional Wiener process. Jentzen and Pusnik [14] proposed a numerical method for approximating a class of semilinear stochastic evolution equations with non-globally Lipschitz continuous nonlinearities. Qi and Wang [31] derived strong approximation errors for a stochastic Allen-Cahn equation driven by an additive infinite-dimensional Wiener process. Liu and Qiao [21] established a general theory of the numerical analysis for the stochastic Allen-Cahn equations driven by a multiplicative infinite-dimensional Wiener process. Qi et al. [32] proposed an efficient approximation method to the stochastic Allen-Cahn equation with random diffusion coefficient field and multiplicative noise.
The main motivation of this work includes two components:
  • To the best of our knowledge, existing works on the numerical analysis of the three-dimensional stochastic Allen-Cahn equation have not adequately addressed scenarios involving non-smooth initial data. This scenario is of practical significance, as initial phases may include randomly distributed particles. The inclusion of non-smooth initial data introduces heightened complexity to numerical analysis compared to cases with smooth initial data. Firstly, the theoretical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise and non-smooth initial data is relatively scarce (see, e.g., [18, 19]). Secondly, the singularity introduced by non-smooth initial data significantly complicates the process of numerical analysis.
  • In the existing literature, the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise primarily focuses on the spatial $L^{2}$-norm. However, error estimates using the more general spatial $L^{q}$-norms are relatively scarce.
The main contributions of this work are as follows:
  • For the model problem Eq. 1.1 with non-smooth initial data $v\in L^{\infty}$, we derive some regularity estimates within the framework of the general spatial $L^{q}$-norms, which are crucial for the numerical analysis. Specifically, we affirm that
    $y\in L^{p}(\Omega;C([0,T];L^{q})),\quad\forall p,q\in(2,\infty).$
    Our analysis uses a suite of primary tools, including: (i) stochastic integration and Itô’s formula in UMD spaces [7]; (ii) the stochastic maximal $L^{p}$-regularity estimates developed in [25, 26]; (iii) the theory of stochastic partial differential equations with Lyapunov conditions from [18, 19]; (iv) the maximal $L^{p}$-regularity theory for deterministic parabolic equations [37].
  • To handle the singularity caused by the non-smooth initial data, by decomposing the error into a singular part and a relatively non-singular part, we derive the error estimate
    $\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\big{(}h^{\alpha/p}+h^{2/p}t^{-1/p}\big{)},\quad\forall t\in(0,T],$
    where $p\in(2,\infty)$ and $\alpha\in(1,2)$ can be arbitrarily chosen. This result is a stochastic counterpart of the well-known error estimate in [35, Theoerem 3.2].
  • Given smooth initial data $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we mainly use the standard theory of finite element methods (see, e.g., [1, 35]), the discrete maximal $L^{p}$-regularity [16], and the discrete stochastic maximal $L^{p}$-regularity [17] to derive two rigorous error estimates within the framework of general spatial $L^{q}$-norms. Specifically, for arbitrary parameters $p\in(2,\infty)$ and $q\in[2,\infty)$, and for any chosen $\epsilon>0$, we establish the following error estimates:
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},$
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
To achieve a second-order spatial convergence rate, prior numerical analyses in the literature have typically enforced stringent regularity assumptions on the diffusion coefficients. In contrast, this study imposes relatively moderate regularity conditions on the diffusion coefficient $F$ (see Section 2). While our numerical analysis assumes deterministic initial data, extending the results to accommodate random initial data is straightforward.
The remainder of this paper is organized as follows. Section 2 introduces the notational conventions and definitions that are used throughout the paper. Section 3 presents the main results of this work, including the regularity properties of the mild solution to model problem 1.1 with non-smooth initial data, and establishes the convergence of the finite element-based spatial semidiscretization for both non-smooth and smooth initial conditions. A rigorous proof of these results is provided in Section 4. Finally, Section 5 concludes the paper by summarizing the key findings of this study.
2 Preliminaries
Notational Conventions. Throughout this paper, we will use the following conventions. For Banach spaces $E_{1}$ and $E_{2}$, the space of all bounded linear operators from $E_{1}$ to $E_{2}$ is denoted by $\mathcal{L}(E_{1},E_{2})$, and we abbreviate this to $\mathcal{L}(E_{1})$ when $E_{1}=E_{2}$. The identity operator is indicated by $I$. Given $\theta\in(0,1)$ and $p\in(1,\infty)$, the notation $(\cdot,\cdot)_{\theta,p}$ is used to refer to the real interpolation space constructed via the $K$-method, as detailed in Chapter 1 of [22]. The conjugate exponent of each $p\in[0,\infty]$ is denoted by $p^{\prime}$. The symbol $\mathcal{O}$ represents a bounded convex domain in $\mathbb{R}^{3}$ with a boundary $\partial\mathcal{O}$ that belongs to the $C^{2,0}$ class (see Section 6.2 of [11]). For $q\in[1,\infty]$, we abbreviate $L^{q}(\mathcal{O})$ to $L^{q}$, and use $W_{0}^{1,q}(\mathcal{O})$ and $W^{2,q}(\mathcal{O})$ to denote two standard Sobolev spaces (see, e.g., [34]). Let $(\Omega,\mathcal{F},\mathbb{P})$ represent a complete probability space equipped with a right-continuous filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$, on which a sequence of independent $\mathbb{F}$-adapted Brownian motions $(\beta_{n})_{n\in\mathbb{N}}$ is defined. The expectation operator associated with this probability space is denoted by $\mathbb{E}$. Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$. The $H$-cylindrical Brownian motion $W_{H}$ is defined such that, for each $t\in\mathbb{R}_{+}$, $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ is defined by
$W_{H}(t)h:=\sum_{n\in\mathbb{N}}\beta_{n}(t)(h,h_{n})_{H},\quad\forall h\in H.$
For $p\in(1,\infty)$ and a Banach space $X$, the space of all $\mathbb{F}$-adapted processes in $L^{p}(\Omega\times(0,T);X)$ is denoted by $L^{p}_{\mathbb{F}}(\Omega\times(0,T);X)$. For any Banach space $X$, the notation $\gamma(H,X)$ represents the space of all $\gamma$-radonifying operators from $H$ to $X$, as defined in [13, Chapter 9]. For a comprehensive treatment of the theory of stochastic integration with respect to the $H$-cylindrical Wiener process $W_{H}$, we refer the reader to [24].
Definition of the Operator $F$. Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of continuous functions mapping from $\overline{\mathcal{O}}\times\mathbb{R}$ to $\mathbb{R}$, which satisfy the following conditions for every $n\in\mathbb{N}$:
$f_{n}(x,0)=0,$ $\forall x\in\partial\mathcal{O},$ (2.1)
$\lvert{f_{n}(x,y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert),$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.2)
$\lvert{f_{n}(x,y_{1})-f_{n}(x,y_{2})}\rvert\leqslant C_{F}\lvert{y_{1}-y_{2}}\rvert,$ $\forall x\in\mathcal{O},\,\forall y_{1},y_{2}\in\mathbb{R},$ (2.3)
$\lvert{f_{n}(x_{1},y)-f_{n}(x_{2},y)}\rvert\leqslant C_{F}(1+\lvert{y}\rvert)\lvert{x_{1}-x_{2}}\rvert,$ $\forall x_{1},x_{2}\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.4)
where $C_{F}$ is a positive constant independent of $n$, $x$, $y$, $x_{1}$, $x_{2}$, $y_{1}$, and $y_{2}$. Hereafter, with a slight abuse of notation, we will also employ $f_{n}$ to denote the Nemytskii operator defined by
$f_{n}(v):=f_{n}(\cdot,v(\cdot)),\quad\forall v\in L^{1}.$
Let $(\lambda_{n})_{n\in\mathbb{N}}$ be a sequence of non-negative numbers satisfying the summability condition
$\sum_{n\in\mathbb{N}}\lambda_{n}<\infty.$ (2.5)
For any $v\in L^{2}$, we define $F(v)\in\gamma(H,L^{2})$ as
$F(v):=\sum_{n=0}^{\infty}\sqrt{\lambda}_{n}h_{n}\otimes f_{n}(v).$ (2.6)
Accordingly, the model equation (1.1) can be explicitly formulated as follows:
$\begin{cases}\mathrm{d}y(t)=(\Delta y+y-y^{3})(t)\,\mathrm{d}t+\sum_{n=0}^{\infty}\sqrt{\lambda_{n}}f_{n}(y(t))\,\mathrm{d}\beta_{n}(t),\quad 0\leqslant t\leqslant T,\\ y(0)=v.\end{cases}$
Mild Solutions. Let $p,q\in[2,\infty)$. We introduce the deterministic convolution operator $S_{0}$ acting on functions $g\in L^{p}(0,T;L^{q})$ as follows:
$(S_{0}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad\text{for }0\leqslant t\leqslant T.$ (2.7)
Additionally, for any process $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, we define the stochastic convolution operator $S_{1}$ by
$(S_{1}g)(t):=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad\text{for }0\leqslant t\leqslant T.$ (2.8)
Here, $e^{t\Delta}$, for $t\in[0,\infty)$, denotes the analytic semigroup generated by the Laplace operator $\Delta$ subject to homogeneous Dirichlet boundary conditions. A process $y\in L_{\mathbb{F}}^{6}(\Omega\times(0,T);L^{6})$ is called a mild solution to equation (1.1) if it satisfies the equality
$y=G+S_{0}(y-y^{3})+S_{1}F(y),$ (2.9)
where the function $G$ is given by
$G(t):=e^{t\Delta}v,\quad\text{for }t\in[0,T].$ (2.10)
Note that the conditions in the above definition of a mild solution can be relaxed (see, e.g., [29, Chapter 6]). However, the present formulation is adequate for the purposes of this study.
Spatial Semidiscretization. Consider $\mathcal{K}_{h}$ to be a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, composed of three-dimensional simplexes. We introduce the finite element space $X_{h}$ as
$X_{h}:=\Bigl{\{}v_{h}\in C(\overline{\mathcal{O}})\bigm{|}v_{h}=0\text{ on }(\mathcal{O}\setminus\mathcal{O}_{h})\cup\partial\mathcal{O}_{h},\text{ and }v_{h}|_{K}\in P_{r}(K)\text{ for each }K\in\mathcal{K}_{h}\Bigr{\}},$
where $C(\overline{\mathcal{O}})$ denotes the space of continuous functions on the closure of $\mathcal{O}$, $\mathcal{O}_{h}$ is the closure of the union of the simplexes in $\mathcal{K}_{h}$, $r\geqslant 1$ is a positive integer, and $P_{r}(K)$ is the space of polynomials on simplex $K$ up to degree $r$. Let the spatial mesh size $h$ be the maximum diameter of the elements in $\mathcal{K}_{h}$. Define $P_{h}$ as the $L^{2}$-orthogonal projection operator onto $X_{h}$. The discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ is defined by
$\int_{\mathcal{O}}(\Delta_{h}u_{h})\cdot v_{h}\,\mathrm{d}\mu=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}\mu,\quad\text{for all }u_{h},v_{h}\in X_{h},$
where $\mu$ denotes the three-dimensional Lebesgue measure on $\mathcal{O}$. We consider the following spatial semidiscretization:
$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}(y_{h}^{3})\right)(t)\,\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (2.11a)
$y_{h}(0)=P_{h}v,$ (2.11b)
where $v$ is the given initial value. Assume that $v\in L^{2}$. Since this spatial semidiscretization is a finite-dimensional stochastic differential equation (SDE), we can use the classical theory of finite-dimensional SDEs (see, e.g., [28, Theorem 3.21]) to conclude that Eq. 2.11 has a unique strong solution.
3 Main results
In the remainder of this paper, the symbol $c$ denotes a generic positive constant that depends only on $T$, $\mathcal{O}$, $\sum_{n\in\mathbb{N}}\lambda_{n}$, the constant $C_{F}$ in Eqs. 2.1, 2.2, 2.4 and 2.3, the indices of the relevant Lebesgue, Sobolev, and interpolation spaces, the initial value $v$ given in equations Eq. 1.1 and Eq. 2.11, the regularity parameters of the triangulation $\mathcal{K}_{h}$, and the exponents associated with the spatial mesh size $h$. The value of $c$ may differ at each occurrence. For any $\alpha\in[0,\infty)$ and $q\in(1,\infty)$, let $\dot{H}^{\alpha,q}$ denote the domain of $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the standard norm defined as follows:
$\|v\|_{\dot{H}^{\alpha,q}}:=\|(-\Delta)^{\alpha/2}v\|_{L^{q}}\quad\text{for all $v\in\dot{H}^{\alpha,q}$}.$
The dual space of $\dot{H}^{\alpha,q}$ is denoted by $\dot{H}^{-\alpha,q^{\prime}}$. It is well-established that $\Delta$ has a unique extension as a bounded linear operator from $\dot{H}^{\alpha,q}$ to $\dot{H}^{\alpha-2,q}$ for all $\alpha\in[0,2)$. Specifically, when $\alpha=2$, the space $\dot{H}^{2,q}$ is identical to the intersection $W_{0}^{1,q}(\mathcal{O})\cap W^{2,q}(\mathcal{O})$ and is endowed with a norm equivalent to that of $W^{2,q}(\mathcal{O})$ (see [11, Theorem 9.15]). Moreover, for $\alpha=1$, the space $\dot{H}^{1,q}$ coincides with the standard Sobolev space $W_{0}^{1,q}(\mathcal{O})$, and the two norms are equivalent (see [38, Theorem 16.14]).
The main results of this paper are summarized in the following three theorems.
Theorem 3.1.
Assume that $v\in L^{\infty}$. Then equation Eq. 1.1 admits a unique mild solution $y$, which satisfies that
$y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\qquad\forall p,q\in(2,\infty).$ (3.1)
Furthermore, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the solution $y$ possesses the following regularity properties:
$y\in L^{p}\big{(}\Omega;C([0,T];L^{q})\big{)},$ (3.2)
$y\in L^{p}\big{(}\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}\big{)},$ (3.3)
$y-G\in L^{p}\big{(}\Omega;C([0,T];\dot{H}^{1-\epsilon,q})\big{)},$ (3.4)
$y-G\in L^{p}\big{(}\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q}\big{)},$ (3.5)
where $G$ is defined by Eq. 2.10.
Theorem 3.2.
Assume that $v\in L^{\infty}$. Let $y$ be the mild solution of Eq. 1.1, and let $y_{h}$ be the strong solution to the spatial semidiscretization Eq. 2.11. Then, for any $\alpha\in(1,2)$ and $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\left(h^{\alpha/p}+h^{2/p}t^{-1/p}\right).$ (3.6)
Theorem 3.3.
Assume that $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Let $y$ be the mild solution of Eq. 1.1, and let $y_{h}$ be the strong solution of Eq. 2.11. Then, for any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$, the following error estimates hold:
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},$ (3.7)
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$ (3.8)
It is worth noting that the aforementioned theoretical results remain valid under the assumption that the operator $F$ possesses the growth and Lipschitz continuity properties outlined in Lemma 4.5.
Remark 3.1.
Liu [18] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to our model yields
$y\in L^{p}(\Omega;C([0,T];\dot{H}^{1,2})),\quad\forall p\in(2,\infty),$
provided that $v\in\dot{H}^{1,2}$. Furthermore, Liu and Qiao [21] have established certain regularity estimates in the context of the spatial $L^{2}$-norm, and some regularity estimates with the spatial $L^{\infty}$-norm were derived by using Sobolev’s embedding theorem. Regarding the stochastic Allen-Cahn equation with additive noise, Bréhier et al. [5, Lemma 3.1] provided some a priori estimates involving general spatial $L^{q}$ norms.
Remark 3.2.
For Theorems 3.2 and 3.3, we have the following comments.
  1. 1.
    In the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, the existing literature often imposes strict regularity conditions on the diffusion coefficients. For instance, the study in [23] demands that the diffusion coefficients possess bounded second-order derivatives. Our choice of diffusion coefficient $F$ adheres to the criteria outlined in Assumption 2.2 of [21]; however, it does not meet the more stringent requirements stated in Assumption 2.3 of the same reference. Investigating the numerical approximations of Eq. 1.1 under less restrictive conditions on $F$, especially the loosening of condition (2.1), has important theoretical and practical implications.
  2. 2.
    In Section 4.3, we show that the norm of $y_{h}$ in $L^{p}(\Omega\times(0,T);L^{q})$ is uniformly bounded with respect to the spatial mesh size $h$ for all $p,q\in[2,\infty)$, assuming $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ (see Eq. 4.53). It can be rigorously demonstrated that this result also holds when $v\in W_{0}^{\epsilon,\infty}(\mathcal{O})$, where $\epsilon>0$ can be any arbitrarily small positive number. This finding is critical for the numerical analysis of the fully discrete approximations of Eq. 1.1. However, it remains an open question whether this conclusion applies when $v\in L^{\infty}$.
4 Proofs
The objective of this section is to affirm the validity of Theorems 3.1, 3.2, and 3.3. The structure of this section is as follows:
  • Subsection 4.1 introduces key lemmas that serve as foundational results.
  • Subsection 4.2 focuses on the proof of Theorem 3.1, which establishes the regularity of the mild solution of Eq. 1.1.
  • Subsection 4.3 is dedicated to the proofs of Theorems 3.2 and 3.3, which concern the error estimates of the spatial semidiscretization.
4.1 Preliminary Lemmas
It is a well-established result that for any $q\in(1,\infty)$ and $s\in(0,2)$, the space $\dot{H}^{s,q}$ coincides with the complex interpolation space $[L^{q},\dot{H}^{2,q}]_{s/2}$, equipped with equivalent norms. This result is documented, for example, in Chapter 16 of [38]. The combination of Proposition 1.4 from [22] and Theorem C.4.1 in [12] yields the following continuous embeddings.
Lemma 4.1.
Let $0\leqslant\beta<\gamma\leqslant 1$ and $q\in(1,\infty)$. Then the following continuous embeddings hold:
  • (i)
    The space $\dot{H}^{\gamma,q}$ is continuously embedded in $(L^{q},\dot{H}^{1,q})_{\beta,q}$;
  • (ii)
    The space $(L^{q},\dot{H}^{1,q})_{\gamma,q}$ is continuously embedded in $\dot{H}^{\beta,q}$.
Theorem 2.12 and inequality (2.128) in [38] provide the following standard estimate for the growth of the analytic semigroup generated by the Laplace operator with homogeneous Dirichlet boundary conditions.
Lemma 4.2.
Let $0\leqslant\alpha\leqslant\beta<\infty$ and $q\in(1,\infty)$. For all $t\in(0,T]$, the following estimate holds:
$\lVert{e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})}\leqslant ct^{(\alpha-\beta)/2}.$
For the deterministic convolution operator $S_{0}$, defined by Eq. 2.7, we have the following celebrated deterministic maximal $L^{p}$-regularity estimate; see Theorem 6.1.11 in [30] and Corollary 1.14 from [22].
Lemma 4.3.
Let $p,q\in(1,\infty)$. Then for any $g\in L^{p}(0,T;L^{q})$, the following inequality holds:
$\big{\lVert}\frac{\mathrm{d}}{\mathrm{d}t}S_{0}g\big{\rVert}_{L^{p}(0,T;L^{q})}+\lVert{S_{0}g}\rVert_{L^{p}(0,T;\dot{H}^{2,q})}+\lVert{S_{0}g}\rVert_{C([0,T];(L^{q},\dot{H}^{2,q})_{1-1/p,p})}\leqslant c\lVert{g}\rVert_{L^{p}(0,T;L^{q})}.$
For the stochastic convolution operator $S_{1}$, defined by Eq. 2.8, we present the following fundamental stochastic maximal $L^{p}$-regularity estimate; see Theorems 1.1 and 1.2 in [26].
Lemma 4.4.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, the following inequality holds:
$\lVert{S_{1}g}\rVert_{L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1}g}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1,q})}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}.$
The operator $F$, defined by Eq. 2.6, possesses the following growth and Lipschitz continuity properties.
Lemma 4.5.
Let $q\in(1,\infty)$. The following inequalities are valid:
$\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\left(1+\lVert{v}\rVert_{L^{q}}\right),$ for all $\text{for all }v\in L^{q};$ (4.1)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c\left(1+\lVert{v}\rVert_{\dot{H}^{1,q}}\right),$ for all $\text{for all }v\in\dot{H}^{1,q};$ (4.2)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{s_{1},q})}\leqslant c\left(1+\lVert{v}\rVert_{\dot{H}^{s_{2},q}}\right),$ for $\text{for all }v\in\dot{H}^{s_{2},q},\text{ where }0\leqslant s_{1}<s_{2}<1;$ (4.3)
$\lVert{F(u)-F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},$ for all $\text{for all }u,v\in L^{q}.$ (4.4)
Proof.
We focus our demonstration on Eq. 4.3; the proofs of Eq. 4.1, Eq. 4.2, and Eq. 4.4 follow similar, yet more straightforward, lines of reasoning. For the definition of $\gamma$-radonifying operators, we direct the reader to Section 9.1 of [13]. Consider $\Omega_{\gamma}$ as a probability space supporting a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of independent standard Gaussian random variables. Let $\mathbb{E}_{\gamma}$ denote the expectation operator with respect to the probability space $\Omega_{\gamma}$. The subsequent argument is structured in two main steps.
Step 1. We proceed to demonstrate that for any $\theta\in(0,1)$ and $N\in\mathbb{N}_{>0}$, the following inequality holds:
$\lVert{A_{N}v}\rVert_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}\big{)},\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q},$ (4.5)
where
$A_{N}v:=\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}f_{n}(v),\quad\forall v\in L^{q}.$
For any $v\in\dot{H}^{1,q}$, we deduce that
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\bigg{[}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}\mu\bigg{]}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\sqrt{\lambda_{n}}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q}$
$=c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lambda_{n}(1+\lvert{v}\rvert+\lvert{\nabla v}\rvert)^{2}\Big{)}^{q/2}\,\mathrm{d}\mu\bigg{]}^{1/q},$
where step $(\mathrm{i})$ uses the famous Kahane-Khintchine inequality (see Theorem 6.2.6 from [13]), and step $(\mathrm{ii})$ relies on Lipschitz conditions Eqs. 2.3 and 2.4. Using the summability condition Eq. 2.5 and the standard inequality
$\left(\int_{\mathcal{O}}|v|^{q}+|\nabla v|^{q}\,\mathrm{d}\mu\right)^{1/q}\leqslant c\lVert{v}\rVert_{\dot{H}^{1,q}},$
we further deduce that
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}).$
Analogously, employing the growth condition Eq. 2.2 yields
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),\quad\forall v\in L^{q},$
and utilizing the Lipschitz continuous condition Eq. 2.3 results in
$\lVert{A_{N}u-A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},\quad\forall u,v\in L^{q}.$
Thus, following the methodology outlined in Lemma 28.1 of [34], one obtains
$\lVert{A_{N}v}\rVert_{(L^{q}(\Omega_{\gamma};L^{q}),L^{q}(\Omega_{\gamma};\dot{H}^{1,q}))_{\theta,q}}\leqslant c(1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}),\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q}.$
Combining this result with the continuous embedding property
$\left(L^{q}(\Omega_{\gamma};L^{q}),L^{q}(\Omega_{\gamma};\dot{H}^{1,q})\right)_{\theta,q}\hookrightarrow L^{q}\left(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q}\right),$
which can be easily verified by definition, yields
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}\big{)},\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q}.$
Applying the Kahane-Khintchine inequality once again, we arrive at the assertion in Eq. 4.5.
Step 2. For any $\theta\in(0,1)$ and $v\in(L^{q},\dot{H}^{1,q})_{\theta,q}$, we consider the equality
$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}=\Big{\lVert}\lim_{N\to\infty}A_{N}v\Big{\rVert}_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})},$
which follows directly from the definition of $F$ as stated in (2.6) and is justified by Theorem 9.1.19 in [13]. Using the inequality (4.5), we apply Fatou’s Lemma to deduce that
$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c(1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}).$
Consequently, the desired inequality (4.3) is established through the application of Lemma 4.1. This concludes the proof. ∎
Remark 4.1.
As established by Theorem 4.36 in [22], the interpolation space $(L^{2},\dot{H}^{1,2})_{s,2}$ coincides with $\dot{H}^{s,2}$, sharing equivalent norms for all $s\in(0,1)$. Consequently, we refine the inequality Eq. 4.3 to the following form:
$\|F(v)\|_{\gamma(H,\dot{H}^{s,2})}\leqslant c(1+\|v\|_{\dot{H}^{s,2}}),\quad\text{for all $v\in\dot{H}^{s,2}$, where $s\in(0,1)$}.$
4.2 Proof of Theorem 3.1
It is well known that the space $L^{2}$ admits an orthonormal basis $(\phi_{n})_{n\in\mathbb{N}}$, consisting of eigenvectors of the negative Laplace operator $-\Delta$ with homogeneous Dirichlet boundary conditions. This basis $(\phi_{n})_{n\in\mathbb{N}}$ forms a complete orthogonal system in the space $\dot{H}^{s,2}$ for all $s\in(0,\infty)$.
Lemma 4.6.
For each $n\in\mathbb{N}$ and $v_{n}\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}$, the stochastic Allen-Cahn equation
$\mathrm{d}y_{n}(t)=\big{(}\Delta y_{n}+y_{n}-y_{n}^{3}\big{)}(t)\,\mathrm{d}t+F(y_{n}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (4.6a)
$y_{n}(0)=v_{n}$ (4.6b)
has a unique mild solution. Moreover, for all $p,q\in(2,\infty)$, the following inequality holds:
$\lVert{y_{n}}\rVert_{L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})}\leqslant c(1+\lVert{v_{n}}\rVert_{L^{q}}).$ (4.7)
Proof.
We split the proof into the following two steps.
Step 1. Define $G_{n}(t):=e^{t\Delta}v_{n}$ for $t\in[0,T]$. Given that $v_{n}\in\dot{H}^{2,q}$ for every $q\in(2,\infty)$, it follows from Lemma 4.2 with $\alpha=\beta=2$ that
$G_{n}\in L^{p}(0,T;\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.8)
In view of Eq. 4.2 and the fact that $v_{n}\in\dot{H}^{1,2}$, we use [18, Theorem 1.1] (see also, [21, Lemma 2.3 and Theorem 3.1]) to deduce that Eq. 4.6 has a unique mild solution $y_{n}$ and
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,2}),\quad\forall p\in(2,\infty).$ (4.9)
The continuous embedding of $\dot{H}^{1,2}$ in $L^{6}$ then leads to
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{6}),\quad\forall p\in(2,\infty),$
which in turn implies that $y_{n}-y_{n}^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for all $p\in(2,\infty)$. Applying Lemma 4.3, we obtain
$S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.10)
Since (4.2) and (4.9) ensure that $F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for every $p\in(2,\infty)$, Lemma 4.4 allows us to conclude that
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.11)
Considering the mild solution representation
$y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n}),$
and combining (4.8), (4.10), and (4.11), we deduce that
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$
The continuous embeddings of $\dot{H}^{2,2}$ into $\dot{H}^{1,6}$ and $\dot{H}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$ yield
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\quad\forall p,q\in(2,\infty),$
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,6}),\quad\forall p\in(2,\infty).$
Hence, Eqs. 4.10 and 4.11 can be refined to
$S_{0}(y_{n}-y_{n}^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$ (4.12)
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$
Combining these refined regularity results with (4.8), we arrive at
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty),$
which, through the continuous embedding of $\dot{H}^{2,6}$ into $\dot{H}^{1,q}$ for all $q\in(2,\infty)$, implies that
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q}),\quad\forall p,q\in(2,\infty).$
This regularity result allows us to further improve (4.11) to
$S_{1}F(y_{n})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$
which, in conjunction with (4.8) and (4.12), yields
$y_{n}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.13)
Step 2. Let $p,q\in(2,\infty)$ be arbitrary but fixed. Considering the regularity assertion Eq. 4.13 and the fact
$y_{n}(t)=v_{n}+\int_{0}^{t}(\Delta y_{n}+y_{n}-y_{n}^{3})(s)\,\mathrm{d}s+\int_{0}^{t}F(y_{n}(s))\,\mathrm{d}W_{H}(s),\quad t\in[0,T],$
we use Itô’s formula in UMD Banach spaces (see [7, Theorem 2.4]) to conclude that, for any $t\in(0,T]$,
$\mathbb{E}\Big{[}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}-\lVert{v_{n}}\rVert_{L^{q}}^{p}\Big{]}$
[…]
$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\Biggl{[}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lambda_{k}\lvert{f_{k}(y_{n}(s))}\rvert^{2}\,\mathrm{d}\mu\,\mathrm{d}s\Biggr{]}$
$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\Biggl{\{}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\lambda_{k}\bigg{[}\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)f_{k}(y_{n}(s))\,\mathrm{d}\mu\bigg{]}^{2}\,\mathrm{d}s\Biggr{\}}$
[…]
Considering the inequality
$\int_{\mathcal{O}}\lvert{y_{n}(s)}\rvert^{q-2}y_{n}(s)(\Delta y_{n}-y_{n}^{3})(s)\,\mathrm{d}\mu$
[…]
we infer the bound for $I_{1}$:
$I_{1}\leqslant p\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
For $I_{2}+I_{3}$, using Hölder’s inequality in conjunction with Young’s inequality, and employing the summability condition Eq. 2.5 along with the growth condition Eq. 2.2, we derive the following bound:
$I_{2}+I_{3}\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
Combining the above bounds for $I_{1}$ and $I_{2}+I_{3}$, we arrive at
$\mathbb{E}\lVert{y_{n}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{v_{n}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y_{n}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]},\quad\forall t\in[0,T].$
Applying Gronwall’s lemma to the above inequality yields
$\lVert{y_{n}}\rVert_{C([0,T];L^{p}(\Omega;L^{q}))}\leqslant c\big{(}1+\lVert{v_{n}}\rVert_{L^{q}}\big{)},$
which implies the desired inequality Eq. 4.7. This completes the proof. ∎
Proof of Theorem 3.1. We split the proof into the following two steps.
Step 1. We proceed to show that equation Eq. 1.1 has a unique mild solution $y$, and that Eq. 3.1 is satisfied. Let $p,q\in(2,\infty)$ be fixed but arbitrary. For each $n\in\mathbb{N}$, denote by $y_{n}$ the mild solution of equation Eq. 4.6, where $v_{n}$ is defined as
$v_{n}:=\mathop{\mathrm{argmin}}_{w\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}}\lVert{v-w}\rVert_{L^{3q}}.$
For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y_{m}-y_{n}$. By definition, we infer that
$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+y_{n}^{3}-y_{m}^{3})(t)\,\mathrm{d}t+\big{(}F(y_{m}(t))-F(y_{n}(t))\big{)}\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
Observing that the integral over $\mathcal{O}$ of $\lvert{e_{m,n}(t)}\rvert^{q-2}e_{m,n}(t)(y_{n}^{3}-y_{m}^{3})(t)$ is non-positive for all $t\in[0,T]$, and following the proof of Eq. 4.7, we can deduce that
$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{v_{m}-v_{n}}\rVert_{L^{3q}}.$ (4.14)
On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $(\phi_{n})_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $(v_{n})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$, converging to $v$. Consequently, there exists a unique process $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
$\lim_{n\to\infty}y_{n}=y\quad\text{ in }L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q}).$ (4.15)
For any $n\in\mathbb{N}$, applying Hölder’s inequality yields
$\lVert{y_{n}^{3}-y^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
Therefore, using the convergence established in Eq. 4.15, we deduce that
$\lim_{n\to\infty}y_{n}^{3}=y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),$
which consequently leads to
$\lim_{n\to\infty}y_{n}-y_{n}^{3}=y-y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Using Lemma 4.3, this directly implies the following convergence of $S_{0}(y_{n}-y_{n}^{3})$:
$\lim_{n\to\infty}S_{0}(y_{n}-y_{n}^{3})=S_{0}(y-y^{3})\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Furthermore, invoking Eqs. 4.4 and 4.15 yields
$\lim_{n\to\infty}F(y_{n})=F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q})),$
and subsequently by Lemma 4.4 we infer the following convergence for $S_{1}F(y_{n})$:
$\lim_{n\to\infty}S_{1}F(y_{n})=S_{1}F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Additionally, applying Lemma 4.2 with $\alpha=\beta=0$ and noting the convergence $\lim_{n\in\infty}\lVert{v_{n}-v}\rVert_{L^{q}}=0$, we establish the following convergence for $G_{n}$:
$\lim_{n\to\infty}G_{n}=G\quad\text{ in }L^{p}(0,T;L^{q}),$
where $G$ is defined by Eq. 2.10, and $G_{n}(t):=e^{t\Delta}v_{n}$ for each $t\in[0,T]$. Finally, combining the derived convergences of $y_{n}$, $G_{n}$, $S_{0}(y_{n}-y_{n}^{3})$, and $S_{1}F(y_{n})$, and recalling that $y_{n}=G_{n}+S_{0}(y_{n}-y_{n}^{3})+S_{1}F(y_{n})$, we obtain by passing to the limit $n\to\infty$ that
$y=G+S_{0}(y-y^{3})+S_{1}F(y)\quad\text{in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
This confirms that $y$ is a mild solution of Eq. 1.1, and it is evident that Eq. 3.1 holds. To prove the uniqueness of the mild solution, assume $\widetilde{y}$ is another mild solution. Consider the difference $e=y-\widetilde{y}$. Then $e$ satisfies the equation
$\begin{cases}\mathrm{d}e(t)=(\Delta e+e+\widetilde{y}^{3}-y^{3})(t)\,\mathrm{d}t+\big{(}F(y(t))-F(\widetilde{y}(t))\big{)}\,\mathrm{d}W_{H}(t),&t\in[0,T],\\ e(0)=0.\end{cases}$
By an argument similar to that used in Eq. 4.14, we find that
$\lVert{e}\rVert_{L^{2}(\Omega\times(0,T);L^{2})}\leqslant 0,$
which implies that $e=0$ almost surely, confirming the uniqueness of the mild solution to Eq. 1.1.
Step 2. Fix any $p\in(2,\infty)$ and $q\in[2,\infty)$. By virtue of the regularity assertion in Eq. 3.1 and the growth estimate Eq. 4.1 for $F$, we deduce that $y-y^{3}$ belongs to $L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})$ and $F(y)$ is an element of $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$. Hence, the application of Lemma 4.3, in conjunction with Lemma 4.1, gives that
$S_{0}(y-y^{3})\in L^{p}\big{(}\Omega;L^{p}(0,T;\dot{H}^{2,q})\cap C([0,T];\dot{H}^{2-2/p-\epsilon,q})\big{)},\quad\forall\epsilon>0,$ (4.16)
while Lemma 4.4 implies $S_{1}F(y)\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q})$. Additionally, Lemma 4.2 ensures that $G\in L^{p}(0,T;\dot{H}^{2/p-\epsilon,q})$ for all $\epsilon>0$. Consequently, by the mild solution representation Eq. 2.9, we arrive at
$y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}),\quad\forall\epsilon>0,$
which confirms the regularity assertion Eq. 3.3. In conjunction with Eq. 4.3, we further infer that
$F(y)\in L_{\mathbb{F}}^{p}\big{(}\Omega\times(0,T);\gamma(H,\dot{H}^{2/p-\epsilon,q})\big{)},\quad\forall\epsilon>0,$
leading to the enhanced conclusion via Lemmas 4.1 and 4.4 that
$S_{1}F(y)\in L^{p}\big{(}\Omega;L^{p}(0,T;\dot{H}^{1+2/p-\epsilon,q})\cap C([0,T];\dot{H}^{1-\epsilon,q})\big{)},\quad\forall\epsilon>0.$ (4.17)
Finally, combining Eqs. 2.9, 4.16 and 4.17 establishes the regularity assertions in Eqs. 3.4 and 3.5. Moreover, the regularity assertion in Eq. 3.2 follows directly from Eq. 3.4 and the observation that $G\in C([0,T];L^{q})$. This completes the proof of Theorem 3.1.
$\blacksquare$
4.3 Proof of Theorems 3.2 and 3.3
We begin by summarizing several well-established results. As known from [8], it holds that
$\|P_{h}\|_{\mathcal{L}(L^{q})}\text{ is uniformly bounded with respect to }h\text{ for all }q\in(1,\infty).$ (4.18)
Consequently, the operator $P_{h}$ possesses the following approximation property:
$\|I-P_{h}\|_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha}\quad\text{for all }\alpha\in[0,2]\text{ and }q\in(1,\infty).$ (4.19)
Inequality (2.5) from [1], combined with the observation that
$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,2})}\leqslant 1,$
yields, through interpolation, that for any $q\in[2,\infty)$, the operator norm
$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,q})}$
is uniformly bounded with respect to $h$. Utilizing this fundamental stability estimate, the approximation property Eq. 4.19, the interpolation theory as stated in [22, Theorem 2.6], and the fact that $\Delta^{-1}$ is a bounded linear operator mapping from $\dot{H}^{-s,q}$ to $\dot{H}^{2-s,q}$ for all $s\in[0,2]$, standard arguments then lead to the conclusion that
$\|P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}\|_{\mathcal{L}(\dot{H}^{\beta-2,q},\dot{H}^{\alpha,q})}\leqslant ch^{\beta-\alpha}\quad\text{for all }0\leqslant\alpha\leqslant 1\leqslant\beta\leqslant 2.$ (4.20)
Let $e^{t\Delta_{h}}$, for $t\in[0,\infty)$, denote the analytic semigroup generated by the discrete Laplace operator $\Delta_{h}$. Using the established estimates for the resolvents of $\Delta_{h}$ from [1] and applying (2.128) from [38], we obtain the well-known growth estimate: for any $\alpha\in[0,1]$, $q\in(1,\infty)$, and $t\in(0,T]$,
$\lVert{(-\Delta_{h})^{\alpha}e^{t\Delta_{h}}}\rVert_{\mathcal{L}(L^{q})}\leqslant ct^{-\alpha}.$ (4.21)
Furthermore, utilizing techniques analogous to those found in Theorem 6.10 and Lemma 20.7 of [35], one can readily derive the following well-known inequality: for any $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t\in[0,T]$,
$\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}+\lVert{(I-P_{h})e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant c\min\Big{\{}h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\Big{\}}.$ (4.22)
For the special case where $q=2$, the reader is referred to Theorem 3.2 of [35] for a detailed treatment of this result.
Let $q\in(1,\infty)$. For any $\alpha\in\mathbb{R}$, denote by $\dot{H}_{h}^{\alpha,q}$ the space $X_{h}$ equipped with the norm
$\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}:=\|(-\Delta_{h})^{\alpha/2}v_{h}\|_{L^{q}}\quad\text{for all }v_{h}\in X_{h}.$
According to Theorem 3.1 (ii) in [17], the boundedness of the purely imaginary powers of $-\Delta_{h}$ in $\mathcal{L}(\dot{H}_{h}^{0,q})$ is independent of the spatial mesh size $h$. Consequently, by applying Theorem 4.17 from [22], we deduce that for any $\alpha\in(0,2)$,
$c_{0}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant\|v_{h}\|_{[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}}\leqslant c_{1}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h},$ (4.23)
where $c_{0}$ and $c_{1}$ are positive constants independent of $h$. Here, $[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}$ denotes the complex interpolation space between $\dot{H}_{h}^{0,q}$ and $\dot{H}_{h}^{2,q}$ (see [22, Chapter 2]). This result is important as it enables the application of interpolation theory to the space $\dot{H}_{h}^{s,q}$ for all $s\in\mathbb{R}$. For example, given the standard inverse estimate
$\|v_{h}\|_{\dot{H}_{h}^{2,q}}\leqslant ch^{-2}\|v_{h}\|_{L^{q}}\quad\text{for all }v_{h}\in X_{h},$
we can invoke [22, Theorem 2.6] to derive the following inverse estimate: for any $0\leqslant\alpha<\beta<\infty$,
$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant ch^{\alpha-\beta}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h}.$
Furthermore, for any $\alpha\in[0,1]$, there exist positive constants $c_{0}$ and $c_{1}$, independent of $h$, such that
$c_{0}\|v_{h}\|_{\dot{H}^{\alpha,q}}\leqslant\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant c_{1}\|v_{h}\|_{\dot{H}^{\alpha,q}}\quad\text{for all }v_{h}\in X_{h}.$
Henceforth, the aforementioned theoretical results will be utilized without further explicit reference.
We now introduce a discrete deterministic maximal $L^{p}$-regularity estimate and its stochastic counterpart. For a function $g_{h}\in L^{1}(0,T;\dot{H}_{h}^{0,2})$, the discrete deterministic convolution operator $S_{0,h}$ is defined as
$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s,\quad\text{for }t\in[0,T].$ (4.24)
For a process $g_{h}\in L_{\mathbb{F}}^{2}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}W_{H}(s),\quad\text{for }t\in[0,T].$ (4.25)
The proofs of the following two lemmas can be found in [16] for the first and in [17] for the second.
Lemma 4.7.
Suppose $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;X_{h})$, the following discrete deterministic maximal $L^{p}$-regularity estimate is valid:
$\Big{\lVert}{\frac{\mathrm{d}}{\mathrm{d}t}S_{0,h}g_{h}}\Big{\rVert}_{L^{p}(0,T;\dot{H}_{h}^{0,q})}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
Lemma 4.8.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. For any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))$, the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
$\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,q})}$ (4.26)
[…]
Additionally, for the discrete deterministic convolution operator $S_{0,h}$, we establish the following property.
Lemma 4.9.
Let $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;\dot{H}_{h}^{0,q})$, the following inequality holds for all $\epsilon>0$:
$\|S_{0,h}g_{h}(t)\|_{C([0,T];\dot{H}_{h}^{2-2/p-\epsilon,q})}\leqslant c\|g_{h}\|_{L^{p}(0,T;L^{q})}.$
Proof.
For any $t\in[0,T]$, from the identity $S_{0,h}g_{h}(t)=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s$, we deduce that
$\|S_{0,h}g_{h}(t)\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}$ $\leqslant\int_{0}^{t}\left\|e^{(t-s)\Delta_{h}}g_{h}(s)\right\|_{\dot{H}_{h}^{2-2/p-\epsilon,q}}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\text{(i)}}}{{\leqslant}}c\int_{0}^{t}(t-s)^{-1+\frac{1}{p}+\frac{\epsilon}{2}}\|g_{h}(s)\|_{L^{q}}\,\mathrm{d}s$
$\stackrel{{\scriptstyle\text{(ii)}}}{{\leqslant}}c\Big{(}\int_{0}^{t}(t-s)^{(-1+\frac{1}{p}+\frac{\epsilon}{2})\frac{p}{p-1}}\,\mathrm{d}s\Big{)}^{1-\frac{1}{p}}\|g_{h}\|_{L^{p}(0,T;L^{q})},$
where step (i) uses the estimate Eq. 4.21 with $\alpha=2-2/p-\epsilon$, and step (ii) applies Hölder’s inequality. Since the integral
$\int_{0}^{t}(t-s)^{\left(-1+\frac{1}{p}+\frac{\epsilon}{2}\right)\frac{p}{p-1}}\,\mathrm{d}s$
is uniformly bounded with respect to $t\in[0,T]$, the desired inequality follows, thereby completing the proof of the lemma. ∎
We proceed to present some auxiliary estimates.
Lemma 4.10.
Suppose that $v\in L^{\infty}$. Consider the mild solution $y$ to equation (1.1), the strong solution $y_{h}$ to equation (2.11), and the function $G$ defined by equation (2.10). Additionally, let $G_{h}(t):=e^{t\Delta_{h}}P_{h}v$ for $t\in[0,T]$. For notational simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$. For any $p\in(2,\infty)$, $q\in[2,\infty)$, and $\varepsilon>0$, the following inequalities hold:
$\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}+\lVert{P_{h}G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}\leqslant ch^{\frac{2}{p}},$ (4.27)
$\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\varepsilon},$ (4.28)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\varepsilon},$ (4.29)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{1-\varepsilon},$ (4.30)
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\varepsilon}.$ (4.31)
Furthermore, for any $\beta\in\Big{(}1,\frac{3p}{3p-4}\Big{)}$, the following inequalities hold:
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p},$ (4.32)
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\varepsilon}.$ (4.33)
Here, $\beta^{\prime}$ denotes the conjugate exponent of $\beta$.
Proof.
The inequality Eq. 4.27 is evident by Eq. 4.22, and the inequality Eq. 4.28 follows by Eq. 4.19 and the regularity assertion Eq. 3.5. The inequality Eq. 4.29 is derived as follows:
[…]
The inequality Eq. 4.30 is derived in a similarly manner:
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
[…]
The inequality Eq. 4.31 is straightforward to verify by applying Hölder’s inequality, in conjunction with the approximation property of $P_{h}$ from Eq. 4.19 and the regularity assertions in Theorem 3.1.
Now let us prove Eq. 4.32. By invoking Eq. 4.18 and replicating the reasoning behind [19, Equation (2.2)], we establish that
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{\dot{H}^{1,2}}^{2}\,\mathrm{d}t\bigg{]}$
is uniformly bounded with respect to $h$. Using the continuous embedding of $\dot{H}^{1,2}$ into $L^{6}$, this uniform bound extends to
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2}\,\mathrm{d}t\bigg{]}.$
Employing Hölder’s inequality, we deduce for all $t\in[0,T]$ that
$\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{\frac{4\beta}{3\beta-3}}\leqslant\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2},$
which in turn confirms the uniform bound on
$\lVert{y_{h}}\rVert_{L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})}.$
Subsequently, applying Hölder’s inequality once more yields
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}$
$\leqslant\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}\lVert{y_{h}}\rVert_{L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})}^{p}$
$\leqslant c\lVert{P_{h}G-G_{h}}\rVert_{L^{\frac{4p\beta}{4\beta-(3\beta-3)p}}(\Omega\times(0,T);L^{2\beta^{\prime}})}^{p}.$
Invoking the inequality Eq. 4.27 with $p$ replaced by $4p\beta/(4\beta-(3\beta-3)p)$, we find
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}(1-\frac{1}{\beta})p}.$
Given that Theorem 3.1 ensures $y$ belongs to $L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})$, we can replicate the previous argument to conclude that
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}(1-1/\beta)p}.$
Hence, Eq. 4.32 follows by combining the aforementioned inequalities. To deduce Eq. 4.33, we rely on Eq. 4.29 and proceed analogously to the proof of Eq. 4.32. This completes the proof. ∎
Proof of Theorem 3.2. Fix any $\alpha\in(1,2)$ and $p\in(2,\infty)$. Let
$z(t)$ $:=y(t)-G(t),\quad 0\leqslant t\leqslant T,$ (4.34)
$z_{h}(t)$ $:=y_{h}(t)-G_{h}(t),\quad 0\leqslant t\leqslant T,$ (4.35)
where $G$ is defined by Eq. 2.10, and $G_{h}$ is given by
$G_{h}(t):=e^{t\Delta_{h}}P_{h}v,\quad t\in[0,T].$ (4.36)
From Eq. 1.1 and the fact $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
$\mathrm{d}z(t)=\big{(}\Delta z+z+G-y^{3}\big{)}(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Consequently,
$\mathrm{d}P_{h}z(t)=P_{h}(\Delta z+z+G-y^{3})(t)\,\mathrm{d}t+P_{h}F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Similarly, from Eq. 2.11 and the fact $G_{h}^{\prime}(t)=\Delta_{h}G_{h}(t)$ for $t\in[0,T]$, it follows that
$\mathrm{d}z_{h}(t)=(\Delta_{h}z_{h}\!+\!z_{h}\!+\!G_{h}\!-\!P_{h}y_{h}^{3})(t)\,\mathrm{d}t+P_{h}F\big{(}y_{h}(t)\big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Defining $e_{h}:=z_{h}-P_{h}z$, we then obtain
$\mathrm{d}e_{h}(t)$ $=\Big{(}\Delta_{h}e_{h}+\Delta_{h}P_{h}z-P_{h}\Delta z+e_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.37)
$\quad P_{h}\Big{(}F\big{(}y_{h}(t)\big{)}-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Defining $\eta_{h}:=e_{h}-\xi_{h}$, where $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}z-P_{h}\Delta z)$, and considering the differential equation
$\Big{(}\frac{\mathrm{d}}{\mathrm{d}t}-\Delta_{h}\Big{)}\xi_{h}(t)=(\Delta_{h}P_{h}z-P_{h}\Delta z)(t),\quad t\geqslant 0,$
we use Eq. 4.37 to conclude that
$\mathrm{d}\eta_{h}(t)$ $=\Big{(}\Delta_{h}\eta_{h}+\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.38)
$\quad P_{h}\Big{(}F\big{(}y_{h}(t))-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
By applying Itô’s formula and considering the initial condition $\eta_{h}(0)=0$ and the definition of $F$ given by Eq. 2.6, we deduce that for any $0\leqslant t\leqslant T$,
$\mathbb{E}\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ $=p\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Big{(}\Delta_{h}\eta_{h}(s)+\xi_{h}(s)+\eta_{h}(s)$
$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{)}\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]}$
$\qquad{}+\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]}$
$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{)}^{2}\,\mathrm{d}s\biggr{]}$
$=I_{1}+I_{2}+I_{3}+I_{4}+I_{5}+I_{6},$
where
$\allowdisplaybreaks I_{1}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{2}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{3}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{4}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}\mu\,\mathrm{d}s\biggr{]},$
$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]},$
$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\lambda_{k}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}\mu\Big{)}^{2}\,\mathrm{d}s\biggr{]}.$
Next, let us estimate $I_{1}$, $I_{2}$, $I_{3}$, and $I_{4}$. For the term $I_{1}$, by the definition of $\Delta_{h}$, we establish that
$I_{1}=-p\mathbb{E}\biggl{[}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\|{\eta_{h}(s)}\|_{\dot{H}^{1,2}}^{2}\,\mathrm{d}s\biggr{]}.$ (4.39)
Regarding $I_{2}$, by Hölder’s inequality, Young’s inequality, and the inequalities Eq. 4.27 and Eq. 4.29, we can deduce that
$I_{2}\leqslant ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,$ (4.40)
For $I_{3}$, by employing Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ as stated in Eq. 4.18, a straightforward calculation yields
$I_{3}$ $\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\Big{(}\lVert{y(s)}\rVert_{L^{8}}^{2}+\lVert{P_{h}y(s)}\rVert_{L^{8}}^{2}\Big{)}\,\mathrm{d}s\bigg{]}$
$\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\lVert{y(s)}\rVert_{L^{8}}^{2}\,\mathrm{d}s\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\,\mathrm{d}s\bigg{]}.$
By utilizing Eq. 4.31, we arrive at the following bound:
$I_{3}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,\quad\forall\epsilon>0.$ (4.41)
To estimate $I_{4}$, we proceed as follows. Using the definitions in (4.34) and (4.35), along with the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we deduce that $P_{h}y-y_{h}=-\xi_{h}-\eta_{h}+P_{h}G-G_{h}$. For any $s\in[0,t]$, we can then write:
$\int_{\mathcal{O}}\eta_{h}(s)\left[(P_{h}y)^{3}-y_{h}^{3}\right](s)\,\mathrm{d}\mu$
[…]
Hence, applying Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ from (4.18), we obtain the following bound for $I_{4}$:
$I_{4}$ $\leqslant c\mathbb{E}\int_{0}^{t}\bigg{[}\|\eta_{h}(s)\|_{L^{2}}^{p}+\left(\|\xi_{h}(s)\|^{p}_{L^{2\beta^{\prime}}}+\|(P_{h}G-G_{h})(s)\|^{p}_{L^{2\beta^{\prime}}}\right)$
×$\qquad\qquad\qquad{}\times\Big{(}\|y(s)\|^{p}_{L^{2\beta}}+\|y_{h}(s)\|^{p}_{L^{2\beta}}\Big{)}\bigg{]}\,\mathrm{d}s,\quad\forall\beta\in\big{(}1,\frac{3p}{3p-4}\big{)}.$
Using (4.32) and (4.33), we further simplifies this bound to:
$I_{4}\leqslant c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s+ch^{2-\frac{3}{2}p\left(1-\frac{1}{\beta}\right)},\quad\forall\beta\in\left(1,\frac{3p}{3p-4}\right).$
By selecting $\beta$ sufficiently small within $\big{(}1,\frac{3p}{3p-4}\big{)}$ for any $\epsilon>0$, we arrive at the final bound for $I_{4}$:
$I_{4}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s.$ (4.42)
Now, let us proceed to estimate $I_{5}+I_{6}$. Using Hölder’s inequality and the stability property of $P_{h}$ from Eq. 4.18, we obtain the following bound:
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lambda_{k}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
Invoking the Lipschitz continuity condition Eq. 2.3 and the summability condition Eq. 2.5, we further deduce that
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
Subsequently, by Young’s inequality, we obtain
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}.$
Recalling the definitions in Eqs. 4.34 and 4.35 and considering the relation $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
$(y_{h}-y)(s)=\xi_{h}(s)+\eta_{h}(s)-(z-P_{h}z)(s)-(G-G_{h})(s),\quad\forall s\in[0,T],$ (4.43)
which, in conjunction with Eqs. 4.27, 4.28 and 4.29, yields
$\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}\leqslant ch^{2}+\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
Consequently, we obtain the following bound for $I_{5}+I_{6}$:
$I_{5}+I_{6}\leqslant ch^{2}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$ (4.44)
Finally, by integrating the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as given by Eqs. 4.39, 4.40, 4.41, 4.42 and 4.44, we obtain the following inequality:
$\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}^{p}+\mathbb{E}\bigg{[}\int_{0}^{t}\|\eta_{h}(s)\|_{L^{2}}^{p-2}\|\eta_{h}(s)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}s\bigg{]}$
[…]
Applying Gronwall’s lemma to the above inequality, we infer that
$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\left[\mathbb{E}\int_{0}^{T}\|\eta_{h}(t)\|_{L^{2}}^{p-2}\|\eta_{h}(t)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}t\right]^{1/p}\leqslant ch^{\alpha/p}.$
Moreover, for any $\epsilon>0$ and $t\in(0,T]$, we establish the following bounds:
$\lVert{(z-P_{h}z)(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{1-\epsilon}$ (by Eqs. 4.19 and 3.4),
[…]
Since the equality Eq. 4.43 implies
$\|(y-y_{h})(t)\|_{L^{p}(\Omega;L^{2})}$ $\leqslant\|(z-P_{h}z)(t)\|_{L^{p}(\Omega;L^{2})}+\|(G-G_{h})(t)\|_{L^{p}(\Omega;L^{2})}$
$\qquad{}+\|\xi_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})},\quad\forall t\in[0,T],$
the combination of these bounds yields the desired error estimate Eq. 3.6. This completes the proof of Theorem 3.2.
$\blacksquare$
Proof of Theorem 3.3. Let $G$ be defined by Eq. 2.10, and let $z$ be defined by Eq. 4.34. Extending the reasoning from the proof of Theorem 3.1, we establish the following enhanced regularity properties for the processes $y$ and $z$:
$y\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0;$ (4.45)
$z\in L^{p}(\Omega\times(0,T);\dot{H}^{2,q}),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty);$ (4.46)
$z\in L^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p\in(2,\infty),\,\forall q\in[2,\infty),\,\forall\epsilon>0.$ (4.47)
Let $G_{h}$ be defined by Eq. 4.36. We adopt the same notation as introduced in the proof of Theorem 3.2:
$z_{h}$ $:=y_{h}-G_{h},$ $e_{h}:=z_{h}-P_{h}z,$
$\xi_{h}$ $:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z,$ $\eta_{h}:=e_{h}-\xi_{h}.$
The remainder of the proof is structured into the following five steps.
Step 1. Let us present some key estimates that form the foundation of our subsequent analysis. We begin by establishing the following estimates:
$\left\|z-P_{h}z\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty),$ (4.48)
$\left\|G-G_{h}\right\|_{C([0,T];L^{q})}+\left\|P_{h}G-G_{h}\right\|_{C([0,T];L^{q})}\leqslant ch^{2},\quad\forall q\in[2,\infty).$ (4.49)
Estimate (4.48) is a direct consequence of (4.19) combined with the regularity property (4.46). Estimate (4.49) follows from (4.22) and the condition $v\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. In analogy to the estimate (4.29), we utilize Lemma 4.7, along with (4.20) and the regularity assertion (4.46), to derive the following bound for $\xi_{h}$:
$\left\|\xi_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.50)
Following the reasoning from Theorem 3.2, we obtain the following estimate for $\eta_{h}$:
$\max_{t\in[0,T]}\left\|\eta_{h}(t)\right\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.51)
In light of the decomposition Eq. 4.43, combining Eqs. 4.50, 4.51, 4.49 and 4.48 leads to the error estimate
$\left\|y-y_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.52)
Furthermore, for any $p\in(2,\infty)$ and $q\in[2,\infty)$, we deduce the following chain of inequalities:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}{}$ $\lVert{P_{h}y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{3(1/q-1/2)}\lVert{P_{h}(y-y_{h})}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{3(1/q-1/2)}\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}+ch^{2+3(1/q-1/2)},$
where in (i) we apply the inverse estimate $\lVert{v_{h}}\rVert_{L^{q}}\leqslant ch^{3(1/q-1/2)}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$; in (ii) we utilize the stability property of $P_{h}$ as stated in Eq. 4.18; and in (iii) we employ the error estimate Eq. 4.52. By invoking the regularity assertion Eq. 4.45, we subsequently conclude a uniform boundedness with respect to the spatial mesh size $h$:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\text{ is uniformly bounded with respect to $h$ for all $p,q\in[2,\infty)$}.$ (4.53)
Step 2. Let us prove
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.54)
Using the equality Eq. 4.38, along with the definitions of the discrete convolution operators $S_{0,h}$ and $S_{1,h}$ given in Eqs. 4.24 and 4.25, we deduce the following decomposition:
$\eta_{h}=II_{1}+II_{2}+II_{3},$ (4.55)
where
$II_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)},$
$II_{2}$ $:=S_{0,h}P_{h}(y^{3}-y_{h}^{3}),$
$II_{3}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
Fix an arbitrary $p\in(2,\infty)$. For the term $II_{1}$, we establish the following bound for any $q\in[2,\infty)$:
$\lVert{II_{1}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,2})}$
$\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\leqslant ch^{2},$
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,2}}$ for all $v_{h}\in X_{h}$; the subsequent inequality relies on Lemma 4.7; and the final inequality is justified by Eqs. 4.51, 4.49 and 4.50. Turning to $II_{2}$, we establish the bound: for any $q\in[2,\infty)$,
$\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$ $\leqslant c\lVert{S_{0,h}P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
$\leqslant c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{5/3})}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2})}\Big{(}\lVert{y}\rVert_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}+\lVert{y_{h}}\rVert_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}\Big{)}$
$\leqslant ch^{2},$
where the first inequality uses the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{2,5/3}}$ for all $v_{h}\in X_{h}$; the second inequality uses Lemma 4.7 and Eq. 4.18; the third inequality uses Hölder’s inequality; and the last inequality uses Eqs. 4.45, 4.52 and 4.53. For $II_{3}$ we establish the following bound:
$\lVert{II_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant c\lVert{S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,2})}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
[…]
where the first inequality employs the discrete Sobolev’s inequality $\lVert{v_{h}}\rVert_{L^{6}}\leqslant c\lVert{v_{h}}\rVert_{\dot{H}_{h}^{1,2}}$ for all $v_{h}\in X_{h}$, and the second inequality invokes Lemma 4.8, along with Eqs. 4.18 and 4.4. In view of the decomposition Eq. 4.55, combining the above bounds for $II_{1}$, $II_{2}$, and $II_{3}$ yields Eq. 4.54.
Step 3. We proceed to establish the error estimate (3.7). Considering the decomposition of $y_{h}-y$ given in (4.43), we integrate the inequalities (4.48), (4.49), (4.50), and (4.54) to validate (3.7) for the specific case where $p\in(2,\infty)$ and $q=6$. Building upon this outcome and employing a similar bounding technique as utilized for $II_{3}$ in Step 2, we obtain an enhanced bound for $II_{3}$:
$\lVert{II_{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$
Integrating this refined bound with the previously established bounds for $II_{1}$ and $II_{2}$ from Step 2, and incorporating the decomposition (4.55), we arrive at the following conclusion:
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.56)
Ultimately, revisiting the decomposition (4.43) and synthesizing the inequalities (4.48), (4.49), (4.50), and (4.56), we affirm the validity of the error estimate (3.7) for all $p\in(2,\infty)$ and $q\in[2,\infty)$.
Step 4. Let us demonstrate the following inequality:
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2},\quad\forall p\in(2,\infty),\,\forall q\in[2,\infty).$ (4.57)
To this end, let $p\in(2,\infty)$ and $q\in[2,\infty)$ be arbitrarily chosen. Invoking the definitions of $II_{1}$ and $II_{2}$ from (4.55), in conjunction with Lemma 4.9 and the continuous embedding of $\dot{H}_{h}^{1,q}$ into $L^{q}$, we deduce
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant c\lVert{\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
Relying on (4.49), (4.50), and (4.56), it follows that
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{P_{h}(y^{3}-y_{h}^{3})}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
which, combined with the stability property of $P_{h}$ as stated in (4.18), leads to
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}.$
Moreover, using Hölder’s inequality yields
$\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
which, in light of the regularity result (4.45) and the stability assertion (4.53), simplifies to
$\lVert{y^{3}-y_{h}^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
Consequently, we obtain the following bound for $II_{1}+II_{2}$:
$\lVert{II_{1}+II_{2}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}+c\lVert{y-y_{h}}\rVert_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
We proceed to bound the term $II_{3}$ from Eq. 4.55 as follows:
$\lVert{II_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
$\leqslant c\lVert{F(y_{h})-F(y)}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}$
$\leqslant c\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})},$
where the first inequality is justified by Lemma 4.8 and the continuous embedding of $(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}$ in $L^{q}$; the second inequality is a consequence of the stability of the projection $P_{h}$ from Eq. 4.18; and the final inequality relies on the Lipschitz continuity of $F$ as stated in Eq. 4.4. Combining the above bounds for $II_{1}+II_{2}$ and $II_{3}$, we arrive at
$\lVert{II_{1}+II_{2}+II_{3}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
[…]
Given the decomposition Eq. 4.55 and the derived error estimate Eq. 3.7 for all $p\in(2,\infty)$ and $q\in[2,\infty)$, the desired inequality Eq. 4.57 is thereby established.
Step 5. We proceed to prove the error estimate Eq. 3.8. Fix arbitrarily $p\in(2,\infty)$, $q\in[2,\infty)$, and $\epsilon>0$. Using the regularity assertion Eq. 4.46 and following the proof technique used for Eq. 4.30, we deduce the following bound for $\xi_{h}$:
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
Employing Eq. 4.19 and the regularity assertion Eq. 4.47, we establish the following bound for $z-P_{h}z$:
$\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\epsilon}.$
Furthermore, by combining Eqs. 4.49 and 4.57, we obtain the following bound:
$\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2}.$
Since the decomposition Eq. 4.43 implies
$\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant\|\xi_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}+\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}$
$\qquad{}+\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))},$
the desired error estimate Eq. 3.8 follows by combining the above bounds. This concludes the proof of Theorem 3.3.
$\blacksquare$
5 Conclusions
In this work, we have rigorously investigated the regularity properties of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, particularly in the presence of non-smooth initial data. For the spatial semidiscretization using the finite element method, we have established a convergence rate of $O(h^{\alpha/p}+h^{2/p}t^{-1/p})$ with non-smooth initial data. Meanwhile, under the condition of smooth initial data, employing the discrete stochastic maximal $L^{p}$-regularity theory, we have achieved a convergence rate of $O(h^{2})$ in the $L^{p}(\Omega\times(0,T);L^{q})$ norm, along with a pathwise uniform convergence rate of $O(h^{2-\epsilon})$. Significantly, these findings cover a variety of spatial $L^{q}$-norms, enhancing their adaptability and universality. The theoretical results derived here lay a foundation for the numerical analysis of fully discrete approximations based on the finite element method for the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, especially within the context of general spatial $L^{q}$-norms.