### v8

Abstract
This paper studies the convergence of a spatial semidiscretization of a three-dimensional stochastic Allen-Cahn equation with multiplicative noise. For non-smooth initial data, the regularity of the mild solution is investigated, and an error estimate is derived within the spatial $L^{2}$-norm setting. In the case of smooth initial data, two error estimates are established within the framework of general spatial $L^{q}$-norms.
Keywords: stochastic Allen-Cahn equation, spatial semidiscretization, discrete stochastic maximal $L^{p}$-regularity
1 Introduction
Let $\mathcal{O}$ be a bounded, convex domain in $\mathbb{R}^{3}$ with a smooth boundary $\partial\mathcal{O}$, and let $\Delta$ denote the Laplace operator equipped with homogeneous Dirichlet boundary conditions. We consider the stochastic Allen–Cahn equation given by
$\mathrm{d}y(t)=\left(\Delta y+y-y^{3}\right)(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (1.1a)
$y(0)=y_{0},$ (1.1b)
where $y_{0}$ is the given initial condition, $W_{H}$ is an $H$-cylindrical Brownian motion, and the operator $F$ will be specified later.
The stochastic Allen-Cahn equations are widely used to model phase transitions in materials science; see e.g., [4, 11, 18, 26, 29]. They are also an important class of nonlinear stochastic partial differential equations from a mathematical perspective. Recently, researchers in the field of numerical analysis have shown considerable interest in this topic; see [2, 3, 5, 6, 10, 17, 22, 23, 25, 26, 28, 35] and the references therein. We briefly summarize some closely related works of numerical analysis of three-dimensional stochastic Allen-Cahn equations as follows. Feng et al. [10] analyzed finite element approximations of a stochastic Allen-Cahn equation with gradient-type multiplicative noise. Majee and Prohl [25] derived strong convergence rates for a stochastic Allen-Cahn equation driven by a multiplicative Wiener process. Bréhier et al. [5] analyzed an explicit temporal splitting numerical scheme for a stochastic Allen-Cahn equation driven by an additive Wiener process. Jentzen and Pusnik [17] proposed a numerical method for approximating a class of semilinear stochastic evolution equations with non-globally Lipschitz continuous nonlinearities. Qi and Wang [28] derived strong approximation errors for a stochastic Allen-Cahn equation driven by an additive Wiener process. Liu and Qiao [23] established a general theory of the numerical analysis for the stochastic Allen-Cahn equations with multiplicative noise.
The motivation for this work stems from two important gaps in the existing literature:
  • To the best of our knowledge, the numerical analysis of the three-dimensional stochastic Allen–Cahn equation with non-smooth initial data remains underdeveloped, despite its practical significance. Existing studies often impose restrictive regularity assumptions on the initial data. For instance, the analysis in [23] requires that the initial condition $y_{0}$ belongs to the space $\dot{H}^{\alpha,2}$ for some $\alpha\in[1,2]$, while [28] assumes $\dot{H}^{1,2}$-regularity; the precise definition of these spaces is given in Section 2. The treatment of non-smooth initial data introduces substantial theoretical and numerical challenges. On the one hand, there is a lack of rigorous results for the three-dimensional stochastic Allen–Cahn equation with multiplicative noise and rough initial conditions (see, e.g., [20, 21]). On the other hand, the presence of singularities due to non-smooth initial data significantly complicates the numerical analysis.
  • Current numerical studies on the three-dimensional stochastic Allen–Cahn equation with multiplicative noise predominantly focus on error estimates in the $L^{2}$-norm in space. However, the analysis of numerical approximations in more general spatial $L^{q}$-norms (for $q\geqslant 2$) remains largely unaddressed, which limits the scope and applicability of the theoretical results.
The main contributions of this work are summarized as follows:
  • For the model problem Eq. 1.1 with non-smooth initial data $y_{0}\in L^{\infty}(\mathcal{O})$, we establish essential regularity estimates within the general framework of spatial $L^{q}$-norms. These estimates are crucial for the subsequent numerical analysis. Specifically, we show that the mild solution satisfies
    $y\in L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q}(\mathcal{O})))\quad\text{for all }p,q\in(2,\infty).$
    Our analysis combines key tools from both stochastic and deterministic PDE theory, including: stochastic integration and Itô’s formula in UMD Banach spaces [8]; stochastic maximal $L^{p}$-regularity estimates [32, 33]; the Lyapunov framework for stochastic partial differential equations [20, 21]; and maximal $L^{p}$-regularity estimates for deterministic evolution equations [36].
  • To address the singularity induced by the non-smooth initial data, we employ a decomposition technique that separates the error into a singular and a more regular component. This allows us to derive the error estimate
    $\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2}(\mathcal{O}))}\leqslant c\big{(}h^{2/p-\epsilon}+h^{2/p}t^{-1/p}\big{)},\quad\forall t\in(0,T],\,p\in(2,\infty),$
    where $y_{h}$ is the strong solution of the spatial semidiscretization, $\epsilon>0$ can be arbitrarily small, and $c>0$ is a constant independent of $t$ and the spatial mesh size $h$. This result can be viewed as a stochastic counterpart of the classical deterministic error estimate given in [31, Theorem 3.2].
  • For smooth initial data $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$, we establish the optimal error estimate
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q}(\mathcal{O}))}\leqslant ch^{2},$
    and the nearly optimal pathwise uniform error estimate
    $\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}(\mathcal{O}))}\leqslant ch^{2-\epsilon}\quad\forall\epsilon>0,$
    where $p,q\in(2,\infty)$ and $c>0$ is a generic constant independent of the spatial mesh size $h$. The analysis combines classical finite element theory [1, 31] with discrete maximal $L^{p}$-regularity [12] and discrete stochastic maximal $L^{p}$-regularity [19].
To the best of our knowledge, this work provides the first result on the spatial convergence of numerical approximations for the three-dimensional stochastic Allen–Cahn equation with multiplicative noise, established within the general framework of spatial $L^{q}$-norms. Although we assume deterministic initial data, the analysis can be naturally extended to the case of random initial conditions. The main contribution of this study lies in developing a rigorous and general numerical framework for spatial semidiscretizations of nonlinear stochastic parabolic equations under the framework of general spatial $L^{q}$-norms, thereby distinguishing it from previous works.
The remainder of this paper is structured as follows. Section 2 introduces notation, defines the operator $F$, presents the mild solution to the model problem Eq. 1.1, and describes its spatial semidiscretization. Section 3 states the main results of this work, analyzing the regularity properties of the mild solution under non-smooth initial data, and establishes the convergence of the spatial semidiscretization for both non-smooth and smooth initial conditions. Section 4 provides the rigorous proofs of these results. Finally, Section 5 concludes the paper with a summary of the key findings.
2 Preliminaries
Notational Conventions. For Banach spaces $E_{1}$ and $E_{2}$, we denote by $\mathcal{L}(E_{1},E_{2})$ the space of all bounded linear operators from $E_{1}$ to $E_{2}$; in the case $E_{1}=E_{2}$, we write $\mathcal{L}(E_{1})$ for brevity. The identity operator on any given space is denoted by $I$. For $\theta\in(0,1)$ and $p\in(1,\infty)$, the notation $(\cdot,\cdot)_{\theta,p}$ refers to the real interpolation space defined via the $K$-method, as presented in Chapter 1 of [24].
Let $\mathcal{O}\subset\mathbb{R}^{3}$ denote a bounded convex domain with $\mathcal{C}^{2}$-regular boundary $\partial\mathcal{O}$, as defined in [37, p. 39]. For $q\in[1,\infty]$, we adopt the abbreviation $L^{q}:=L^{q}(\mathcal{O})$ for Lebesgue spaces and denote by $W_{0}^{1,q}(\mathcal{O})$ and $W^{2,q}(\mathcal{O})$ the standard Sobolev spaces, following the conventions established in [14, Chapter 7]. The Laplace operator $\Delta$ is equipped with homogeneous Dirichlet boundary conditions. For $\alpha\geqslant 0$ and $q\in(1,\infty)$, define $\dot{H}^{\alpha,q}$ as the domain of the operator $(-\Delta)^{\alpha/2}$ in $L^{q}$, endowed with the natural norm:
$\|v\|_{\dot{H}^{\alpha,q}}:=\left\|(-\Delta)^{\alpha/2}v\right\|_{L^{q}},\quad v\in\dot{H}^{\alpha,q}.$
The dual space of $\dot{H}^{\alpha,q}$ is denoted by $\dot{H}^{-\alpha,q^{\prime}}$, where $q^{\prime}$ is the Hölder conjugate of $q$. When $\alpha=2$, $\dot{H}^{2,q}$ coincides with $W_{0}^{1,q}(\mathcal{O})\cap W^{2,q}(\mathcal{O})$, and its norm is equivalent to the $W^{2,q}(\mathcal{O})$-norm (cf. [14, Theorem 9.15]). For $\alpha=1$, $\dot{H}^{1,q}$ agrees with $W_{0}^{1,q}(\mathcal{O})$, with equivalent norms (see [37, Theorem 16.14]).
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a complete probability space endowed with a filtration $\mathbb{F}:=(\mathcal{F}_{t})_{t\geqslant 0}$ satisfying the usual conditions. On this stochastic basis, we define a sequence $(W_{n})_{n\in\mathbb{N}}$ of independent real-valued $\mathbb{F}$-adapted Brownian motions. The associated expectation operator is denoted by $\mathbb{E}$. Let $H$ be a separable Hilbert space with inner product $(\cdot,\cdot)_{H}$ and an orthonormal basis $(h_{n})_{n\in\mathbb{N}}$. The $H$-cylindrical Brownian motion $W_{H}$ is defined such that, for each $t\in\mathbb{R}_{+}$, $W_{H}(t)\in\mathcal{L}(H,L^{2}(\Omega))$ is given by
$W_{H}(t)x:=\sum_{n\in\mathbb{N}}W_{n}(t)(x,h_{n})_{H},\quad\forall x\in H.$
We refer the reader to [34] for the stochastic integration with respect to $W_{H}$.
Let $p\in(1,\infty)$ and let $E$ be a Banach space. Denote by $L^{p}_{\mathbb{F}}(\Omega\times(0,T);E)$ the closure in the Bochner space $L^{p}(\Omega\times(0,T);E)$ of the linear span of the set
[…]
where $\mathbbm{1}_{(s,t]\times F}$ denotes the indicator function of the set $(s,t]\times F$. Let $L^{0}_{\mathbb{F}}(\Omega;C([0,T];E))$ denote the space of all continuous, $\mathbb{F}$-adapted, $E$-valued processes. Two processes $g_{1}$ and $g_{2}$ in $L_{\mathbb{F}}^{0}(\Omega;C([0,T];E))$ are considered identical if $g_{1}=g_{2}$ in $C([0,T];E)$ holds $\mathbb{P}$-almost surely. We denote by $L^{p}_{\mathbb{F}}(\Omega;C([0,T];E))$ the subspace of all processes $g\in L^{0}_{\mathbb{F}}(\Omega;C([0,T];E))$ satisfying
$\mathbb{E}\big{[}\|g\|_{C([0,T];E)}^{p}\big{]}<\infty,$
where $C([0,T];E)$ is the space of continuous $E$-valued functions on $[0,T]$, equipped with the supremum norm. The space $\gamma(H,E)$ denotes the space of $\gamma$-radonifying operators from the Hilbert space $H$ to $E$ (see, e.g., [16, Chapter 9]). If $E$ is itself a Hilbert space, then $\gamma(H,E)$ is isometrically isomorphic to the space of Hilbert–Schmidt operators from $H$ to $E$.
Definition of the Operator $F$. Let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of real-valued, continuously differentiable functions defined on $\overline{\mathcal{O}}\times\mathbb{R}$, satisfying the following conditions:
$\sum_{n\in\mathbb{N}}|f_{n}(x,0)|^{2}=0,$ $\forall x\in\partial\mathcal{O},$ (2.1)
$\sum_{n\in\mathbb{N}}\big{(}|f_{n}(x,y)|^{2}+|\nabla_{x}f_{n}(x,y)|^{2}\big{)}\leqslant C_{F}(1+|y|^{2}),$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.2)
$\sum_{n\in\mathbb{N}}|\nabla_{y}f_{n}(x,y)|^{2}\leqslant C_{F},$ $\forall x\in\mathcal{O},\,\forall y\in\mathbb{R},$ (2.3)
where $C_{F}>0$ is a constant independent of $x$ and $y$. Here and in the sequel, we slightly abuse notation by identifying each $f_{n}$ with the corresponding Nemytskii operator defined pointwise by
$f_{n}(v):=f_{n}(\cdot,v(\cdot)),\quad\text{for all }v\in L^{2}.$
Recall that $(h_{n})_{n\in\mathbb{N}}$ denotes a fixed orthonormal basis of the separable Hilbert space $H$. Given condition Eq. 2.2, for each $v\in L^{2}$, define the operator $F(v)\in\gamma(H,L^{2})$ by
$F(v):=\sum_{n\in\mathbb{N}}h_{n}\otimes f_{n}(v),$ (2.4)
where the elementary tensor $h_{n}\otimes f_{n}(v)\in\gamma(H,L^{2})$ acts on $x\in H$ via
$(h_{n}\otimes f_{n}(v))(x):=(x,h_{n})_{H}f_{n}(v).$
In this setting, the model equation (1.1) takes the explicit form of the stochastic evolution equation:
$\begin{cases}\mathrm{d}y(t)=\left(\Delta y+y-y^{3}\right)(t)\,\mathrm{d}t+\displaystyle\sum_{n\in\mathbb{N}}f_{n}(y(t))\,\mathrm{d}W_{n}(t),\quad t\in[0,T],\\ y(0)=y_{0},\end{cases}$
where $\Delta$ denotes the Laplace operator with homogeneous Dirichlet boundary conditions.
Mild Solutions. We define the two convolution operators:
$(S_{0}g)(t)$ $:=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}s,\quad 0\leqslant t\leqslant T,$ (2.5)
$(S_{1}g)(t)$ $:=\int_{0}^{t}e^{(t-s)\Delta}g(s)\,\mathrm{d}W_{H}(s),\quad 0\leqslant t\leqslant T,$ (2.6)
whenever the integrals are well-defined. Here, $(e^{t\Delta})_{t\geqslant 0}$ denotes the analytic semigroup generated by the Laplace operator $\Delta$ under homogeneous Dirichlet boundary conditions.
Assume that $y_{0}:(\Omega,\mathcal{F}_{0},\mathbb{P})\to L^{4}$ is strongly measurable. A process $y\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$ is called a mild solution to equation (1.1) if it satisfies, $\mathbb{P}$-almost surely for all $t\in[0,T]$,
$y(t)=G(t)+\big{[}S_{0}(y-y^{3})\big{]}(t)+\big{[}S_{1}F(y)\big{]}(t),$ (2.7)
where
$G(t):=e^{t\Delta}y_{0},\quad t\in[0,T].$ (2.8)
Spatial Semidiscretization. Let $\mathcal{K}_{h}$ be a conforming and quasi-uniform triangulation of the domain $\mathcal{O}$, consisting of three-dimensional simplices. We define the finite element space $X_{h}$ as
$X_{h}:=\Bigl{\{}v_{h}\in C(\overline{\mathcal{O}})\Bigm{|}v_{h}=0\text{ on }\overline{\mathcal{O}}\setminus\mathcal{O}_{h},\text{ and }v_{h}\text{ is piecewise linear with respect to each }K\in\mathcal{K}_{h}\Bigr{\}},$
where $C(\overline{\mathcal{O}})$ denotes the space of continuous functions on $\overline{\mathcal{O}}$, and $\mathcal{O}_{h}$ is the closure of the union of all elements in $\mathcal{K}_{h}$. We assume that every boundary vertex of $\mathcal{O}_{h}$ lies on $\partial\mathcal{O}$. The spatial mesh size $h$ is defined as the maximum diameter over all elements in $\mathcal{K}_{h}$. Let $P_{h}$ denote the $L^{2}$-orthogonal projection onto $X_{h}$, and define the discrete Laplace operator $\Delta_{h}:X_{h}\to X_{h}$ via the identity
$\int_{\mathcal{O}}(\Delta_{h}u_{h})v_{h}\,\mathrm{d}x=-\int_{\mathcal{O}}\nabla u_{h}\cdot\nabla v_{h}\,\mathrm{d}x,\quad\text{for all }u_{h},v_{h}\in X_{h}.$
We consider the following spatial semidiscretization of the model equation:
$\mathrm{d}y_{h}(t)=\left(\Delta_{h}y_{h}+y_{h}-P_{h}y_{h}^{3}\right)(t)\,\mathrm{d}t+P_{h}F(y_{h}(t))\,\mathrm{d}W_{H}(t),\quad t\in[0,T],$ (2.9a)
$y_{h}(0)=P_{h}y_{0}.$ (2.9b)
Suppose that $y_{0}:(\Omega,\mathcal{F}_{0},\mathbb{P})\to L^{2}(\mathcal{O})$ is strongly measurable. A process $y_{h}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];X_{h}))$, where $X_{h}$ is equipped with the $L^{2}$-norm, is called a strong solution to (2.9) if, $\mathbb{P}$-almost surely, the identity
$y_{h}(t)=y_{h}(0)+\int_{0}^{t}\left(\Delta_{h}y_{h}+y_{h}-P_{h}y_{h}^{3}\right)(s)\,\mathrm{d}s+\int_{0}^{t}P_{h}F(y_{h}(s))\,\mathrm{d}W_{H}(s)$
holds for all $t\in[0,T]$. By the standard stability estimate for $P_{h}$ (see, e.g., Eq. 4.19), together with the growth and Lipschitz conditions on $F$ given in Lemma 4.5, the classical theory of stochastic differential equations (see, e.g., [27, Theorem 3.21]) ensures the existence and uniqueness of a strong solution $y_{h}$ to the spatial semidiscretization (2.9).
3 Main results
In the remainder of this paper, we denote by $c$ a generic positive constant that is independent of the spatial mesh size $h$. It may, however, depend on the final time $T$, the domain $\mathcal{O}$, the constant $C_{F}$ from Eqs. 2.2 and 2.3, the indices of the relevant Lebesgue, Sobolev, and interpolation spaces, the initial data $y_{0}$, the shape-regularity parameters of the triangulation $\mathcal{K}_{h}$, and the exponents associated with $h$, unless otherwise specified. The value of $c$ may differ at each occurrence.
Theorem 3.1.
Assume that $y_{0}\in L^{\infty}$ is deterministic. Then equation Eq. 1.1 admits a unique mild solution $y$, which satisfies that
$y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\qquad\forall p,q\in(2,\infty).$ (3.1)
Furthermore, for any $p,q\in(2,\infty)$ and $\epsilon>0$, the solution $y$ possesses the following regularity properties:
$y\in L_{\mathbb{F}}^{p}\big{(}\Omega;C([0,T];L^{q})\big{)},$ (3.2)
$y\in L_{\mathbb{F}}^{p}\big{(}\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}\big{)},$ (3.3)
$y-G\in L_{\mathbb{F}}^{p}\big{(}\Omega;C([0,T];\dot{H}^{1-\epsilon,q})\big{)},$ (3.4)
$y-G\in L_{\mathbb{F}}^{p}\big{(}\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q}\big{)},$ (3.5)
where $G$ is defined by Eq. 2.8.
Theorem 3.2.
Assume that $y_{0}\in L^{\infty}$ is deterministic. Let $y$ be the mild solution of Eq. 1.1, and let $y_{h}$ be the strong solution to the spatial semidiscretization Eq. 2.9. Then, for any $p\in(2,\infty)$, the following error estimate holds for all $t\in(0,T]$:
$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant c\left(h^{2/p-\epsilon}+h^{2/p}t^{-1/p}\right),\quad\forall\epsilon>0.$ (3.6)
Theorem 3.3.
Assume that $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ is deterministic. Let $y$ be the mild solution of Eq. 1.1, and let $y_{h}$ be the strong solution of Eq. 2.9. Then the following error estimates hold:
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p,q\in(2,\infty),$ (3.7)
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}))}\leqslant ch^{2-\epsilon},\quad\forall p\in(2,\infty),\,\forall\epsilon>0.$ (3.8)
It is worth noting that the aforementioned theoretical results remain valid under the assumption that the operator $F$ possesses the growth and Lipschitz continuity properties outlined in Lemma 4.5.
Remark 3.1.
Liu [20] proposed a theoretical framework for a broad spectrum of nonlinear stochastic partial differential equations. Applying this framework to the model problem Eq. 1.1 yields
$y\in L^{p}(\Omega;C([0,T];\dot{H}^{1,2})),\quad\forall p\in(2,\infty),$
provided that $y_{0}\in L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};\dot{H}^{1,2})$, where $p\in[2,\infty)$. Furthermore, Liu and Qiao [23] have established certain regularity estimates in the context of the spatial $L^{2}$-norm, and some regularity estimates with the spatial $L^{\infty}$-norm were derived by using Sobolev’s embedding theorem. Regarding the stochastic Allen-Cahn equation with additive noise, Bréhier et al. [5, Lemma 3.1] provided some a priori estimates involving general spatial $L^{q}$ norms.
Remark 3.2.
In the numerical analysis of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, the existing literature often imposes strict regularity conditions on the diffusion coefficients. For instance, the study in [25] demands that the diffusion coefficients possess bounded second-order derivatives. Our choice of diffusion coefficient $F$ adheres to the criteria outlined in Assumption 2.2 of [23]; however, it does not meet the more stringent requirements stated in Assumption 2.3 of the same reference. Investigating the numerical approximations of Eq. 1.1 under less restrictive conditions on $F$, especially the loosening of condition (2.1), has important theoretical and practical implications.
4 Proofs
The objective of this section is to affirm the validity of Theorems 3.1, 3.2, and 3.3. The structure of this section is as follows:
  • Subsection 4.1 introduces key lemmas that serve as foundational results.
  • Subsection 4.2 focuses on the proof of Theorem 3.1, which establishes the regularity of the mild solution of Eq. 1.1.
  • Subsection 4.3 is dedicated to the proofs of Theorems 3.2 and 3.3, which concern the error estimates of the spatial semidiscretization.
4.1 Preliminary Lemmas
It is a well-established result that for any $\alpha\in(0,2)$ and $q\in(1,\infty)$, the space $\dot{H}^{\alpha,q}$ coincides with the complex interpolation space $[L^{q},\dot{H}^{2,q}]_{\alpha/2}$, equipped with equivalent norms. This result is documented, for example, in Chapter 16 of [37]. The combination of Propositions 1.3 and 1.4 and Theorem 4.17 from [24] and Theorem C.4.1 in [15] yields the following continuous embeddings.
Lemma 4.1.
Let $0\leqslant\alpha<\beta\leqslant 2$ and $p,q\in(1,\infty)$. Then the following continuous embeddings hold:
  • (i)
    If $\gamma>(1-\theta)\alpha+\theta\beta$ for some $\theta\in(0,1)$, then $\dot{H}^{\gamma,q}$ is continuously embedded in the real interpolation space $(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})_{\theta,p}$.
  • (ii)
    If $\gamma<(1-\theta)\alpha+\theta\beta$ for some $\theta\in(0,1)$, then the real interpolation space $(\dot{H}^{\alpha,q},\dot{H}^{\beta,q})_{\theta,p}$ is continuously embedded in $\dot{H}^{\gamma,q}$.
Theorem 2.12 and inequality (2.128) in [37] provide the following standard estimate for the growth of the analytic semigroup generated by the Laplace operator with homogeneous Dirichlet boundary conditions.
Lemma 4.2.
Let $\alpha\in[0,\infty)$ and $q\in(1,\infty)$. For all $t>0$, the following estimate holds:
$\|(-\Delta)^{\alpha}e^{t\Delta}\|_{\mathcal{L}(L^{q})}+t\Big{\|}(-\Delta)^{\alpha}\frac{\mathrm{d}}{\mathrm{d}t}e^{t\Delta}\Big{\|}_{\mathcal{L}(L^{q})}\leqslant ct^{-\alpha}.$
For the convolution operator $S_{0}$, defined by Eq. 2.5, we have the following celebrated maximal $L^{p}$-regularity estimate; see Theorem 4.2 in [36] and Corollary 1.14 from [24].
Lemma 4.3.
Let $p,q\in(1,\infty)$. Then for any $g\in L^{p}(0,T;L^{q})$, the following inequality holds:
$\big{\lVert}\frac{\mathrm{d}}{\mathrm{d}t}S_{0}g\big{\rVert}_{L^{p}(0,T;L^{q})}+\lVert{S_{0}g}\rVert_{L^{p}(0,T;\dot{H}^{2,q})}+\lVert{S_{0}g}\rVert_{C([0,T];(L^{q},\dot{H}^{2,q})_{1-1/p,p})}\leqslant c\lVert{g}\rVert_{L^{p}(0,T;L^{q})}.$
For the stochastic convolution operator $S_{1}$, defined by Eq. 2.6, we present the following fundamental stochastic maximal $L^{p}$-regularity estimate; see Theorems 1.1 and 1.2 in [33].
Lemma 4.4.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. Then for any $g\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$, the following inequality holds:
$\lVert{S_{1}g}\rVert_{L^{p}(\Omega;C([0,T];(L^{q},\dot{H}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1}g}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1,q})}\leqslant c\lVert{g}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{q}))}.$
The operator $F$, defined by Eq. 2.4, possesses the following growth and Lipschitz continuity properties.
Lemma 4.5.
Let $q\in[2,\infty)$. The following inequalities are valid:
$\lVert{F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\left(1+\lVert{v}\rVert_{L^{q}}\right),$ for all $\text{for all }v\in L^{q};$ (4.1)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{1,q})}\leqslant c\left(1+\lVert{v}\rVert_{\dot{H}^{1,q}}\right),$ for all $\text{for all }v\in\dot{H}^{1,q};$ (4.2)
$\lVert{F(v)}\rVert_{\gamma(H,\dot{H}^{s_{1},q})}\leqslant c\left(1+\lVert{v}\rVert_{\dot{H}^{s_{2},q}}\right),$ for $\text{for all }v\in\dot{H}^{s_{2},q},\text{ where }0\leqslant s_{1}<s_{2}<1;$ (4.3)
$\lVert{F(u)-F(v)}\rVert_{\gamma(H,L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},$ for all $\text{for all }u,v\in L^{q}.$ (4.4)
Proof.
We focus our demonstration on Eq. 4.3; the proofs of Eq. 4.1, Eq. 4.2, and Eq. 4.4 follow similar, yet more straightforward, lines of reasoning. For the definition of $\gamma$-radonifying operators, we direct the reader to Section 9.1 of [16]. Consider $\Omega_{\gamma}$ as a probability space supporting a sequence $(\gamma_{n})_{n\in\mathbb{N}}$ of independent standard Gaussian random variables. Let $\mathbb{E}_{\gamma}$ denote the expectation operator with respect to the probability space $\Omega_{\gamma}$. The subsequent argument is structured in two main steps.
Step 1. We proceed to demonstrate that for any $\theta\in(0,1)$ and $N\in\mathbb{N}_{>0}$, the following inequality holds:
$\lVert{A_{N}v}\rVert_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}\big{)},\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q},$ (4.5)
where
$A_{N}v:=\sum_{n=0}^{N}\gamma_{n}f_{n}(v),\quad\forall v\in L^{q}.$
For any $v\in\dot{H}^{1,q}$, we deduce that
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}$ $\leqslant c\bigg{[}\mathbb{E}_{\gamma}\int_{\mathcal{O}}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\nabla f_{n}(v)}\Big{\rvert}^{q}\,\mathrm{d}x\bigg{]}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}\mathbb{E}_{\gamma}\Big{\lvert}{\sum_{n=0}^{N}\gamma_{n}\nabla f_{n}(v)}\Big{\rvert}^{2}\Big{)}^{q/2}\,\mathrm{d}x\bigg{]}^{1/q}$
$=c\bigg{[}\int_{\mathcal{O}}\Big{(}\sum_{n=0}^{N}\lvert{\nabla f_{n}(v)}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}x\bigg{]}^{1/q}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}c\bigg{[}\int_{\mathcal{O}}\Big{(}1+\lvert{v}\rvert^{2}+\lvert{\nabla v}\rvert^{2}\Big{)}^{q/2}\,\mathrm{d}x\bigg{]}^{1/q},$
where step $(\mathrm{i})$ uses the famous Kahane-Khintchine inequality (see Theorem 6.2.6 from [16]), and step $(\mathrm{ii})$ relies on conditions Eqs. 2.3 and 2.2. Using the standard inequality
$\left(\int_{\mathcal{O}}|v|^{q}+|\nabla v|^{q}\,\mathrm{d}x\right)^{1/q}\leqslant c\lVert{v}\rVert_{\dot{H}^{1,q}},$
we further deduce that
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};\dot{H}^{1,q})}\leqslant c(1+\lVert{v}\rVert_{\dot{H}^{1,q}}).$
Analogously, employing the condition Eq. 2.2 yields
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c(1+\lVert{v}\rVert_{L^{q}}),\quad\forall v\in L^{q},$
and utilizing the condition Eq. 2.3 results in
$\lVert{A_{N}u-A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};L^{q})}\leqslant c\lVert{u-v}\rVert_{L^{q}},\quad\forall u,v\in L^{q}.$
Thus, following the methodology outlined in Lemma 28.1 of [30], one obtains
$\lVert{A_{N}v}\rVert_{(L^{q}(\Omega_{\gamma};L^{q}),L^{q}(\Omega_{\gamma};\dot{H}^{1,q}))_{\theta,q}}\leqslant c(1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}),\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q}.$
Combining this result with the continuous embedding property
$\left(L^{q}(\Omega_{\gamma};L^{q}),L^{q}(\Omega_{\gamma};\dot{H}^{1,q})\right)_{\theta,q}\hookrightarrow L^{q}\left(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q}\right),$
which can be easily verified by definition, yields
$\lVert{A_{N}v}\rVert_{L^{q}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c\big{(}1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}\big{)},\quad\forall v\in(L^{q},\dot{H}^{1,q})_{\theta,q}.$
Applying the Kahane-Khintchine inequality once again, we arrive at the assertion in Eq. 4.5.
Step 2. For any $\theta\in(0,1)$ and $v\in(L^{q},\dot{H}^{1,q})_{\theta,q}$, we consider the equality
$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}=\Big{\lVert}\lim_{N\to\infty}A_{N}v\Big{\rVert}_{L^{2}(\Omega_{\gamma};(L^{q},\dot{H}^{1,q})_{\theta,q})},$
which follows directly from the definition of $F$ as stated in (2.4) and is justified by Theorem 9.1.19 in [16]. Using the inequality (4.5), we apply Fatou’s Lemma to deduce that
$\lVert{F(v)}\rVert_{\gamma(H,(L^{q},\dot{H}^{1,q})_{\theta,q})}\leqslant c(1+\lVert{v}\rVert_{(L^{q},\dot{H}^{1,q})_{\theta,q}}).$
Consequently, the desired inequality (4.3) is established through the application of Lemma 4.1. This concludes the proof. ∎
Remark 4.1.
As established by Theorem 4.36 in [24], the interpolation space $(L^{2},\dot{H}^{1,2})_{s,2}$ coincides with $\dot{H}^{s,2}$, sharing equivalent norms for all $s\in(0,1)$. Consequently, we refine the inequality Eq. 4.3 to the following form:
$\|F(v)\|_{\gamma(H,\dot{H}^{s,2})}\leqslant c(1+\|v\|_{\dot{H}^{s,2}}),\quad\text{for all $v\in\dot{H}^{s,2}$, where $s\in(0,1)$}.$
It is well known that the space $L^{2}$ admits an orthonormal basis $(\phi_{n})_{n\in\mathbb{N}}$, consisting of eigenvectors of the negative Laplace operator $-\Delta$ with homogeneous Dirichlet boundary conditions. This basis $(\phi_{n})_{n\in\mathbb{N}}$ forms a complete orthogonal system in the space $\dot{H}^{s,2}$ for all $s\in(0,\infty)$.
Lemma 4.6.
For each $n\in\mathbb{N}$ and $y_{0}^{(n)}\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}$, which is deterministic, the stochastic Allen-Cahn equation
$\mathrm{d}y^{(n)}(t)=\left[\Delta y^{(n)}+y^{(n)}-(y^{(n)})^{3}\right](t)\,\mathrm{d}t+F(y^{(n)}(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T,$ (4.6a)
$y^{(n)}(0)=y_{0}^{(n)}$ (4.6b)
has a unique mild solution. Moreover, for all $p,q\in(2,\infty)$, the following inequality holds:
$\lVert{y^{(n)}}\rVert_{L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})}\leqslant c(1+\lVert{y_{0}^{(n)}}\rVert_{L^{q}}).$ (4.7)
Proof.
We split the proof into the following two steps.
Step 1. Define $G^{(n)}(t):=e^{t\Delta}y_{0}^{(n)}$ for $t\in[0,T]$. Given that $y_{0}^{(n)}\in\dot{H}^{2,q}$ for every $q\in(2,\infty)$, it follows from Lemma 4.2 with $\alpha=\beta=2$ that
$G^{(n)}\in L^{p}(0,T;\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.8)
In view of the estimates for $F$ in Eqs. 4.2 and 4.4 and the fact that $y_{0}^{(n)}\in\dot{H}^{1,2}$ is deterministic, we use [20, Theorem 1.1] (see also, [23, Lemma 2.3 and Theorem 3.1]) to deduce that Eq. 4.6 has a unique mild solution $y^{(n)}$ and
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,2}),\quad\forall p\in(2,\infty).$ (4.9)
The continuous embedding of $\dot{H}^{1,2}$ in $L^{6}$ then leads to
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{6}),\quad\forall p\in(2,\infty),$
which in turn implies that $y^{(n)}-(y^{(n)})^{3}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{2})$ for all $p\in(2,\infty)$. Applying Lemma 4.3, we obtain
$S_{0}(y^{(n)}-(y^{(n)})^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.10)
Since (4.2) and (4.9) ensure that $F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}^{1,2}))$ for every $p\in(2,\infty)$, Lemma 4.4 allows us to conclude that
$S_{1}F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$ (4.11)
Considering the mild solution representation
$y^{(n)}=G^{(n)}+S_{0}(y^{(n)}-(y^{(n)})^{3})+S_{1}F(y^{(n)}),$
and combining (4.8), (4.10), and (4.11), we deduce that
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,2}),\quad\forall p\in(2,\infty).$
The continuous embeddings of $\dot{H}^{2,2}$ into $\dot{H}^{1,6}$ and $\dot{H}^{2,2}$ into $L^{q}$ for all $q\in(2,\infty)$ yield
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),\quad\forall p,q\in(2,\infty),$
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,6}),\quad\forall p\in(2,\infty).$
Hence, Eqs. 4.10 and 4.11 can be refined to
$S_{0}(y^{(n)}-(y^{(n)})^{3})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$ (4.12)
$S_{1}F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty).$
Combining these refined regularity results with (4.8), we arrive at
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,6}),\quad\forall p\in(2,\infty),$
which, through the continuous embedding of $\dot{H}^{2,6}$ into $\dot{H}^{1,q}$ for all $q\in(2,\infty)$, implies that
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q}),\quad\forall p,q\in(2,\infty).$
This regularity result allows us to further improve (4.11) to
$S_{1}F(y^{(n)})\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty),$
which, in conjunction with (4.8) and (4.12), yields
$y^{(n)}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),\quad\forall p,q\in(2,\infty).$ (4.13)
Step 2. Let $p,q\in(2,\infty)$ be arbitrary but fixed. Considering the regularity assertion Eq. 4.13 and the fact, $\mathbb{P}$-almost surely,
$y^{(n)}(t)=y_{0}^{(n)}+\int_{0}^{t}\Delta y^{(n)}(s)+y^{(n)}(s)-(y^{(n)}(s))^{3}\,\mathrm{d}s+\int_{0}^{t}F(y^{(n)}(s))\,\mathrm{d}W_{H}(s),\quad t\in[0,T],$
we use Itô’s formula in UMD Banach spaces (see [8, Theorem 2.4]) to conclude that, for any $t\in(0,T]$,
$\mathbb{E}\Big{[}\lVert{y^{(n)}(t)}\rVert_{L^{q}}^{p}-\lVert{y_{0}^{(n)}}\rVert_{L^{q}}^{p}\Big{]}$
[…]
$\quad{}+\frac{p(q-1)}{2}\,\mathbb{E}\Biggl{[}\int_{0}^{t}\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p-q}\,\int_{\mathcal{O}}\lvert{y^{(n)}(s)}\rvert^{q-2}\sum_{k=0}^{\infty}\lvert{f_{k}(y^{(n)}(s))}\rvert^{2}\,\mathrm{d}x\,\mathrm{d}s\Biggr{]}$
$\quad{}+\frac{p(p-q)}{2}\,\mathbb{E}\Biggl{\{}\int_{0}^{t}\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p-2q}\sum_{k=0}^{\infty}\bigg{[}\int_{\mathcal{O}}\lvert{y^{(n)}(s)}\rvert^{q-2}y^{(n)}(s)f_{k}(y^{(n)}(s))\,\mathrm{d}x\bigg{]}^{2}\,\mathrm{d}s\Biggr{\}}$
[…]
Considering the inequality
$\int_{\mathcal{O}}\lvert{y^{(n)}(s)}\rvert^{q-2}y^{(n)}(s)[\Delta y^{(n)}(s)-(y^{(n)}(s))^{3}]\,\mathrm{d}x$
[…]
we infer the bound for $I_{1}$:
$I_{1}\leqslant p\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
For $I_{2}+I_{3}$, using Hölder’s inequality in conjunction with Young’s inequality, and employing condition Eq. 2.2, we derive the following bound:
$I_{2}+I_{3}\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]}.$
Combining the above bounds for $I_{1}$ and $I_{2}+I_{3}$, we arrive at
$\mathbb{E}\lVert{y^{(n)}(t)}\rVert_{L^{q}}^{p}\leqslant\lVert{y_{0}^{(n)}}\rVert_{L^{q}}^{p}+c\,\mathbb{E}\bigg{[}\int_{0}^{t}1+\lVert{y^{(n)}(s)}\rVert_{L^{q}}^{p}\,\mathrm{d}s\bigg{]},\quad\forall t\in[0,T].$
An application of Gronwall’s inequality then yields
$\lVert{y^{(n)}}\rVert_{C([0,T];L^{p}(\Omega;L^{q}))}\leqslant c\big{(}1+\lVert{y_{0}^{(n)}}\rVert_{L^{q}}\big{)},$
which implies the desired estimate Eq. 4.7. This completes the proof. ∎
4.2 Proof of Theorem 3.1
We split the proof into the following three steps.
Step 1. We proceed to show that equation Eq. 1.1 has a unique mild solution $y$, and that Eq. 3.1 is satisfied. Let $p\in(2,\infty)$ and $q\in[4,\infty)$ be arbitrarily chosen. For each $n\in\mathbb{N}$, denote by $y^{(n)}$ the mild solution of equation Eq. 4.6, where $y_{0}^{(n)}$ is defined as
$y_{0}^{(n)}:=\mathop{\mathrm{argmin}}_{w\in\mathrm{span}\{\phi_{0},\ldots,\phi_{n}\}}\lVert{y_{0}-w}\rVert_{L^{3q}}.$
For any $m,n\in\mathbb{N}$, put $e_{m,n}:=y^{(m)}-y^{(n)}$. By definition, we infer that
$\mathrm{d}e_{m,n}(t)=(\Delta e_{m,n}+e_{m,n}+(y^{(n)})^{3}-(y^{(m)})^{3})(t)\,\mathrm{d}t+\big{(}F(y^{(m)}(t))-F(y^{(n)}(t))\big{)}\,\mathrm{d}W_{H}(t),\quad t\in[0,T].$
Observing that the integral over $\mathcal{O}$ of
$\lvert{e_{m,n}(t)}\rvert^{3q-2}e_{m,n}(t)[\Delta e_{m,n}(t)+(y^{(n)}(t))^{3}-(y^{(m)}(t))^{3}]$
is non-positive for all $t\in[0,T]$, and following the proof of Eq. 4.7, we can deduce that
$\lVert{e_{m,n}}\rVert_{L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})}\leqslant c\lVert{y_{0}^{(m)}-y_{0}^{(n)}}\rVert_{L^{3q}}.$ (4.14)
On the other hand, from the fact that $\dot{H}^{2,2}$ is dense in $L^{3q}$ and that $(\phi_{n})_{n\in\mathbb{N}}$ is a complete orthogonal system in $\dot{H}^{2,2}$, we conclude that $(y_{0}^{(n)})_{n\in\mathbb{N}}$ is a Cauchy sequence in $L^{3q}$ and converges to $y_{0}$ in $L^{3q}$, i.e.,
$\lim_{n\to\infty}\lVert{y_{0}^{(n)}-y_{0}}\rVert_{L^{3q}}=0.$ (4.15)
Consequently, there exists a unique process $y\in L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q})$ such that
$\lim_{n\to\infty}y^{(n)}=y\quad\text{ in }L_{\mathbb{F}}^{3p}(\Omega\times(0,T);L^{3q}).$ (4.16)
For any $n\in\mathbb{N}$, applying Hölder’s inequality yields
$\lVert{(y^{(n)})^{3}-y^{3}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}$
[…]
Therefore, using the convergence established in Eq. 4.16, we deduce that
$\lim_{n\to\infty}(y^{(n)})^{3}=y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}),$
which consequently leads to
$\lim_{n\to\infty}y^{(n)}-(y^{(n)})^{3}=y-y^{3}\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q}).$
Using Lemma 4.3, this directly implies the following convergence of $S_{0}(y^{(n)}-(y^{(n)})^{3})$:
$\lim_{n\to\infty}S_{0}(y^{(n)}-(y^{(n)})^{3})=S_{0}(y-y^{3})\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})\cap L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q})).$
Furthermore, invoking Eqs. 4.4 and 4.16 yields
$\lim_{n\to\infty}F(y^{(n)})=F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q})),$
and subsequently by Lemma 4.4 we infer the following convergence for $S_{1}F(y^{(n)})$:
$\lim_{n\to\infty}S_{1}F(y^{(n)})=S_{1}F(y)\quad\text{ in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})\cap L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q})).$
Additionally, applying Lemma 4.2 with $\alpha=\beta=0$ and using Eq. 4.15, we establish the following convergence for $G^{(n)}$:
$\lim_{n\to\infty}G^{(n)}=G\quad\text{ in }C([0,T];L^{q}),$
where $G$ is defined by Eq. 2.8, and $G^{(n)}(t):=e^{t\Delta}y_{0}^{(n)}$ for each $t\in[0,T]$. Finally, combining the derived convergences of $y^{(n)}$, $G^{(n)}$, $S_{0}(y^{(n)}-(y^{(n)})^{3})$, and $S_{1}F(y^{(n)})$, and recalling that $y^{(n)}=G^{(n)}+S_{0}(y^{(n)}-(y^{(n)})^{3})+S_{1}F(y^{(n)})$, we obtain by passing to the limit $n\to\infty$ that
$y=G+S_{0}(y-y^{3})+S_{1}F(y)\quad\text{in }L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})\cap L_{\mathbb{F}}^{p}(\Omega;C([0,T];L^{q})).$
This confirms that $y$ is a mild solution of Eq. 1.1, and it is evident that Eq. 3.1 holds.
Step 2. We now consider the uniqueness of the mild solution $y$, which can be proved by a standard localization argument. Suppose $\widetilde{y}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$ is another mild solution of Eq. 1.1, and define the difference $e:=y-\widetilde{y}$. Then $e$ satisfies the stochastic evolution equation
$\begin{cases}\mathrm{d}e(t)=\left(\Delta e(t)+e(t)+\widetilde{y}^{3}(t)-y^{3}(t)\right)\mathrm{d}t\\ \qquad\qquad{}+\left(F(y(t))-F(\widetilde{y}(t))\right)\mathrm{d}W_{H}(t)\quad\text{in $\dot{H}^{-1,2}$,}\quad t\in[0,T],\\ e(0)=0.\end{cases}$
For $m\geqslant 1$, define the stopping time
$\sigma_{m}:=\inf\left\{t\in[0,T]\colon\max\left(\|y(t)\|_{L^{4}},\|\widetilde{y}(t)\|_{L^{4}}\right)\geqslant m\right\},$
with $\sigma_{m}=T$ if the set is empty. Set $e_{m}(t):=e(t\wedge\sigma_{m})$ for $t\in[0,T]$. Then $e_{m}$ satisfies the stopped equation
$\begin{cases}\mathrm{d}e_{m}(t)=\mathbbm{1}_{[0,\sigma_{m}]}(t)\left(\Delta e_{m}(t)+e_{m}(t)+\widetilde{y}^{3}(t\wedge\sigma_{m})-y^{3}(t\wedge\sigma_{m})\right)\mathrm{d}t\\ \qquad\qquad{}+\mathbbm{1}_{[0,\sigma_{m}]}(t)\left(F(y(t))-F(\widetilde{y}(t))\right)\mathrm{d}W_{H}(t)\quad\text{in $\dot{H}^{-1,2}$},\quad t\in[0,T],\\ e_{m}(0)=0.\end{cases}$
Applying Itô’s formula to $\|e_{m}(t)\|_{L^{2}}^{2}$ and taking expectation yields, for all $t\in[0,T]$,
$\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}$ $=2\mathbb{E}\bigg{[}\mathbbm{1}_{[0,\sigma_{m}]}(s)\big{\langle}\Delta e_{m}(s)+e_{m}(s)+\widetilde{y}^{3}(s)-y^{3}(s),\,e_{m}(s)\big{\rangle}\,\mathrm{d}s\bigg{]}$
$\quad+\mathbb{E}\bigg{[}\int_{0}^{t}\mathbbm{1}_{[0,\sigma_{m}]}(s)\|F(y(s))-F(\widetilde{y}(s))\|_{\gamma(H,L^{2})}^{2}\,\mathrm{d}s\bigg{]},$
[…]
$\mathbbm{1}_{[0,\sigma_{m}]}(s)\big{\langle}\Delta e_{m}(s)+e_{m}(s)+\widetilde{y}^{3}(s)-y^{3}(s),\,e_{m}(s)\big{\rangle}\leqslant\mathbbm{1}_{[0,\sigma_{m}]}(s)\|e_{m}(s)\|_{L^{2}}^{2},$
it follows that
$\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}\leqslant\mathbb{E}\bigg{[}\int_{0}^{t}\mathbbm{1}_{[0,\sigma_{m}]}(s)\left(\|e_{m}(s)\|_{L^{2}}^{2}+\|F(y(s))-F(\widetilde{y}(s))\|_{\gamma(H,L^{2})}^{2}\right)\mathrm{d}s\bigg{]}.$
Using the Lipschitz condition Eq. 4.4 of $F$, we deduce
$\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}\leqslant c\mathbb{E}\int_{0}^{t}\|e_{m}(s)\|_{L^{2}}^{2}\,\mathrm{d}s,\quad\forall t\in[0,T].$
Gronwall’s inequality then yields
$\sup_{t\in[0,T]}\mathbb{E}\|e_{m}(t)\|_{L^{2}}^{2}=0,\quad\forall m\geqslant 1.$
Since $e_{m}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$, this implies $\|e_{m}\|_{C([0,T];L^{4})}=0$ holds $\mathbb{P}$-almost surely. Consequently,
$y=\widetilde{y}\quad\text{in }C([0,\sigma_{m}];L^{4})\text{ holds }\mathbb{P}\text{-almost surely, }\forall m\geqslant 1.$
As $y,\widetilde{y}\in L_{\mathbb{F}}^{0}(\Omega;C([0,T];L^{4}))$, we have $\lim_{m\to\infty}\mathbb{P}(\sigma_{m}=T)=1$. Thus, $y=\widetilde{y}$ in $C([0,T];L^{4})$ holds $\mathbb{P}$-almost surely, proving uniqueness for the mild solution to Eq. 1.1.
Step 3. Fix any $p,q\in(2,\infty)$. By the regularity assertion in Eq. 3.1 and the growth estimate Eq. 4.1 for $F$, we deduce that $y-y^{3}$ belongs to $L_{\mathbb{F}}^{p}(\Omega\times(0,T);L^{q})$ and $F(y)$ is an element of $L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,L^{q}))$. Hence, the application of Lemma 4.3, in conjunction with Lemma 4.1, gives that
$S_{0}(y-y^{3})\in L^{p}\big{(}\Omega;L^{p}(0,T;\dot{H}^{2,q})\cap C([0,T];\dot{H}^{2-2/p-\epsilon,q})\big{)},\quad\forall\epsilon>0,$ (4.17)
while Lemma 4.4 implies $S_{1}F(y)\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{1,q})$. Additionally, the initial condition $y_{0}\in L^{\infty}$ ensures by Lemma 4.2 that $G\in L^{p}(0,T;\dot{H}^{2/p-\epsilon,q})$ for all $\epsilon>0$. Consequently, by the mild solution representation Eq. 2.7, we arrive at
$y\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2/p-\epsilon,q}),\quad\forall\epsilon>0,$
which confirms the regularity assertion Eq. 3.3. In conjunction with Eq. 4.3, we further infer that
$F(y)\in L_{\mathbb{F}}^{p}\big{(}\Omega\times(0,T);\gamma(H,\dot{H}^{2/p-\epsilon,q})\big{)},\quad\forall\epsilon>0,$
leading to the enhanced conclusion via Lemmas 4.1 and 4.4 that
$S_{1}F(y)\in L^{p}\big{(}\Omega;L^{p}(0,T;\dot{H}^{1+2/p-\epsilon,q})\cap C([0,T];\dot{H}^{1-\epsilon,q})\big{)},\quad\forall\epsilon>0.$ (4.18)
Finally, combining Eqs. 2.7, 4.17 and 4.18 establishes the regularity assertions in Eqs. 3.4 and 3.5. Moreover, the regularity assertion in Eq. 3.2 follows directly from Eq. 3.4 and the observation that $G\in C([0,T];L^{q})$, which is justified by the initial condition $y_{0}\in L^{\infty}$ and Lemma 4.2. This completes the proof of Theorem 3.1.
$\blacksquare$
4.3 Proof of Theorems 3.2 and 3.3
We begin by summarizing several well-known results. It is known from [9] that
$\|P_{h}\|_{\mathcal{L}(L^{q})}\text{ is uniformly bounded in }h\text{ for all }q\in[1,\infty].$ (4.19)
This uniform stability of $P_{h}$ implies the following approximation property:
$\|I-P_{h}\|_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha}\quad\text{for all }\alpha\in[0,2]\text{ and }q\in(1,\infty).$ (4.20)
Furthermore, it is a standard result (see, e.g., [13, Equations (12) and (13)]) that
$\lVert{I-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},L^{q})}$ $\leqslant ch^{2},\quad\forall\,q\in(1,\infty),$ (4.21)
$\lVert{I-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{1,q},L^{q})}$ $\leqslant ch,\quad\forall\,q\in[2,\infty).$ (4.22)
By applying complex interpolation between (4.21) and (4.22), one obtains
$\lVert{I-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha},\quad\forall\,\alpha\in[1,2),\quad\forall\,q\in[2,\infty).$ (4.23)
Combining this with Eq. 4.20 via the triangle inequality yields
$\lVert{P_{h}-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant ch^{\alpha},\quad\forall\alpha\in[1,2],\,\forall q\in[2,\infty),$ (4.24)
which in turn implies the estimate
$\lVert{P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}}\rVert_{\mathcal{L}(\dot{H}^{\alpha-2,q},L^{q})}\leqslant ch^{\alpha},\quad\forall\alpha\in[1,2],\,\forall q\in[2,\infty).$ (4.25)
Remark 4.2.
The projection operator $P_{h}$ can be extended to act as a linear operator from $\dot{H}^{-1,q}$ to $X_{h}$ for any $q\in(1,\infty)$. This extension is defined via the identity
$\int_{\mathcal{O}}(P_{h}v)v_{h}\,\mathrm{d}x=\langle{v,v_{h}}\rangle,$
[…]
Next, we introduce the discrete Sobolev space and recall some standard results. Let $q\in(1,\infty)$. For any $\alpha\in\mathbb{R}$, define $\dot{H}_{h}^{\alpha,q}$ as the space $X_{h}$ equipped with the norm
$\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}:=\|(-\Delta_{h})^{\alpha/2}v_{h}\|_{L^{q}},\quad v_{h}\in X_{h}.$
According to Theorem 3.1 (ii) in [19], the purely imaginary powers of $-\Delta_{h}$ are bounded in $\mathcal{L}(\dot{H}_{h}^{0,q})$ uniformly in $h$. As a consequence, an application of Theorem 4.17 from [24] yields, for any $\alpha\in(0,2)$, the equivalence
$c_{0}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\leqslant\|v_{h}\|_{[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}}\leqslant c_{1}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}},\quad v_{h}\in X_{h},$
with constants $c_{0},c_{1}>0$ independent of $h$. Here, $[\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q}]_{\alpha/2}$ denotes the complex interpolation space between $\dot{H}_{h}^{0,q}$ and $\dot{H}_{h}^{2,q}$; see [24, Chapter 2] for details. Given the standard inverse estimate
$\|v_{h}\|_{\dot{H}_{h}^{2,q}}\leqslant ch^{-2}\|v_{h}\|_{L^{q}},\quad v_{h}\in X_{h},$
complex interpolation yields the following general inverse estimate:
$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant ch^{\alpha-\beta}\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}},\quad v_{h}\in X_{h},\,\forall 0\leqslant\alpha<\beta<\infty.$ (4.26)
Applying this inverse estimate to Eq. 4.25, we obtain the estimate
$\|P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}\|_{\mathcal{L}(\dot{H}^{\alpha-2,q},\dot{H}_{h}^{\beta,q})}\leqslant ch^{\alpha-\beta}\quad\text{for all }\alpha\in[1,2],\,\beta\in[0,1],\,q\in[2,\infty).$ (4.27)
In analogy with Lemma 4.1, we also have the following $h$-uniform embedding result.
Lemma 4.7.
Let $0\leqslant\beta<\gamma\leqslant 1$ and $p,q\in(1,\infty)$. Then
$\|v_{h}\|_{\dot{H}_{h}^{\beta,q}}\leqslant c\|v_{h}\|_{(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{\gamma/2,p}},\quad v_{h}\in X_{h}.$
Additionally, we establish the following $h$-uniform Sobolev-type embeddings into $L^{\infty}$, whose proofs are given in Appendix A.
Lemma 4.8.
The following continuous embeddings hold:
  1. (i)
    If $q\in(3/2,\infty)$, then $\|v_{h}\|_{L^{\infty}}\leqslant c\|v_{h}\|_{\dot{H}_{h}^{2,q}}$ for all $v_{h}\in X_{h}$.
  2. (ii)
    If $\alpha\in(0,2)$ and $q\in(3/\alpha,\infty)\cap[2,\infty)$, then $\|v_{h}\|_{L^{\infty}}\leqslant c\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}$ for all $v_{h}\in X_{h}$.
Now, we introduce the discrete semigroup $e^{t\Delta_{h}}$ and the maximal $L^{p}$-regularity estimates associated with this semigroup. Let $e^{t\Delta_{h}}$, for $t\in[0,\infty)$, denote the analytic semigroup generated by the discrete Laplace operator $\Delta_{h}$. The following fundamental properties hold, with proofs provided in Appendix A.
Lemma 4.9.
The following estimates hold:
  • (i)
    For any $\alpha\in[0,1]$, $q\in(1,\infty)$, and $t>0$,
    $\lVert{(-\Delta_{h})^{\alpha}e^{t\Delta_{h}}}\rVert_{\mathcal{L}(L^{q})}\leqslant ct^{-\alpha}.$ (4.28)
  • (ii)
    For any $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$,
    $\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}+\lVert{(I-P_{h})e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant c\min\Big{\{}h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\Big{\}}.$ (4.29)
For a function $g_{h}\in L^{1}(0,T;\dot{H}_{h}^{0,2})$, the discrete convolution operator $S_{0,h}$ is defined by
$(S_{0,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}s,\quad t\in[0,T].$ (4.30)
For a stochastic process $g_{h}\in L_{\mathbb{F}}^{2}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))$, the discrete stochastic convolution operator $S_{1,h}$ is given by
$(S_{1,h}g_{h})(t):=\int_{0}^{t}e^{(t-s)\Delta_{h}}g_{h}(s)\,\mathrm{d}W_{H}(s),\quad t\in[0,T].$ (4.31)
These operators satisfy the following fundamental discrete maximal $L^{p}$-regularity estimates.
Lemma 4.10.
Suppose $p,q\in(1,\infty)$. For any $g_{h}\in L^{p}(0,T;\dot{H}_{h}^{0,q})$, the following discrete maximal $L^{p}$-regularity estimate is valid:
$\lVert{S_{0,h}g_{h}}\rVert_{C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1-1/p,p}}+\lVert{S_{0,h}g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{2,q})}\leqslant c\lVert{g_{h}}\rVert_{L^{p}(0,T;\dot{H}_{h}^{0,q})}.$
Lemma 4.11.
Let $p\in(2,\infty)$ and $q\in[2,\infty)$. For any $g_{h}\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))$, the following discrete stochastic maximal $L^{p}$-regularity estimate holds:
$\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}+\lVert{S_{1,h}g_{h}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,q})}$ (4.32)
[…]
Remark 4.3.
A proof of Lemma 4.10 can be found in [12, Theorem 3.2]. Since the boundedness constant of the $H^{\infty}$-calculus for $-\Delta_{h}$ is independent of the spatial mesh size $h$, as shown in [19, Theorem 3.1], Lemma 4.11 follows directly from [32, Theorem 3.5]; see [19, Theorem 3.2] for further details.
We proceed to present some auxiliary estimates.
Lemma 4.12.
Suppose that $y_{0}\in L^{\infty}$ is deterministic. Consider the mild solution $y$ to equation (1.1), the strong solution $y_{h}$ to equation (2.9), and the function $G$ defined by equation (2.8). Additionally, let $G_{h}(t):=e^{t\Delta_{h}}P_{h}y_{0}$ for $t\in[0,T]$. For notational simplicity, let $z:=y-G$ and $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z$. For any $p,q\in(2,\infty)$ and $\epsilon>0$, the following inequalities hold:
$\lVert{G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}+\lVert{P_{h}G-G_{h}}\rVert_{L^{p}(0,T;L^{q})}\leqslant ch^{\frac{2}{p}},$ (4.33)
$\lVert{z-P_{h}z}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\epsilon},$ (4.34)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{1+\frac{2}{p}-\epsilon},$ (4.35)
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{1-\epsilon},$ (4.36)
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\epsilon}.$ (4.37)
Furthermore, for any $\beta\in\Big{(}1,\frac{3p}{3p-4}\Big{)}$, the following inequalities hold:
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p},$ (4.38)
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{\xi_{h}(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\big{(}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}+\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\big{)}\,\mathrm{d}t\bigg{]}\leqslant ch^{p+2-\frac{3}{2}\big{(}1-\frac{1}{\beta}\big{)}p-\epsilon}.$ (4.39)
Here, $\beta^{\prime}$ denotes the Hölder conjugate of $\beta$.
Proof.
The inequality Eq. 4.33 is evident by Eq. 4.29, and the inequality Eq. 4.34 follows by Eq. 4.20 and the regularity assertion Eq. 3.5. The inequality Eq. 4.35 is derived as follows:
[…]
$=c\lVert{(P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h})\Delta z}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{0,q})}$
[…]
$=ch^{1+2/p-\epsilon}\lVert{z}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}^{1+2/p-\epsilon,q})}$
[…]
The inequality Eq. 4.36 is derived in a similar manner:
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$
[…]
For Eq. 4.37, we apply Hölder’s inequality with $\alpha\in(1,2)$, yielding
$\mathbb{E}\left[\int_{0}^{T}\lVert{(y-P_{h}y)(t)}\rVert_{L^{4}}^{p}\lVert{y(t)}\rVert_{L^{8}}^{2p}\,\mathrm{d}t\right]$
[…]
Choosing $\alpha>1$ sufficiently close to 1 so that $\frac{2}{\alpha}-\frac{\epsilon}{2}\geqslant 2-\epsilon$, and using the regularity estimate (3.3), we obtain (4.37).
Now, let us prove Eq. 4.38. Fix any $\beta\in(1,\frac{3p}{3p-4})$. By using Eq. 4.19 and following the reasoning behind [21, Equation (2.2)], we establish that
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{\dot{H}^{1,2}}^{2}\,\mathrm{d}t\bigg{]}$
is uniformly bounded with respect to $h$. Using the continuous embedding of $\dot{H}^{1,2}$ into $L^{6}$, this uniform bound extends to
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2}\,\mathrm{d}t\bigg{]}.$
Employing Hölder’s inequality, we deduce for all $t\in[0,T]$ that
$\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{\frac{4\beta}{3\beta-3}}\leqslant\lVert{y_{h}(t)}\rVert_{L^{2}}^{\frac{4\beta}{3\beta-3}-2}\lVert{y_{h}(t)}\rVert_{L^{6}}^{2},$
which in turn confirms the uniform bound on
$\lVert{y_{h}}\rVert_{L^{\frac{4\beta}{3\beta-3}}(\Omega\times(0,T);L^{2\beta})}.$
Subsequently, applying Hölder’s inequality once more yields
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{2\beta}^{p}\,\mathrm{d}t\bigg{]}$
[…]
Using Eq. 4.33 with $p$ replaced by $\frac{4p\beta}{4\beta-(3\beta-3)p}$ and $q$ by $2\beta^{\prime}$, we conclude that
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y_{h}(t)}\rVert_{L^{2\beta}}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3p}{2}(1-\frac{1}{\beta})}.$
Moreover, by Theorem 3.1, we have $y\in L^{\frac{4\beta}{3(\beta-1)}}(\Omega\times(0,T);L^{2\beta})$, so a similar argument yields
$\mathbb{E}\bigg{[}\int_{0}^{T}\lVert{(P_{h}G-G_{h})(t)}\rVert_{L^{2\beta^{\prime}}}^{p}\lVert{y(t)}\rVert_{L^{2\beta}}^{p}\,\mathrm{d}t\bigg{]}\leqslant ch^{2-\frac{3p}{2}(1-\frac{1}{\beta})}.$
Combining both estimates gives Eq. 4.38. The estimate Eq. 4.39 follows by a similar argument using Eq. 4.35. This completes the proof.
Finally, we are in a position to prove Theorems 3.2 and 3.3.
Proof of Theorem 3.2. Fix any $p\in(2,\infty)$. Define
$z(t)$ $:=y(t)-G(t),\quad 0\leqslant t\leqslant T,$ (4.40)
$z_{h}(t)$ $:=y_{h}(t)-G_{h}(t),\quad 0\leqslant t\leqslant T,$ (4.41)
where $G$ is defined by Eq. 2.8, and $G_{h}$ is given by
$G_{h}(t):=e^{t\Delta_{h}}P_{h}y_{0},\quad t\in[0,T].$ (4.42)
From Eq. 1.1 and the fact that $G^{\prime}(t)=\Delta G(t)$ for $t\in[0,T]$, we deduce that
$\mathrm{d}z(t)=\big{(}\Delta z+z+G-y^{3}\big{)}(t)\,\mathrm{d}t+F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Consequently,
$\mathrm{d}P_{h}z(t)=P_{h}(\Delta z+z+G-y^{3})(t)\,\mathrm{d}t+P_{h}F(y(t))\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Similarly, from Eq. 2.9 and the fact that $G_{h}^{\prime}(t)=\Delta_{h}G_{h}(t)$ for $t\in[0,T]$, it follows that
$\mathrm{d}z_{h}(t)=(\Delta_{h}z_{h}+z_{h}+G_{h}-P_{h}y_{h}^{3})(t)\,\mathrm{d}t+P_{h}F\big{(}y_{h}(t)\big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Define $e_{h}:=z_{h}-P_{h}z$. Then we obtain
$\mathrm{d}e_{h}(t)$ $=\Big{(}\Delta_{h}e_{h}+\Delta_{h}P_{h}z-P_{h}\Delta z+e_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.43)
$\quad P_{h}\Big{(}F\big{(}y_{h}(t)\big{)}-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
Now define $\eta_{h}:=e_{h}-\xi_{h}$, where $\xi_{h}:=S_{0,h}(\Delta_{h}P_{h}z-P_{h}\Delta z)$, and consider the differential equation
$\Big{(}\frac{\mathrm{d}}{\mathrm{d}t}-\Delta_{h}\Big{)}\xi_{h}(t)=(\Delta_{h}P_{h}z-P_{h}\Delta z)(t),\quad t\geqslant 0.$
Using Eq. 4.43, we conclude that
$\mathrm{d}\eta_{h}(t)$ $=\Big{(}\Delta_{h}\eta_{h}+\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}\big{(}y^{3}-y_{h}^{3}\big{)}\Big{)}(t)\,\mathrm{d}t+{}$ (4.44)
$\quad P_{h}\Big{(}F\big{(}y_{h}(t)\big{)}-F\big{(}y(t)\big{)}\Big{)}\,\mathrm{d}W_{H}(t),\quad 0\leqslant t\leqslant T.$
By applying Itô’s formula to $\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ and using the initial condition $\eta_{h}(0)=0$ together with the definition of $F$ given by Eq. 2.4, we deduce that for any $0\leqslant t\leqslant T$,
$\mathbb{E}\lVert{\eta_{h}(t)}\rVert_{L^{2}}^{p}$ $=p\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Big{(}\Delta_{h}\eta_{h}(s)+\xi_{h}(s)+\eta_{h}(s)$
$\qquad\qquad\qquad\qquad\qquad{}+(G_{h}-P_{h}G)(s)+\big{(}y^{3}-y_{h}^{3}\big{)}(s)\Big{)}\,\mathrm{d}x\,\mathrm{d}s\biggr{]}$
$\qquad{}+\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]}$
$\qquad{}+\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}x\Big{)}^{2}\,\mathrm{d}s\biggr{]}$
$=I_{1}+I_{2}+I_{3}+I_{4}+I_{5}+I_{6},$
where
$\allowdisplaybreaks I_{1}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\Delta_{h}\eta_{h}(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
$I_{2}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot(\xi_{h}+\eta_{h}+G_{h}-P_{h}G)(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
$I_{3}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}y^{3}-(P_{h}y)^{3}\big{]}(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
$I_{4}$ $:=p\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{[}(P_{h}y)^{3}-y_{h}^{3}\big{]}(s)\,\mathrm{d}x\,\mathrm{d}s\biggr{]},$
$I_{5}$ $:=\frac{p}{2}\,\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lVert{P_{h}\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}}\rVert_{L^{2}}^{2}\,\mathrm{d}s\biggr{]},$
$I_{6}$ $:=\frac{p(p-2)}{2}\mathbb{E}\biggl{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-4}\sum_{k=0}^{\infty}\Big{(}\int_{\mathcal{O}}\eta_{h}(s)\cdot\big{(}f_{k}(y_{h}(s))-f_{k}(y(s))\big{)}\,\mathrm{d}x\Big{)}^{2}\,\mathrm{d}s\biggr{]}.$
Next, we estimate the terms $I_{1}$, $I_{2}$, $I_{3}$, and $I_{4}$. For the term $I_{1}$, by the definition of $\Delta_{h}$, we obtain
$I_{1}=-p\mathbb{E}\biggl{[}\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{2}}^{p-2}\|{\eta_{h}(s)}\|_{\dot{H}^{1,2}}^{2}\,\mathrm{d}s\biggr{]}.$ (4.45)
Regarding $I_{2}$, applying Hölder’s inequality, Young’s inequality, and the bounds given in Eq. 4.33 and Eq. 4.35, we deduce that
$I_{2}\leqslant ch^{2}+c\int_{0}^{t}\|{\eta_{h}(s)}\|_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$ (4.46)
For $I_{3}$, using Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ stated in Eq. 4.19, a straightforward computation gives
$I_{3}$ $\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\Big{(}\lVert{y(s)}\rVert_{L^{8}}^{2}+\lVert{P_{h}y(s)}\rVert_{L^{8}}^{2}\Big{)}\,\mathrm{d}s\bigg{]}$
$\leqslant c\,\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-1}\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}\lVert{y(s)}\rVert_{L^{8}}^{2}\,\mathrm{d}s\bigg{]}$
$\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\left(\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y-P_{h}y)(s)}\rVert_{L^{4}}^{p}\lVert{y(s)}\rVert_{L^{8}}^{2p}\right)\,\mathrm{d}s\bigg{]}.$
Using the estimate Eq. 4.37, we then arrive at the bound
$I_{3}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s,\quad\forall\epsilon>0.$ (4.47)
To estimate $I_{4}$, we proceed as follows. Using the definitions in (4.40) and (4.41), together with the identity $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we find that
$P_{h}y-y_{h}=-\xi_{h}-\eta_{h}+P_{h}G-G_{h}.$
Thus, for any $s\in[0,t]$, we compute:
$\int_{\mathcal{O}}\eta_{h}(s)\left[(P_{h}y)^{3}-y_{h}^{3}\right](s)\,\mathrm{d}x$
[…]
Hence, applying Hölder’s inequality, Young’s inequality, and the stability property of $P_{h}$ from (4.19), we derive the following estimate for $I_{4}$:
$I_{4}$ $\leqslant c\mathbb{E}\int_{0}^{t}\bigg{[}\|\eta_{h}(s)\|_{L^{2}}^{p}+\left(\|\xi_{h}(s)\|^{p}_{L^{2\beta^{\prime}}}+\|(P_{h}G-G_{h})(s)\|^{p}_{L^{2\beta^{\prime}}}\right)$
×$\qquad\qquad\qquad{}\times\Big{(}\|y(s)\|^{p}_{L^{2\beta}}+\|y_{h}(s)\|^{p}_{L^{2\beta}}\Big{)}\bigg{]}\,\mathrm{d}s,\quad\forall\beta\in\left(1,\frac{3p}{3p-4}\right),$
where $\beta^{\prime}$ denotes the Hölder conjugate of $\beta$. Using the estimates (4.38) and (4.39), this simplifies further to
$I_{4}\leqslant c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s+ch^{2-\frac{3}{2}p\left(1-\frac{1}{\beta}\right)},\quad\forall\beta\in\left(1,\frac{3p}{3p-4}\right).$
By choosing $\beta$ sufficiently close to 1 within the interval $\left(1,\frac{3p}{3p-4}\right)$, for any $\epsilon>0$, we finally obtain
$I_{4}\leqslant ch^{2-\epsilon}+c\int_{0}^{t}\|\eta_{h}(s)\|^{p}_{L^{p}(\Omega;L^{2})}\,\mathrm{d}s,\quad\forall\epsilon>0.$ (4.48)
Now, we proceed to estimate $I_{5}+I_{6}$. Using Hölder’s inequality and the stability property of $P_{h}$ given in Eq. 4.19, we obtain the following bound:
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\sum_{k=0}^{\infty}\lVert{f_{k}(y_{h}(s))-f_{k}(y(s))}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
Invoking the condition Eq. 2.3, we further deduce that
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p-2}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{2}\,\mathrm{d}s\bigg{]}.$
Applying Young’s inequality, we then have
$I_{5}+I_{6}\leqslant c\mathbb{E}\bigg{[}\int_{0}^{t}\left(\lVert{\eta_{h}(s)}\rVert_{L^{2}}^{p}+\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\right)\,\mathrm{d}s\bigg{]}.$
Recalling the definitions in Eqs. 4.40 and 4.41 and using the identity $z_{h}-P_{h}z=\xi_{h}+\eta_{h}$, we infer that
$(y_{h}-y)(s)=\xi_{h}(s)+\eta_{h}(s)-(z-P_{h}z)(s)-(G-G_{h})(s),\quad\forall s\in[0,T].$ (4.49)
Combining this with the estimates Eq. 4.33, Eq. 4.34, and Eq. 4.35, we find that
$\mathbb{E}\bigg{[}\int_{0}^{t}\lVert{(y_{h}-y)(s)}\rVert_{L^{2}}^{p}\,\mathrm{d}s\bigg{]}\leqslant ch^{2}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$
As a result, we arrive at the following bound for $I_{5}+I_{6}$:
$I_{5}+I_{6}\leqslant ch^{2}+c\int_{0}^{t}\lVert{\eta_{h}(s)}\rVert_{L^{p}(\Omega;L^{2})}^{p}\,\mathrm{d}s.$ (4.50)
Finally, by combining the estimates for $I_{1}$, $I_{2}$, $I_{3}$, $I_{4}$, and $I_{5}+I_{6}$, as established in Eqs. 4.45, 4.46, 4.47, 4.48 and 4.50, we obtain
$\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}^{p}+\mathbb{E}\bigg{[}\int_{0}^{t}\|\eta_{h}(s)\|_{L^{2}}^{p-2}\|\eta_{h}(s)\|_{\dot{H}_{h}^{1,2}}^{2}\,\mathrm{d}s\bigg{]}$
[…]
An application of Gronwall’s inequality then yields
$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{(2-\epsilon)/p},\quad\forall\epsilon>0.$
It follows that
$\max_{t\in[0,T]}\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2/p-\epsilon},\quad\forall\epsilon>0.$ (4.51)
Moreover, for any $\epsilon>0$ and $t\in(0,T]$, the following bounds hold:
$\lVert{(z-P_{h}z)(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{1-\epsilon},$ (see Eqs. 4.20 and 3.4),
[…]
Now, from the decomposition Eq. 4.49, we deduce that
$\|(y-y_{h})(t)\|_{L^{p}(\Omega;L^{2})}$ $\leqslant\|(z-P_{h}z)(t)\|_{L^{p}(\Omega;L^{2})}+\|(G-G_{h})(t)\|_{L^{p}(\Omega;L^{2})}$
$\qquad{}+\|\xi_{h}(t)\|_{L^{p}(\Omega;L^{2})}+\|\eta_{h}(t)\|_{L^{p}(\Omega;L^{2})},\quad\forall t\in[0,T].$
Combining the above estimates, we conclude that for all $t\in(0,T]$,
$\lVert{(y-y_{h})(t)}\rVert_{L^{p}(\Omega;L^{2})}\leqslant ch^{2/p-\epsilon}+ch^{1-\epsilon}+ch^{2/p}t^{-1/p},\quad\forall\epsilon>0.$
Since $h^{1-\epsilon}$ is absorbed into $h^{2/p-\epsilon}$ for $p>2$, the desired error estimate Eq. 3.6 follows immediately. This completes the proof of Theorem 3.2.
$\blacksquare$
Proof of Theorem 3.3. Let $G$ and $z$ be defined by Eqs. 2.8 and 4.40, respectively. Since the initial condition $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$ implies that $y_{0}\in\dot{H}^{2,q}$ for all $q\in(2,\infty)$, it follows from Lemma 4.2 that $G\in C([0,T];\dot{H}^{2,q})$ for all $q\in(2,\infty)$. Consequently, by adapting the argument in Step 3 of the proof of Theorem 3.1, we obtain the following improved regularity properties for the processes $y$ and $z$:
$y\in L_{\mathbb{F}}^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1);$ (4.52)
$z\in L_{\mathbb{F}}^{p}(\Omega\times(0,T);\dot{H}^{2,q}),$ $\forall p,q\in(2,\infty);$ (4.53)
$z\in L_{\mathbb{F}}^{p}(\Omega;C([0,T];\dot{H}^{2-\epsilon,q})),$ $\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$ (4.54)
Let $G_{h}$ be defined as in Eq. 4.42. We retain the notation introduced in the proof of Theorem 3.2:
$z_{h}$ $:=y_{h}-G_{h},$ $e_{h}:=z_{h}-P_{h}z,$
$\xi_{h}$ $:=S_{0,h}(\Delta_{h}P_{h}-P_{h}\Delta)z,$ $\eta_{h}:=e_{h}-\xi_{h}.$
The remainder of the proof is structured into the following three steps.
Step 1. Let us present some key estimates that form the foundation of our subsequent analysis. We begin by establishing the following estimates:
$\left\|z-P_{h}z\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p,q\in(2,\infty),$ (4.55)
$\left\|G-G_{h}\right\|_{C([0,T];L^{q})}+\left\|P_{h}G-G_{h}\right\|_{C([0,T];L^{q})}\leqslant ch^{2},\quad\forall q\in(2,\infty).$ (4.56)
Estimate (4.55) is a direct consequence of (4.20) combined with the regularity property (4.53). Estimate (4.56) follows from (4.29) and the initial condition $y_{0}\in W_{0}^{1,\infty}(\mathcal{O})\cap W^{2,\infty}(\mathcal{O})$. Given the regularity Eq. 4.53, estimate Eq. 4.35 improves to
$\left\|\xi_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant ch^{2},\quad\forall p,q\in(2,\infty),$ (4.57)
and estimate Eq. 4.36 improves to
$\lVert{\xi_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-2/p-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$ (4.58)
Following the derivation of estimate Eq. 4.51 for $\eta_{h}$, we obtain the improved bound
$\max_{0\leqslant t\leqslant T}\left\|\eta_{h}(t)\right\|_{L^{p}(\Omega;L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.59)
Therefore, combining estimates Eqs. 4.57, 4.59, 4.56 and 4.55 via the decomposition Eq. 4.49 yields the error estimate
$\left\|y-y_{h}\right\|_{L^{p}(\Omega\times(0,T);L^{2})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.60)
Furthermore, for any $p\in(2,\infty)$, we deduce the following chain of inequalities:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}$
[…]
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\leqslant}}{}$ $\lVert{P_{h}y}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}+ch^{-3/2}\lVert{P_{h}(y-y_{h})}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}+ch^{-3/2}\lVert{y-y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{2})}$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\leqslant}}{}$ $c\lVert{y}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}+ch^{1/2},$
where in (i) we apply the inverse estimate $\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant ch^{-3/2}\lVert{v_{h}}\rVert_{L^{2}}$ for all $v_{h}\in X_{h}$ (see, e.g., [7, Theorem 4.5.11]); in (ii) we utilize the stability property of $P_{h}$ as stated in Eq. 4.19; and in (iii) we employ the error estimate Eq. 4.60. By Eq. 4.52 and the continuous embedding $\dot{H}^{2-\epsilon,q}\hookrightarrow L^{\infty}$ for sufficiently small $\epsilon>0$, we conclude $h$-uniform boundedness:
$\lVert{y_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}\text{ is uniformly bounded with respect to $h$ for all $p\in(2,\infty)$}.$ (4.61)
Step 2. We now prove the error estimate Eq. 3.7. The proof comprises two parts, (a) and (b).
Part (a). We establish the estimate
$\|\eta_{h}\|_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.62)
Utilizing (4.44) and the definitions of the discrete convolution operators $S_{0,h}$ and $S_{1,h}$ given in Eqs. 4.30 and 4.31, we derive the decomposition
$\eta_{h}=II_{1}+II_{2},$ (4.63)
where
$II_{1}$ $:=S_{0,h}\big{(}\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})\big{)},$
$II_{2}$ $:=S_{1,h}P_{h}\big{(}F(y_{h})-F(y)\big{)}.$
Fix $p\in(2,\infty)$. For $II_{1}$, applying Lemma 4.8(i) and Lemma 4.10 yields
$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}$ $\leqslant c\|II_{1}\|_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{2,5/3})}$
$\leqslant c\|\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{5/3})}$
$\leqslant c\|\xi_{h}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}+c\|\eta_{h}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}$
$\quad{}+c\|G_{h}-P_{h}G\|_{L^{p}(\Omega\times(0,T);L^{5/3})}+c\|P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{5/3})}.$
Employing the estimates Eqs. 4.59, 4.56 and 4.57 and the embedding $L^{q}\hookrightarrow L^{5/3}$ for $q>5/3$, we obtain
$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2}+\|P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{5/3})}.$
Since $P_{h}$ is bounded on $L^{5/3}$ (see (4.19)), it follows that
$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2}+\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}.$
To estimate $\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}$, we use the identity
$y^{3}-y_{h}^{3}=(y-y_{h})(y^{2}+yy_{h}+y_{h}^{2})$
and apply Hölder’s inequality, leading to
$\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}\leqslant c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2})}\Big{(}\|y\|_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}+\|y_{h}\|_{L^{4p}(\Omega\times(0,T);L^{20})}^{2}\Big{)}.$
By the regularity of $y$ (see (4.52)), the stability of $y_{h}$ (see (4.61)), and the error estimate (4.60), we deduce
$\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{5/3})}\leqslant ch^{2}.$ (4.64)
Therefore,
$\|II_{1}\|_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2}.$ (4.65)
For $II_{2}$, the continuous embedding $\dot{H}^{1,2}\hookrightarrow L^{6}$ and the isometric embedding of $\dot{H}_{h}^{1,2}$ into $\dot{H}^{1,2}$ imply
$\|II_{2}\|_{L^{p}(\Omega\times(0,T);L^{6})}\leqslant c\|II_{2}\|_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,2})}.$
Hence, applying Lemma 4.11, the ideal property of $\gamma$-radonifying operators [16, Theorem 9.1.10], the boundedness of $P_{h}$ on $L^{2}$, the Lipschitz condition (4.4), and the estimate (4.60), we obtain
$\|II_{2}\|_{L^{p}(\Omega\times(0,T);L^{6})}$ $\leqslant c\|P_{h}\big{(}F(y_{h})-F(y)\big{)}\|_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,2}))}$
$\leqslant c\|F(y_{h})-F(y)\|_{L^{p}(\Omega\times(0,T);\gamma(H,L^{2}))}$
$\leqslant c\|y-y_{h}\|_{L^{p}(\Omega\times(0,T);L^{2})}$
$\leqslant ch^{2}.$
Combining (4.65) with this estimate via the decomposition (4.63) establishes (4.62).
Part (b). We consider the decomposition of $y_{h}-y$ given in Eq. 4.49, and combine it with the estimates Eq. 4.55, Eq. 4.56, Eq. 4.57, and Eq. 4.62 to establish Eq. 3.7 for the particular case $p\in(2,\infty)$, $q=6$. Based on this, we now apply Lemma 4.8(ii) and adapt the same bounding strategy used for $II_{2}$ in part (a). This yields an improved estimate:
$\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}$ $\leqslant c\lVert{II_{2}}\rVert_{L^{p}(\Omega\times(0,T);\dot{H}_{h}^{1,6})}$
$\leqslant c\lVert{P_{h}(F(y_{h})-F(y))}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,6}))}$
$\leqslant c\lVert{F(y_{h})-F(y)}\rVert_{L^{p}(\Omega\times(0,T);\gamma(H,L^{6}))}$
$\leqslant c\lVert{y_{h}-y}\rVert_{L^{p}(\Omega\times(0,T);L^{6})}$
$\leqslant ch^{2},\quad\forall p\in(2,\infty).$
Combining this refined bound with the estimate for $II_{1}$ given in Eq. 4.65, and recalling the decomposition Eq. 4.63, we deduce that
$\lVert{\eta_{h}}\rVert_{L^{p}(\Omega\times(0,T);L^{\infty})}\leqslant ch^{2},\quad\forall p\in(2,\infty).$ (4.66)
Finally, revisiting the decomposition Eq. 4.49 and combining the estimates Eq. 4.55, Eq. 4.56, Eq. 4.57, and Eq. 4.66, we conclude the error estimate Eq. 3.7 for all $p,q\in(2,\infty)$.
Step 3. We now establish the error estimate Eq. 3.8, proceeding in three parts, (a)–(c).
Part (a). We aim to prove the inequality
$\|\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}\leqslant ch^{2},\quad\forall p\in(4,\infty).$ (4.67)
To this end, fix $p\in(4,\infty)$ and let $q\in(6,\infty)$ be arbitrary. Recall the decomposition Eq. 4.63 for $\eta_{h}$. By Lemma 4.8(ii) and Lemma 4.7, we have
$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant\|II_{1}\|_{L^{p}(\Omega;C([0,T];\dot{H}_{h}^{2-3/p,q}))}$
$\leqslant\|II_{1}\|_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}.$
An application of Lemma 4.10 then gives
$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant c\|\xi_{h}+\eta_{h}+G_{h}-P_{h}G+P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
Using the estimates Eq. 4.56, Eq. 4.57, and Eq. 4.66, we deduce
$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|P_{h}(y^{3}-y_{h}^{3})\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
Since the projection operator $P_{h}$ is bounded on $L^{q}$ (cf. Eq. 4.19), it follows that
$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
As in Eq. 4.64, we use the identity
$y^{3}-y_{h}^{3}=(y-y_{h})(y^{2}+yy_{h}+y_{h}^{2})$
and apply Hölder’s inequality to obtain
$\|y^{3}-y_{h}^{3}\|_{L^{p}(\Omega\times(0,T);L^{q})}\leqslant c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
Therefore,
$\|II_{1}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
Next, we estimate $II_{2}$ from Eq. 4.63. Applying Lemma 4.7 and Lemma 4.8(ii), we get
$\|II_{2}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant\|II_{2}\|_{L^{p}(\Omega;C([0,T];(\dot{H}_{h}^{0,q},\dot{H}_{h}^{2,q})_{1/2-1/p,p}))}$
$\leqslant c\|P_{h}(F(y_{h})-F(y))\|_{L^{p}(\Omega\times(0,T);\gamma(H,\dot{H}_{h}^{0,q}))},$
where the second inequality follows from Lemma 4.11. As in the estimate of $II_{2}$ in Step 2, Part (a), we then obtain
$\|II_{2}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}\leqslant c\|y-y_{h}\|_{L^{p}(\Omega\times(0,T);L^{q})}.$
Combining the bounds for $II_{1}$ and $II_{2}$, and recalling the decomposition Eq. 4.63, we arrive at
$\|\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2}+c\|y-y_{h}\|_{L^{p}(\Omega\times(0,T);L^{q})}+c\|y-y_{h}\|_{L^{2p}(\Omega\times(0,T);L^{2q})}.$
Now, since the error estimate Eq. 3.7 has already been established for all $p,q\in(2,\infty)$, we conclude that the right-hand side is of order $h^{2}$. This completes the proof of Eq. 4.67.
Part (b). We now prove the estimate
$\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-2/p-\varepsilon},\quad\forall p,q\in(2,\infty),\ \forall\varepsilon\in(0,1).$ (4.68)
By the approximation property Eq. 4.20 and the regularity result Eq. 4.54, we have
$\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2-\varepsilon},\quad\forall p,q\in(2,\infty),\ \forall\varepsilon\in(0,1).$ (4.69)
Moreover, combining the approximation result Eq. 4.56 with the bound Eq. 4.67, we deduce
$\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}\leqslant ch^{2},\quad\forall p,q\in(2,\infty).$ (4.70)
Using the decomposition Eq. 4.49, we write
$\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant\|\xi_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}+\|z-P_{h}z\|_{L^{p}(\Omega;C([0,T];L^{q}))}$
$\quad+\|G-G_{h}+\eta_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}.$
Substituting the bounds Eqs. 4.69 and 4.70 into this inequality, we obtain
$\|y-y_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant ch^{2-\varepsilon}+\|\xi_{h}\|_{L^{p}(\Omega;C([0,T];L^{q}))},\quad\forall p,q\in(2,\infty),\ \forall\varepsilon\in(0,1).$
Finally, invoking the estimate Eq. 4.58, we conclude the desired result Eq. 4.68.
Part (c). Let $\Pi_{h}$ denote the standard Lagrange interpolation operator onto the finite element space $X_{h}$. Combining the regularity result Eq. 4.52 with standard interpolation theory yields
$\lVert{y-\Pi_{h}y}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant ch^{2-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$
The triangle inequality applied to this estimate and Eq. 4.68 then gives
$\lVert{y_{h}-\Pi_{h}y}\rVert_{L^{p}(\Omega;C([0,T];L^{q}))}$ $\leqslant ch^{2-2/p-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$
Applying the standard finite element inverse estimate (cf. [7, Theorem 4.5.11])
$\lVert{v_{h}}\rVert_{L^{\infty}}\leqslant ch^{-3/q}\lVert{v_{h}}\rVert_{L^{q}},\quad\forall v_{h}\in X_{h},\,q\in(2,\infty),$
to the preceding bound implies
$\lVert{y_{h}-\Pi_{h}y}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2-2/p-3/q-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$
Furthermore, the regularity Eq. 4.52 combined with standard interpolation estimates yields
$\lVert{y-\Pi_{h}y}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2-3/q-\epsilon},\quad\forall p\in(2,\infty),\,\forall q\in(3,\infty),\,\forall\epsilon\in(0,1).$
Applying the triangle inequality to these last two estimates now provides
$\lVert{y-y_{h}}\rVert_{L^{p}(\Omega;C([0,T];L^{\infty}))}$ $\leqslant ch^{2-2/p-3/q-\epsilon},\quad\forall p,q\in(2,\infty),\,\forall\epsilon\in(0,1).$
Since $p,q\in(2,\infty)$ can be chosen arbitrarily large, the term $2/p+3/q$ can be made arbitrarily small and is therefore absorbed into $\epsilon>0$. This yields the desired error estimate Eq. 3.8, completing the proof of Theorem 3.3.
$\blacksquare$
5 Conclusions
In this work, we have rigorously investigated the regularity properties of the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, in the presence of non-smooth initial data. For the spatial semidiscretization using the finite element method, we have established a convergence rate of $O(h^{2/p-\epsilon}+h^{2/p}t^{-1/p})$ with non-smooth initial data. Meanwhile, under the condition of smooth initial data, employing the discrete stochastic maximal $L^{p}$-regularity theory, we have achieved a convergence rate of $O(h^{2})$ in the $L^{p}(\Omega\times(0,T);L^{q})$ norm, along with a pathwise uniform convergence rate of $O(h^{2-\epsilon})$.
The theoretical results derived here lay a foundation for the numerical analysis of fully discrete approximations based on the finite element method for the three-dimensional stochastic Allen-Cahn equation with multiplicative noise, within the context of general spatial $L^{q}$-norms.
Appendix A Proofs of Lemmas 4.8 and 4.9
The discrete Laplace operator $\Delta_{h}$ has the following fundamental property; see [1].
Lemma A.1.
For any $q\in(1,\infty)$ and $\theta\in(\pi/2,\pi)$, there exists a constant $c>0$ independent of $h$, depending only on $\theta$, $q$, and the regularity constants of $\mathcal{K}_{h}$, such that
$\left\|(re^{i\omega}-\Delta_{h})^{-1}\right\|_{\mathcal{L}(\dot{H}_{h}^{0,q})}\leqslant\frac{c}{1+r},\quad\forall r\geqslant 0,\quad\forall\omega\in[-\theta,\theta].$
Lemma A.2.
For any $\alpha\in[0,2]$ and $q\in(1,\infty)$, the operator norm $\|P_{h}\|_{\mathcal{L}(\dot{H}^{\alpha,q},\dot{H}_{h}^{\alpha,q})}$ is uniformly bounded with respect to $h$.
Proof.
Fix $q\in(1,\infty)$. By the estimates Eqs. 4.20 and 4.21, and applying the triangle inequality, we obtain
$\lVert{P_{h}-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},L^{q})}\leqslant ch^{2}.$
Using the inverse estimate (4.26), it follows that
$\lVert{P_{h}-\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},\dot{H}_{h}^{2,q})}\leqslant c.$
On the other hand, the stability estimate (4.19) gives
$\lVert{\Delta_{h}^{-1}P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},\dot{H}_{h}^{2,q})}=\lVert{P_{h}\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},L^{q})}\leqslant c\lVert{\Delta}\rVert_{\mathcal{L}(\dot{H}^{2,q},L^{q})}=c,$
where we have used the fact that $\Delta:\dot{H}^{2,q}\to L^{q}$ is an isometric isomorphism. Combining the above bounds and applying the triangle inequality, we conclude that
$\lVert{P_{h}}\rVert_{\mathcal{L}(\dot{H}^{2,q},\dot{H}_{h}^{2,q})}\leqslant c.$
This proves the uniform boundedness for $\alpha=2$. The case $\alpha=0$ is already established by (4.19). The general case $\alpha\in[0,2]$ follows from complex interpolation between the endpoint cases $\alpha=0$ and $\alpha=2$, which completes the proof. ∎
By the approximation property of $P_{h}$ in Eq. 4.20 and the smoothing estimate for $e^{t\Delta}$ in Lemma 4.2, a direct calculation yields the following lemma.
Lemma A.3.
For any $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$,
$\left\|(I-P_{h})e^{t\Delta}\right\|_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant c\min\left\{h^{\alpha},\,\frac{h^{2}}{t^{1-\alpha/2}}\right\}.$ (A.1)
A.1 Proof of Lemma 4.8
We only prove (ii), as (i) follows by similar arguments using Eq. 4.21 instead of Eq. 4.23. Fix $\alpha\in(0,2)$ and $q\in(3/\alpha,\infty)\cap[2,\infty)$, and let $v_{h}\in X_{h}$ be arbitrary. By Eq. 4.23, we have
$\|v_{h}-\Delta^{-1}\Delta_{h}v_{h}\|_{L^{q}}=\|-(I-\Delta_{h}^{-1}P_{h}\Delta)\Delta^{-1}\Delta_{h}v_{h}\|_{L^{q}}\leqslant ch^{\alpha}\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
Moreover, Eq. 4.20 gives
$\|(I-P_{h})\Delta^{-1}\Delta_{h}v_{h}\|_{L^{q}}\leqslant ch^{\alpha}\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
Combining these two estimates via the triangle inequality yields
$\|v_{h}-P_{h}\Delta^{-1}\Delta_{h}v_{h}\|_{L^{q}}\leqslant ch^{\alpha}\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
Applying a standard inverse estimate [7, Theorem 4.5.11], we deduce
$\|v_{h}-P_{h}\Delta^{-1}\Delta_{h}v_{h}\|_{L^{\infty}}\leqslant ch^{\alpha-3/q}\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
Since $\alpha>3/q$, it follows that
$\|v_{h}-P_{h}\Delta^{-1}\Delta_{h}v_{h}\|_{L^{\infty}}\leqslant c\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
The Sobolev embedding $\dot{H}^{\alpha,q}\hookrightarrow L^{\infty}$, which holds due to $\alpha q>3$, together with the $L^{\infty}$-stability of $P_{h}$ (see Eq. 4.19), implies
$\|P_{h}\Delta^{-1}\Delta_{h}v_{h}\|_{L^{\infty}}\leqslant c\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}.$
Applying the triangle inequality once more, we obtain
$\|v_{h}\|_{L^{\infty}}\leqslant c\|\Delta^{-1}\Delta_{h}v_{h}\|_{\dot{H}^{\alpha,q}}=c\|\Delta_{h}v_{h}\|_{\dot{H}^{\alpha-2,q}}.$
To estimate $\|\Delta_{h}v_{h}\|_{\dot{H}^{\alpha-2,q}}$, we use the dual characterization
$\|\Delta_{h}v_{h}\|_{\dot{H}^{\alpha-2,q}}=\sup_{\begin{subarray}{c}\varphi\in\dot{H}^{2-\alpha,q^{\prime}}\\ \varphi\neq 0\end{subarray}}\frac{\langle\Delta_{h}v_{h},\varphi\rangle}{\|\varphi\|_{\dot{H}^{2-\alpha,q^{\prime}}}}.$
Since $\Delta_{h}v_{h}\in X_{h}$, we may replace $\varphi$ with $P_{h}\varphi$, yielding
$\|\Delta_{h}v_{h}\|_{\dot{H}^{\alpha-2,q}}=\sup_{\begin{subarray}{c}\varphi\in\dot{H}^{2-\alpha,q^{\prime}}\\ \varphi\neq 0\end{subarray}}\frac{\int_{\mathcal{O}}(-\Delta_{h})^{\alpha/2}v_{h}\cdot(-\Delta_{h})^{1-\alpha/2}P_{h}\varphi\,\mathrm{d}x}{\|\varphi\|_{\dot{H}^{2-\alpha,q^{\prime}}}}.$
By Hölder’s inequality and Lemma A.2, we deduce
$\|\Delta_{h}v_{h}\|_{\dot{H}^{\alpha-2,q}}\leqslant\sup_{\begin{subarray}{c}\varphi\in\dot{H}^{2-\alpha,q^{\prime}}\\ \varphi\neq 0\end{subarray}}\frac{\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}\|P_{h}\varphi\|_{\dot{H}_{h}^{2-\alpha,q^{\prime}}}}{\|\varphi\|_{\dot{H}^{2-\alpha,q^{\prime}}}}\leqslant c\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}}.$
Combining the above estimates, we conclude
$\|v_{h}\|_{L^{\infty}}\leqslant c\|v_{h}\|_{\dot{H}_{h}^{\alpha,q}},$
which completes the proof of (ii), and hence the lemma.
$\blacksquare$
A.2 Proof of Lemma 4.9
Property (i) follows directly from Proposition 2.5 and Equation (2.128) in [37], in light of Lemma A.1. To prove property (ii), let $\alpha\in[0,2]$, $q\in(1,\infty)$, and $t>0$ be arbitrary. By Lemma A.3, the desired estimate for the second term on the left-hand side of Eq. 4.29 is already established. It therefore remains to show that
$\lVert{e^{t\Delta_{h}}P_{h}-P_{h}e^{t\Delta}}\rVert_{\mathcal{L}(\dot{H}^{\alpha,q},L^{q})}\leqslant c\min\left\{h^{\alpha},\frac{h^{2}}{t^{1-\alpha/2}}\right\}.$ (A.2)
This estimate is well known and can be derived using arguments found in [31, Theorem 6.10, Lemma 20.7]. For the sake of completeness, we provide a concise proof of (A.2) in two steps.
Step 1. We establish the estimate
$\left\|e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v\right\|_{L^{q}}\leqslant ch^{2}\|v\|_{\dot{H}^{2,q}},\quad\forall v\in\dot{H}^{2,q}.$ (A.3)
Fix $v\in\dot{H}^{2,q}$. Utilizing the Laplace inversion formula, we represent the semigroups as:
$e^{t\Delta}v$ $=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}\lambda^{-1}(\lambda-\Delta)^{-1}\Delta v\,\mathrm{d}\lambda+v,$
$e^{t\Delta_{h}}P_{h}v$ $=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}\lambda^{-1}(\lambda-\Delta_{h})^{-1}\Delta_{h}P_{h}v\,\mathrm{d}\lambda+P_{h}v,$
where the contour $\Gamma$, oriented such that the negative real axis lies to its left, is defined by
$\Gamma=\left\{re^{\pm i\theta}\mid\,r\in[t^{-1},\infty)\right\}\cup\left\{t^{-1}e^{i\omega}|\,\omega\in[-\theta,\theta]\right\},$
with $\theta\in(\pi/2,\pi)$ and $i$ denoting the imaginary unit. Subtracting these expressions yields:
$e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}\lambda^{-1}\left[(\lambda-\Delta_{h})^{-1}\Delta_{h}P_{h}-P_{h}(\lambda-\Delta)^{-1}\Delta\right]v\,\mathrm{d}\lambda.$ (A.4)
A direct computation establishes the identity
$\left[(\lambda-\Delta_{h})^{-1}\Delta_{h}P_{h}-P_{h}(\lambda-\Delta)^{-1}\Delta\right]v=(\lambda-\Delta_{h})^{-1}\Delta_{h}\left(P_{h}\Delta^{-1}-\Delta_{h}^{-1}P_{h}\right)\lambda(\lambda-\Delta)^{-1}\Delta v.$
This implies
$\left\|\left((\lambda-\Delta_{h})^{-1}\Delta_{h}P_{h}-P_{h}(\lambda-\Delta)^{-1}\Delta\right)v\right\|_{L^{q}}$
[…]
where the second inequality follows from Lemma A.1 and Eq. 4.25 with $\alpha=2$, and the third inequality uses the uniform boundedness of $\|\lambda(\lambda-\Delta)^{-1}\|_{\mathcal{L}(L^{q})}$ for $\lambda\in\Gamma$. Applying this bound to the integral representation in (A.4), we deduce
$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant ch^{2}\lVert{v}\rVert_{\dot{H}^{2,q}}\int_{\Gamma}\big{|}e^{t\lambda}\big{|}\,\frac{|\mathrm{d}\lambda|}{|\lambda|}\leqslant ch^{2}\lVert{v}\rVert_{\dot{H}^{2,q}},$
which establishes (A.3).
Step 2. For arbitrary $v\in L^{q}$, the Laplace inversion formula gives
$e^{t\Delta}v$ $=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}(\lambda-\Delta)^{-1}v\,\mathrm{d}\lambda,$
$e^{t\Delta_{h}}P_{h}v$ $=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}(\lambda-\Delta_{h})^{-1}P_{h}v\,\mathrm{d}\lambda,$
where $\Gamma$ is the contour defined in Step 1. Therefore, the difference is
$e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v$ $=\frac{1}{2\pi i}\int_{\Gamma}e^{t\lambda}\bigl{[}(\lambda-\Delta_{h})^{-1}P_{h}-P_{h}(\lambda-\Delta)^{-1}\bigr{]}v\,\mathrm{d}\lambda.$
Mimicking the derivation of (A.3) leads to the low-regularity estimate
$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant ch^{2}t^{-1}\lVert{v}\rVert_{L^{q}},\quad\forall v\in L^{q}.$ (A.5)
Applying complex interpolation between (A.3) and (A.5) yields
$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant ch^{2}t^{\alpha/2-1}\lVert{v}\rVert_{\dot{H}^{\alpha,q}},\quad\forall v\in\dot{H}^{\alpha,q}.$ (A.6)
Simultaneously, the following stability estimates hold for all $v\in L^{q}$:
$\lVert{e^{t\Delta_{h}}P_{h}v}\rVert_{L^{q}}$ $\leqslant c\lVert{P_{h}v}\rVert_{L^{q}}\leqslant c\lVert{v}\rVert_{L^{q}},$
$\lVert{P_{h}e^{t\Delta}v}\rVert_{L^{q}}$ $\leqslant c\lVert{e^{t\Delta}v}\rVert_{L^{q}}\leqslant c\lVert{v}\rVert_{L^{q}}.$
These rely on the boundedness of $P_{h}$ on $L^{q}$ (see Eq. 4.19), the boundedness of $e^{t\Delta_{h}}$ on $L^{q}$ (see Eq. 4.28), and the boundedness of $e^{t\Delta}$ on $L^{q}$ (see Lemma 4.2). Consequently,
$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant c\lVert{v}\rVert_{L^{q}},\quad\forall v\in L^{q}.$ (A.7)
Complex interpolation between (A.3) and (A.7) then gives
$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant ch^{\alpha}\lVert{v}\rVert_{\dot{H}^{\alpha,q}},\quad\forall v\in\dot{H}^{\alpha,q}.$ (A.8)
Combining estimates (A.6) and (A.8), we conclude that
$\lVert{e^{t\Delta_{h}}P_{h}v-P_{h}e^{t\Delta}v}\rVert_{L^{q}}\leqslant c\min\left\{h^{\alpha},\,h^{2}t^{\alpha/2-1}\right\}\lVert{v}\rVert_{\dot{H}^{\alpha,q}},\quad\forall v\in\dot{H}^{\alpha,q},$
which establishes (A.2). This completes the proof of property (ii).
$\blacksquare$